### v1

# EquiReg: Equivariance Regularized Diffusion for Inverse Problems

Abstract
Diffusion models represent the state-of-the-art for solving inverse problems such as image restoration tasks. In the Bayesian framework, diffusion-based inverse solvers incorporate a likelihood term to guide the prior sampling process, generating data consistent with the posterior distribution. However, due to the intractability of the likelihood term, many current methods rely on isotropic Gaussian approximations, which lead to deviations from the data manifold and result in inconsistent, unstable reconstructions. We propose Equivariance Regularized (EquiReg) diffusion, a general framework for regularizing posterior sampling in diffusion-based inverse problem solvers. EquiReg enhances reconstructions by reweighting diffusion trajectories and penalizing those that deviate from the data manifold. We define a new distribution-dependent equivariance error, empirically identify functions that exhibit low error for on-manifold samples and higher error for off-manifold samples, and leverage these functions to regularize the diffusion sampling process. When applied to a variety of solvers, EquiReg outperforms state-of-the-art diffusion models in both linear and nonlinear image restoration tasks, as well as in reconstructing partial differential equations.
Figure 1: Ground truth, measurement, baseline, and reconstruction. R1 (left-to-right): super-resolution and motion deblur. R2: box and random inpainting. R3: PDE solving and phase retrieval.
1 Introduction
Inverse problems consider the problem of recovering an unknown signal ${\bm{x}}^{*}\in\mathbb{R}^{d}$ from undersampled noisy measurements:
${\bm{y}}={\mathcal{A}}({\bm{x}}^{*})+\bm{\nu}\in\mathbb{R}^{m},$ (1)
where ${\mathcal{A}}$ is a known measurement operator, and $\bm{\nu}$ is an unknown noise [1]. Inverse problems are widely applied and studied in science and engineering, including imaging science [2], medical imaging [3, 4], image restorations [5, 6, 7, 8], computational neuroscience [9], seismology [10], radar imaging [11, 12], source separation [13, 14, 15], and astrophotography [16].
Figure 2: Equivariance Regularized (EquiReg) diffusion for solving inverse problems. a) EquiReg regularizes the posterior sampling trajectory, addressing the errors introduced due to the posterior mean expectation. b) The EquiReg loss function is lower for on- or near-manifold (clean) data and higher for off- or far-from-manifold (blurred and missing pixel) data.
Figure 3: Illustration of off-manifold posterior expectation. a) Manifold interpretation. While samples from $p({\bm{x}}_{0}|{\bm{x}}_{t})$ may lie on the manifold, their posterior expectation, being a linear combination, can fall off the manifold. b) Distribution interpretation. In a bimodal distribution, the posterior expectation lies in a low probability region.
Inverse problems are ill-posed, i.e., the inversion process can have many solutions; hence, they require prior information about the desired solution [17]. Classical approaches use regularization techniques to improve the stability and offer uniqueness of the inversion, formulated as $\min_{{\bm{x}}\in\mathbb{R}^{d}}\ \|{\bm{y}}-{\mathcal{A}}({\bm{x}})\|^{2}+\mathcal{R}({\bm{x}})$, where the first term encapsulates the forward model error and $\mathcal{R}$ is a function regularizing the solution. Bayesian inference formulates the inverse problem as finding ${\bm{x}}$ maximizing the posterior distribution $p({\bm{x}}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}})p({\bm{x}})$ where $p({\bm{y}}|{\bm{x}})$ is the likelihood of the measurements given the signal, and $p({\bm{x}})$ is a prior characterizing the signal structure [18]. Sparsity [19, 20, 21, 22, 23], low-rank [24, 25], and total variation [26] are examples of hand-crafted regularizers, imposing a prior model.
This paper focuses on methods that leverage unconditionally pre-trained score-based generative diffusion models as learned priors [27, 28, 29, 30]. These methods define a sequential noising process ${\bm{x}}_{0}\sim p_{\mathrm{data}}\to{\bm{x}}_{t}\to{\bm{x}}_{T}\sim p_{T}({\bm{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$ and a reverse denoising process parameterized by a neural network score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ [31, 30]. During sampling, these approaches incorporate gradient signals carrying likelihood-related information to solve inverse problems. Applications include image restoration [32, 33, 34, 35, 36, 37, 38, 39, 40, 41], medical imaging [42, 43, 44, 45, 46, 47, 48, 49], and solving partial differential equations (PDEs) [50, 51, 52, 53, 54, 55, 56, 57].
Solving inverse problems with diffusion [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 58, 59] requires computing the conditional score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t}|{\bm{y}})}$, decomposed into $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$. This introduces challenges, as the likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$ is only computationally tractable when $t=0$. To handle the intractable likelihood for $t>0$, many state-of-the-art methods approximate the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ with the isotropic Gaussian distribution [32, 38, 39, 40, 60], where the distribution expectation ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ is computed via Tweedie’s formula using the optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$.
This Gaussian approximation may be inaccurate (e.g., for a bimodal or a highly complex distribution). Moreover, this poor estimation can result in very high errors in likelihood score computation, especially when point estimations are employed as in DPS and PSLD [32, 38]. Conditional expectation, as in the posterior mean expectation, computes a linear combination of all possible ${\bm{x}}_{0}$; hence, from the manifold perspective, the posterior mean expectation may end up being located off-manifold (see Figure 3). The approximation error is further exacerbated when using latent diffusion models [61, 38], and can lead to artifacts in image restoration [32].
To address this error source, prior work explores projection-based methods that project either the diffusion state [62] or the measurement gradient [33, 41] onto the data manifold. However, each has limitations: [62] relies on a learned projector at every step, leading to accumulated errors; [33] assumes a noiseless setting; [41] uses SVD, which may not generalize beyond natural images. Other works incorporate higher-order information[63, 64, 65], but most large-scale applications still adopt the isotropic Gaussian approximation for its efficiency, scalability, and simplicity [40].
Our contributions. We propose Equivariance Regularized (EquiReg) diffusion, a symmetry-driven regularization framework for solving inverse problems with diffusion models (Figure 2).
  • Generalized regularization via Wasserstein gradient flow: We reinterpret the reversed conditional diffusion as a Wasserstein-2 gradient flow minimizing a functional over sample trajectories. This framework suggests employing a regularizer that reweights and penalizes trajectories deviating from the data manifold. This addresses errors arising in the likelihood score due to poor approximations of the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$, as well as prior score errors resulting from off-manifold trajectories. Unlike prior works, this approach provides a generalized regularization strategy over the entire reverse diffusion process. To realize such a regularizer, we seek an approach for manifold alignment via global properties of the data distribution. We take equivariance as one such global property that can enforce geometric symmetries.
  • EquiReg with distribution-dependent equivariance errors: We propose a new class of regularizers grounded in distribution-dependent equivariance error, a formalism that quantifies how symmetry violations vary depending on whether samples lie on- or off-manifold. We introduce Manifold-Preferential Equivariant (MPE) functions, which exhibit low equivariance error on the support of the data manifold (in-distribution) and higher error off-manifold (out-of-distribution). Our proposed regularization, EquiReg, is based on this MPE property.
  • Identifying manifold-preferential equivariant functions: We present two strategies for finding MPE functions: a) training induced, where equivariance emerges in encoder-decoder architectures trained with symmetry-preserving augmentations, and b) data inherent, where MPE arises from inherent data symmetries, commonly observed in physical systems.
  • Two practical equivariance regularizers: We develop two flexible plug-and-play losses, Equi and EquiCon (Equi with an implicit manifold constraint), which can be integrated into a variety of pixel- and latent-space diffusion models to improve solving inverse problems. Both regularizations guide the sampling trajectory toward symmetry-preserving regions which often lie close to the data manifold, improving posterior sampling.
  • Improved image restoration performance: We demonstrate that EquiReg significantly improves performance on linear and nonlinear image restoration tasks (Tables 1, 2 and 3). For example, Equi-DPS improves the FID of DPS [32] by $59\%$ on super-resolution (×$\times 4$). On motion deblur, EquiCon-ReSample improves the LPIPS of ReSample [37] by $51\%$. Moreover, EquiReg consistently shows improvement in performance across many noise levels (Figure 5).
  • Improved PDE reconstruction performance: We show that EquiReg enhances performance on PDE solving tasks (Table 4). For instance, Equi-FunDPS reduces the relative $\ell_{2}$ error of FunDPS [56, 57] from $17.16\%$ to $15.91\%$ on Helmholtz inverse problems and from $8.48\%$ to $7.84\%$ on Navier-Stokes inverse problems.
2 Related Works
Diffusion models for generation. Diffusion generative models [27, 28, 66, 67] are state-of-the-art in computer vision for image [68, 69] and video generation [61, 70, 71, 72, 73], with score-based methods [30] being among the most widely used. Recent work has leveraged pre-trained diffusion models for text-guided image generation [74, 75, 76, 77, 78], while others, including variational approaches [79, 80, 81], have employed diffusion models to solve inverse problems by conditioning on measurements [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 58, 59]. Another line of work uses text conditioning to improve inverse problem solutions [82]. The most widely adopted text-to-image diffusion models are latent diffusion models (LDMs) [61], which typically use classifier-free guidance [83]. Large-scale diffusion models for image [84, 85] and video generation [71] are also based on LDMs.
Diffusion models for inverse problems. Solving inverse problems with pre-trained diffusion models requires approximating the intractable likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$. Diffusion-based inverse solvers differ in how they approximate the likelihood $p_{t}({\bm{y}}|{\bm{x}}_{t})$ and integrate it with the sampling prior $p_{t}({\bm{x}}_{t})$ [63]. To make the likelihood tractable, a common assumption is an isotropic Gaussian approximation of the posterior distribution $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ [32, 86, 60, 38]. Prior work employs the empirical Bayes estimator (known as Tweedie’s formula) [87, 88, 89] to estimate the posterior expectation. Peng et al. [63] provides a unifying perspective on this approach and proposes a non-isotropic Gaussian approximation, aligning with other works using higher-order information [64, 65].
Equivariant deep learning. Equivariance is a strategy for incorporating symmetries into deep learning [90]. Prior work has applied it to graph networks [91], convolutional networks [92, 93], and Lie groups for modeling dynamical systems [94]. The extension of equivariance to probabilistic symmetries [95] has enabled the development of sample efficient diffusion models [96], with applications in molecular generation [97, 98, 99, 100], autonomous driving [101], robotics [102], crystal structure prediction [103], and audio inverse problems [104].
Prior work has used equivariance as a data prior for solving inverse problems such as super-resolution and deblurring [105], either by training the inverter networks with pre-defined equivariance properties [106], or enforcing it on learned denoisers [107]. Their empirical benefits are supported by theoretical analysis in compressed sensing, showing that equivariance enables solving inverse problems with fewer measurements [108]. In diffusion models, equivariance guidance has been applied to improve temporal consistency in video generation [70], and to break symmetry for generation [109]. These prior studies focus on denoising scores for generation, while we use equivariance to correct likelihood-induced errors in inverse problems and restoration.
3 Preliminaries
Diffusion models. Diffusion models [27, 28, 29, 66, 67] generate data via a reverse noising process. The forward noising process transforms the data sample ${\bm{x}}_{0}\sim p_{\text{data}}$ via a series of additive noise into an approximately Gaussian distribution ($p_{\text{data}}\rightarrow p_{t}\rightarrow\mathcal{N}(0,I)$ as $t\rightarrow\infty$), described by the stochastic differential equation (SDE) [30] $\mathrm{d}{\bm{x}}=-\tfrac{\beta_{t}}{2}{\bm{x}}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}{\bm{w}}$, where $\bm{w}$ is a standard Wiener process, and the drift and diffusion coefficients are parameterized by a monotonically increasing noise scheduler $\beta_{t}\in(0,1)$ in time $t$ [27]. Reversing the forward diffusion process is described by [110]
$\mathrm{d}{\bm{x}}=[-\tfrac{\beta_{t}}{2}{\bm{x}}-\beta_{t}\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}]\ \mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{\bar{w}}$ (2)
with $\mathrm{d}t$ moving backward in time. In discrete steps that would be from $T$ to 00. This reverse SDE can be used to generate (sample) data from the distribution $p_{\text{data}}$, where the unknown gradient $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}$ can be approximated by a scoring function $s_{\theta}({\bm{x}}_{t},t)$, parameterized by a neural network, and learned via denoising score matching methods [111, 31]. Solving inverse problems is described as a conditional generation where the data is sampled from the posterior $p({\bm{x}}|{\bm{y}})$, i.e.,
$\mathrm{d}{\bm{x}}=[-\tfrac{\beta_{t}}{2}{\bm{x}}\mathrm{d}t-\beta_{t}(\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})})]\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{\bar{w}}$ (3)
For solving general inverse problems where the diffusion is pre-trained unconditionally, the prior score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}$ can be estimated using $s_{\theta}({\bm{x}}_{t},t)$. However, the likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$ is only known at $t=0$, otherwise it is computationally intractable.
Challenges of training-free diffusion models for inverse problems. Given the intractability of the likelihood, $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$, training-free diffusion models account assumptions on $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$. The most popular approximation is the isotropic Gaussian approximation $\mathcal{N}({\bm{x}}_{0|t},r^{2}_{t}{\bm{I}})$ [32, 38, 60, 39, 40]. With an optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$, the posterior expectation ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ can be computed via Tweedie’s formula. While this offers an MMSE estimation of the posterior expectation, for complex distributions, there is no guarantee that the distribution $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ is concentrated around its mean. From the manifold perspective, this can be interpreted as having a posterior expectation that does not lie on the data manifold $\mathcal{M}$. Error in this estimation can propagate beyond the likelihood score into the whole posterior sampling process. Our proposed EquiReg aims to regularize the process to account for these errors.
Equivariance. An equivariant function respects symmetries under group transformations, i.e.,
Definition 3.1 (Equivariance).
The function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the actions in the group $G$ is equivariant if $S_{g}(f({\bm{z}}))=f(T_{g}({\bm{z}}))$ for all ${\bm{z}}\in\mathcal{Z}$ and $g\in G$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, and $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$, for $g\in G$.
While prior work leverages equivariance to directly incorporate symmetries into deep neural networks [90, 91, 92, 94, 112, 113], recent studies explore approximate equivariant networks to relax strict mathematical symmetries that may not fully hold in real-world data, aiming to improve performance [114]Wang et al. [114] proposes a definition of approximate equivariance (Definition 3.2[114], which reduces to strict equivariance when $\epsilon=0$ (Definition 3.1). For such functions, we define the equivariance error (Definition 3.3) to quantify the deviation from perfect symmetry.
Definition 3.2 (Approximate Equivariance).
Let $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ define the set of transformations under the symmetry group $G$, where $g\in G$. The function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ is $\epsilon$-approximate equivariant if $\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|\leq\epsilon$ for all ${\bm{z}}\in\mathcal{Z}$ and $g\in G$ where $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$.
Definition 3.3 (Equivariance Error).
The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ is $\mathcal{L}_{\text{EE}}(f)=\sup_{{\bm{z}},g}\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating on the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$ for $g\in G$. Hence, $f$ is $\epsilon$-approximate equivariant iff $\mathcal{L}_{\text{EE}}(f)<\epsilon$.
Data manifold hypothesis. Let data ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be in an ambient space of dimension $d$ with support $\mathcal{X}$ distribution. We assume that data are sampled from a low-dimensional manifold $\mathcal{M}$ [115, 116] embedded in a high-dimensional space (Assumption 3.1). This data hypothesis is popular in machine learning [117], and has been studied mathematically in the literature [118, 116, 119, 120]. Moreover, empirical evidence in image processing supports the manifold hypothesis [121, 122], and prior diffusion-based inverse solvers assume this manifold property [62, 33, 32].
Assumption 3.1 (Manifold Hypothesis).
Let ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be a data sample. The support $\mathcal{X}$ of the data distribution lies on a $k$ dimensional manifold $\mathcal{M}$ within an ambient space $\mathbb{R}^{d}$ where $k\ll d$.
4 EquiReg: Equivariance Regularized Diffusion
We first discuss our proposed generalized regularization framework for improving diffusion-based inverse solvers. Then, we introduce a new class of functions with distribution-dependent equivariance errors. Lastly, we propose two objective losses to regularize diffusion models for inverse problems.
Generalized regularization of diffusion models for inverse problems. EquiReg addresses the propagation error caused by the posterior approximation $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ using regularization. This regularization can be used as a plug-in on a diverse range of pixel-based and latent diffusion models. We explain EquiReg in the context of Wasserstein Gradient Flows.
Proposition 4.1.
[…]
The dynamics of $\rho$ remain the same if we replace $\phi({\bm{x}},t)$ with $\phi_{C}({\bm{x}},t):=\phi({\bm{x}},t)-C(t)$ for arbitrary temporal function $C(t)$. Without loss of generality, we assume $\phi_{C}({\bm{x}},t)<0$ for all ${\bm{x}}$ and $t$. In practice, the density function $p_{T-t}$ is not available and thus $\phi_{C}({\bm{x}},t)$ is approximated as $\hat{\phi}$ with $p_{T-t}({\bm{x}}_{T-t}|{\bm{y}})\approx\tilde{C}p_{T-t}({\bm{x}}_{T-t})p({\bm{y}}|\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}])$ where $\tilde{C}$ only depends on ${\bm{y}}$.
Because the conditional expectation $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}]$ is a linear combination of all candidate ${\bm{x}}_{0}$, the approximation remains relatively accurate when $T-t$ is small (i.e., ${\bm{x}}_{T-t}$ stays close to the data manifold under low noise) but may incur high error for larger $T-t$, as shown in Figure 3. To mitigate this, we reweight the contributions to the first term of $\Phi$, down-weighting unreliable estimates, and amplifying the reliable ones. The resulting reweighted functional is
[…]
where $Z_{t}={\int{e^{\frac{\mathcal{R}({\bm{x}})}{\hat{\phi}_{c}({\bm{x}},t)}}\mathrm{d}{\bm{x}}}}$ is the normalizing factor, and $\mathcal{R}({\bm{x}})$ is a positive regularization that is nearly zero near the data manifold and much larger elsewhere. Intuitively, since $\hat{\phi}_{C}<0$, the weight is nearly one for ${\bm{x}}$ near the data manifold and much smaller elsewhere.
Proposition 4.2.
(Informal) The evolution of $\rho$, the probability distribution of ${\bm{x}}_{T-t}$ driven by the practical and regularized reverse dynamics (eq. 5), is an approximation of the Wasserstein-2 gradient flow associated with minimizing $\tilde{\Phi}$.
[…]
In contrast to our work, most prior approaches focus on locally reducing likelihood error and do not adopt a generalized view of correcting the full functional being minimized. Next, we introduce distribution-dependent equivariance error and show how to design a regularizer with low error on or near the manifold and high error off-manifold.
Distribution-dependent equivariance errors. While the literature has primarily studied the equivariance properties of functions for general inputs, we propose a new equivariance error for approximately equivariant functions under the support of an input data distribution (see Definition 4.1).
Definition 4.1 (Distribution-Dependent Equivariance Error).
The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ and under the distribution $p$ is $\mathcal{L}_{\text{EE}}(f,p)=\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|]$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$, for $g\in G$.
The above definition enables us to define the equivariance error of trained function mappings for both on and off the manifold data. We leverage this definition in the next section to find functions whose equivariance error is low for on-manifold data and high elsewhere. Moreover, we introduce a new constrained distribution-dependent equivariance error, where the input is implicitly regularized to lie on the manifold $\mathcal{M}$ in addition to minimizing the equivariance error (Definition 4.2). Both equivariance errors are non-local, defined at the distribution level. When used to regularize the reverse conditional diffusion process, they are computed via local evaluations over the sampled data.
Definition 4.2 (Manifold Constrained Distribution-Dependent Equivariance Error).
Let $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ define the set of transformations under the symmetry group $G$, where $g\in G$. The manifold constrained equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ and under the data distribution $p$ is $\mathcal{L}_{\text{EE}}^{\mathcal{M}}(f,p)=\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|{\bm{z}}-h(S_{g}^{-1}(f(T_{g}({\bm{z}}))))\|]$ where $h:\mathcal{X}\rightarrow\mathcal{Z}$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$ of function $f$, and decoder-encoder composition of $f$-$h$ forms a vanishing-error autoencoder under the manifold $\mathcal{M}$ (see Appendix for the definition of vanishing-error autoencoders and their usage in prior work).
(a) Training induced equivariance.
(b) Data inherent equivariance.
Figure 4: MPE function examples.
Proposed regularization. Given the aforementioned formulated errors, we are now ready to discuss the choice the function $f$. We term a class of manifold-preferential equivariant (MPE) functions, whose equivariance error is lower for samples on the data manifold than for off-manifold samples. We consider two approaches for identifying such functions.
Training induced equivariance In this case, MPE emerges via symmetry-preserving augmentations applied during training. Prior work has studied equivariant properties of learned representations in deep networks [123], showing that data augmentations [124] and representation compression by reducing the model capacity [125] promote equivariant features even when equivariance is not explicitly built into the architecture. Importantly, the trained network is only approximately equivariant, and prior studies have sparsely noted that symmetry-preserving properties degrade for inputs deviating from in-distribution data [126]. A few studies have leveraged this emergence of MPE on trained networks for out-of-distribution detection [127, 128, 129].
Figure 4(a) illustrates the MPE property, emergent via training induced equivariance, of $\mathcal{E}$-$\mathcal{D}$ of a pre-trained variational autoencoder, currently used in LDMs [61]. Specifically, it shows that the equivariance error is lower for natural images and increases when images deviate from the clean data distribution. Based on Definitions 4.1 and 4.2, we propose Equi and EquiCon losses using a pre-trained encoder-decoder for diffusion-based inverse solvers such as DPS, PSLD, and ReSample.
$\text{Equi}_{\text{pixel}}\quad\mathcal{R}({\bm{x}}_{t})$ $=\|S_{g}(\mathcal{E}({\bm{x}}_{0|t}))-\mathcal{E}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$ (6)
$\text{Equi}_{\text{latent}}\quad\mathcal{R}({\bm{z}}_{t})$ $=\|S_{g}(\mathcal{D}({\bm{z}}_{0|t}))-\mathcal{D}(T_{g}({\bm{z}}_{0|t}))\|_{2}^{2}$
$\text{EquiCon}_{\text{latent}}\quad\mathcal{R}({\bm{z}}_{t})$ $=\|{\bm{z}}_{0|t}-\mathcal{E}(S_{g}^{-1}(\mathcal{D}(T_{g}({\bm{z}}_{0|t}))))\|_{2}^{2}$
Algorithm 1 Equi-DPS for Image Restoration Tasks
[…]
2:${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
3:for $t=T-1$ to 00 do
4:    $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
5:    ${{\bm{x}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{x}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
6:    $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
7:    ${\bm{x}}^{\prime}_{t-1}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{{\bm{x}}}_{0|t}+\tilde{\sigma}_{t}\bm{\epsilon}$
8:    ${\bm{x}}_{t-1}\leftarrow{\bm{x}}^{\prime}_{t-1}-\zeta_{t}\nabla_{{\bm{x}}_{t}}\|\bm{y}-\mathcal{A}({{\bm{x}}}_{0|t})\|^{2}_{2}$
9:    ${\bm{x}}_{t-1}\leftarrow{\bm{x}}_{t-1}-\lambda_{t}\nabla_{{\bm{x}}_{t}}\|S_{g}(\mathcal{E}({{\bm{x}}}_{0|t}))-\mathcal{E}(T_{g}({{\bm{x}}}_{0|t}))\|_{2}^{2}$
10:end for
11:return ${{\bm{x}}}_{0}$
We use the above regularization for image restoration tasks (e.g., see Algorithm 1 for Equi-DPS).
Data inherent equivariance MPE can also emerge naturally from symmetries underlying the data. This often occurs in physics systems where the combination of coefficient functions, boundary values, and solution functions of PDEs remains valid after applying an invertible coordinate transformation consistently to them. Let $\mathcal{G}(a)\mapsto u$ be a PDE operator that maps from the initial condition $a$ to the solution $u$. Then, it follows that $S_{g}(\mathcal{G}(a))=\mathcal{G}(T_{g}(a)).$
Table 1: Effect of EquiReg on ReSample [37]. Inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.01$.
[…]
Linear
Gaussian deblur ReSample 0.253 55.65 27.78 0.757
Equi-ReSample 0.197 64.86 29.08 0.825
EquiCon-ReSample 0.156 54.72 28.18 0.777
Motion deblur ReSample 0.160 40.14 30.55 0.854
Equi-ReSample 0.120 46.28 30.92 0.870
EquiCon-ReSample 0.078 37.61 30.73 0.860
Super-res. (×$\times 4$) ReSample 0.204 40.46 28.02 0.790
Equi-ReSample 0.098 43.56 29.74 0.849
EquiCon-ReSample 0.112 40.38 28.27 0.801
Box inpainting ReSample 0.198 108.30 19.91 0.807
Equi-ReSample 0.150 59.69 22.56 0.832
EquiCon-ReSample 0.171 110.70 21.04 0.815
Random inpainting ReSample 0.115 36.12 31.27 0.892
Equi-ReSample 0.047 29.88 31.47 0.908
EquiCon-ReSample 0.047 28.81 31.21 0.904
Nonlinear
HDR ReSample 0.190 49.06 24.88 0.819
Equi-ReSample 0.133 49.52 24.71 0.815
EquiCon-ReSample 0.135 49.98 24.67 0.817
Phase retrieval ReSample 0.237 97.86 27.61 0.750
Equi-ReSample 0.155 85.22 28.16 0.770
EquiCon-ReSample 0.159 88.75 28.11 0.774
Nonlinear deblur ReSample 0.188 56.06 29.54 0.842
Equi-ReSample 0.128 55.09 29.45 0.840
EquiCon-ReSample 0.125 54.62 29.55 0.843
We show that models trained on PDEs with inherent symmetries can indeed learn the equivariance property. Neural operators [130] are a popular architecture for modeling physics systems, as their function space setup can preserve the continuum of physics functions and generalize well to different resolutions. Among these, the Fourier Neural Operator (FNO [131]) has drawn the most attention. We train an FNO on the inverse problem of Navier-Stokes, which respects reflection symmetry (see details in Appendix), without explicit data augmentation. As shown in Figure 4(b), FNOs trained on non-augmented physics data yield lower equivariance error ($\text{Equi}_{\text{pixel}}$ loss: $\mathcal{R}({\bm{x}}_{t})=\|S_{g}(\text{FNO}({\bm{x}}_{0|t}))-\text{FNO}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$) on in-distribution as opposed to out-of-distribution data.
5 Results
This section provides experimental results on the performance of EquiReg for inverse problems, including linear and nonlinear image restoration tasks and reconstructing PDEs. To explore the impact of scaling the noise scheduler, the Appendix includes an experiment on DreamSampler [82], showing preliminary evidence that EquiReg can improve performance on text-to-image guidance.
Experimental setup for image restorations. We evaluate the performance of EquiReg when applied to the following three diffusion-based inverse solvers: PSLD [38], ReSample [37], and DPS [32]. Of these methods, PSLD and ReSample are latent solvers, while DPS is a pixel-based solver. We compare against several other methods including DDRM [34], MCG [33], DiffPIR [39], MPGD-AE [62], and DiffStateGrad [41]. We measure performance via perceptual similarity (LPIPS), distribution alignment (FID), pixel-wise fidelity (PSNR), and structural consistency (SSIM). We test EquiReg on two datasets: a) the FFHQ $256\times 256$ validation dataset [132], and b) the ImageNet $256\times 256$ validation dataset [133]. For pixel-based experiments, we use i) the pre-trained diffusion model from [32] on the FFHQ dataset, and ii) the pre-trained model from [68] on the ImageNet dataset. For latent diffusion experiments, we use i) the unconditional LDM-VQ-4 model [61] on FFHQ, and ii) the Stable Diffusion v1.5 [61] model on ImageNet.
Table 2: Linear inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
Method Gaussian deblur Motion deblur Super-resolution (×$\times 4$)
[…]
Latent
PSLD 0.357 106.2 22.87 0.537 0.322 84.62 24.25 0.615 0.313 89.72 24.51 0.650
Equi-PSLD (ours) 0.344 94.09 24.42 0.657 0.338 99.14 24.83 0.669 0.289 90.88 26.32 0.718
EquiCon-PSLD (ours) 0.320 83.18 24.38 0.664 0.322 89.87 25.14 0.699 0.277 79.39 26.14 0.722
Pixel-based
DDRM 0.239 92.43 24.93 0.732 - - - - 0.282 79.25 26.58 0.782
MCG 0.340 101.2 6.72 0.051 0.702 310.5 6.72 0.055 0.520 87.64 20.05 0.559
DiffPIR 0.236 59.65 27.36 - 0.255 65.78 26.57 - 0.260 65.77 26.64 -
MPGD-AE 0.150 114.9 24.42 0.664 0.120 104.5 25.72 0.723 0.168 137.7 24.01 0.670
DPS 0.145 104.8 25.48 0.709 0.132 99.75 26.75 0.754 0.191 125.4 24.38 0.675
DiffStateGrad-DPS 0.128 52.73 26.29 0.739 0.118 50.14 27.61 0.782 0.186 73.02 24.65 0.683
Equi-DPS (ours) 0.114 48.76 26.32 0.750 0.094 41.71 28.23 0.810 0.120 51.00 27.15 0.776
(a) Deblurring and super-resolution linear inverse problems.
Method Box inpainting Random inpainting
[…]
Latent
PSLD 0.158 43.02 24.22 0.819 0.246 49.77 29.05 0.809
Equi-PSLD (ours) 0.098 31.54 24.19 0.875 0.188 41.61 30.43 0.878
EquiCon-PSLD (ours) 0.092 35.07 24.26 0.884 0.204 40.75 29.99 0.847
Pixel-based
DDRM 0.207 78.62 22.26 0.801 0.218 - 25.75 -
MCG 0.309 40.11 19.97 0.703 0.286 29.26 21.57 0.751
MPGD-AE 0.138 248.7 21.59 0.753 0.172 339.0 25.22 0.732
DPS 0.133 56.89 23.10 0.799 0.113 51.32 29.63 0.849
DiffStateGrad-DPS 0.114 47.53 24.10 0.831 0.107 49.42 30.15 0.852
Equi-DPS (ours) 0.099 40.47 23.39 0.835 0.068 33.65 32.16 0.902
(b) Inpainting linear inverse problems.
Figure 5: EquiReg consistently improves performance across a range of measurement noise levels. Gaussian deblur using FFHQ $256\times 256$ for $\sigma_{{\bm{y}}}=[0,0.8]$.
We evaluate EquiReg on a variety of linear and nonlinear restoration tasks for natural images. We fix sets of $100$ images from FFHQ and ImageNet as our validation sets. All images are normalized from $[0,1]$. For the majority of experiments, we use noise level $\sigma_{{\bm{y}}}=0.05$ (we indicate $\sigma_{{\bm{y}}}$ in our tables). For linear inverse problems, we consider (1) box inpainting, (2) random inpainting, (3) Gaussian deblur, (4) motion deblur, and (5) super-resolution. We apply a random $128\times 128$ pixel box for box inpainting, and a $70\%$ random mask for random inpainting. For Gaussian and motion deblur, we use kernels of size $61\times 61$, with standard deviations of $3.0$ and $0.5$, respectively. For super-resolution, we downscale images by a factor of $4$ using a bicubic resizer. For nonlinear inverse problems, we consider (1) phase retrieval, (2) nonlinear deblur, and (3) high dynamic range (HDR). We use an oversampling rate of $2.0$ for phase retrieval, and due to instability of the task, we generate four independent reconstructions and take the best result (as also done in DPS [32], DAPS [40], and DiffStateGrad [41]). We use the default setting from [134] for nonlinear deblur, and a scale factor of 2 for HDR.
Table 3: Linear inverse problems on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
Method Gaussian deblur Motion deblur Super-resolution (x4) Box inpainting Random inpainting
[…]
PSLD 263.9 20.70 252.1 21.26 224.3 22.29 151.4 16.28 83.22 26.56
EquiCon-PSLD (ours) 214.5 22.01 196.3 22.69 198.5 22.34 137.6 19.25 65.14 27.03
Figure 6: EquiReg reduces failure cases and enhances reconstruction fidelity. Super-resolution using FFHQ $256\times 256$.
For these tasks, we adopt the pre-trained encoder-decoder $\mathcal{E}$-$\mathcal{D}$ as our MPE function. For FFHQ, the symmetry group contains reflection (across vertical axis). For ImageNet, we define a rotation group $G=\{\nicefrac{{\pi}}{{2}},\nicefrac{{3\pi}}{{2}}\}$, and uniformly at random select the group action for each sample. Finally, the loss functions given in Equation 6 are used as regularization terms.
Image restoration performance. Our results (Tables 1, 2 and 3) show that EquiReg improves the performance of several methods across a wide variety of linear and nonlinear tasks on both FFHQ and ImageNet. EquiReg significantly improves performance on box inpainting when applied to ReSample (Table 1). We attribute this improvement in part to the reduction of failure cases (Figure 6).
We observe that EquiReg achieves its largest improvements on perceptual metrics (FID and LPIPS), suggesting that it generates more realistic images that lie closer to the data manifold (see Appendix for supporting qualitative results). Moreover, EquiReg improves performance under high measurement noise (Figure 5). This result aligns with Figure 4(a), which shows the equivariance error is lower on clean images than noisy ones, indicating that EquiReg enforces an effective denoising.
PDE reconstruction tasks. We evaluate the performance of EquiReg on two important PDE problems: the Helmholtz and Navier-Stokes equations (see Appendix). The objective is to solve both forward and inverse problems in sparse sensor settings. The forward problem involves predicting the solution function or the final state using measurements from only $3\%$ of the coefficient field or the initial state. The inverse problem, conversely, aims to predict the input conditions from observations of $3\%$ of the system’s output. This task is challenging due to the nonlinearity of the equations, the complex structure of Gaussian random fields, and the sparsity of observations.
Table 4: Partially-observed PDE reconstruction.
Steps $(N)$ Helmholtz Navier-Stokes
Forward Inverse Forward Inverse
DiffusionPDE 2000 12.64% 19.07% 3.78% 9.63%
FunDPS 500 2.13% 17.16% 3.32% 8.48%
Equi-FunDPS (ours) 500 2.12% 15.91% 3.06% 7.84%
Recent studies [51, 56, 57] have demonstrated the superiority of diffusion-based approaches over deterministic single-forward methods for PDE reconstructions. DiffusionPDE [51] decomposes the conditional log-likelihood into a data prior (from a pre-trained diffusion model) and a measurement score, thereby providing solutions via conditional sampling. FunDPS [56, 57] further extends the sampling process to a more natural infinite-dimensional spaces, achieving better accuracy and speed via function space models.
We integrate EquiReg into the state-of-the-art FunDPS framework [56, 57], computing the Equi loss using data inherent equivariance derived from an FNO trained on the corresponding inverse problem. For our experiments, we use reflection symmetry (i.e., flipping along the $y=x$ axis), and observe no significant performance difference when using other transformations such as rotations or alternating flips. As shown in Table 4, Equi-FunDPS improves performance, measured by relative $\ell_{2}$ loss—across various tasks, especially in inverse problems where a strong data prior is critical.
6 Conclusion
We introduce Equivariance Regularized (EquiReg) diffusion for solving inverse problems. EquiReg corrects errors introduced in posterior approximations by regularizing the sampling trajectories so they lie closer to the data manifold. Our method applies to many inverse problem settings across diverse domains and is flexible to the choice of diffusion model and sampling method.
Limitations and future work. EquiReg’s effectiveness relies on the quality of the pre-trained diffusion prior. In addition, we do not address the approximations of the underlying diffusion models, but we regularize them for improved performance. Moreover, applying EquiReg requires task-specific regularizer design: one must choose an appropriate symmetry group and identify suitable MPE functions for each problem setting. As a result of this task-specific design, EquiReg can be explored in many other problem settings where equivariance is training induced or data inherent. We leave this for future work.
7 Acknowledgements and Contributions
Aditi worked on this project as part of her bachelor’s thesis. Anima was supported by Bren Chair Professor and AI2050 Senior Fellow.
Bahareh proposed and led the project. Aditi and Rayhan implemented the method and executed all the image restoration experiments. Abbas conducted the text-to-image experiments, and Jiachen performed the PDE reconstruction experiments. Bahareh wrote the original draft of the paper with Aditi, Rayhan, and Jiachen writing the results section. The project was originally based on training-induced MPE functions, proposed by Bahareh. Later versions of the project include data-inherent MPE functions, proposed by Jiachen; this is used for the PDE experiments. Bahareh, Abbas, Jiachen, and Chuwei collaborated on formalizing the theory, and the final formulation of Wasserstein gradient flow was proposed and proven by Chuwei. All authors provided editorial comments and feedback on the manuscript. Anima supervised the project and provided valuable feedback.
Appendices
We provide our source code when EquiReg is implemented on DPS. This supplementary materials contain the following:
  • Appendix A includes additional experiments on text-to-image guidance. We regularize DreamSampler [82] with EquiReg for an improved performance (see Figures 7, 8, 9 and 10).
  • Appendix B includes additional image restoration experiments for ImageNet. We show that EquiReg improves LatentDAPS and DPS across several different inverse problems (Table 5, Table 6).
  • Appendix C reports standard deviations on experiments done on PSLD (see Table 7).
  • Appendix D includes qualitative analysis on the performance of methods with and without EquiReg. Results show a reduction of artifacts and an improved perceptual quality of the solution. This section also includes the equivariance error of a pre-trained encoder used in EquiReg (Figure 11(a)), an equivariant loss vs. Gaussian noise plot (Figure 12), and a histogram of Equi’s improvement on DPS (Figure 13).
  • Appendix E demonstrates EquiReg implementation for PSLD, ReSample, DPS, and LatentDAPS (Algorithm 2, Algorithm 3, Algorithm 4, Algorithm 5, Algorithm 6, Algorithm 7). It also contains information about the EquiReg hyperparameters for image restoration tasks.
  • Appendix F contains information on the PDE reconstruction experiment. It discusses the equations along with implementation details and hyperparameters.
  • Appendix G provides theoretical proof of Propositions 4.1 and 4.2.
  • Appendix H contains additional background information on solving inverse problems, vanishing-error autoencoders, and equivariance.
  • Appendix I discloses computing resources used to conduct the experiments.
  • Appendix J credits code assets used for our experiments.
  • Appendix K discusses the broader impacts of this paper, the developed method, the conducted experiments, and their overall implications.
  • Appendix L concludes the appendix with a “responsible release” statement.
Appendix A EquiReg for Text-to-Image Guidance
Given the “source” image, DreamSampler [82] is asked to transform the source image using the prompt. Applying EquiReg to DreamSampler, we observe perceptual improvement of generated images as well as artifact reduction. Figure 7 shows a cat, as the source image, being transformed into “corgi”, “croissant”, “cat statue”, and “fluffy dog”. Equi-DreamSampler generates more realistic images than DreamSampler (e.g., EquiReg modifies the three front legged corgi to two front legs).
Figure 7: EquiReg helps the text-guided DreamSampler generate more realistic images.
Figure 8: Impact of EquiReg parameter $\lambda$, implicit acceleration, and introduction of more image details. Women, red hair, smiling.
We have also observed an implicit acceleration of image generation when EquiReg is imposed (Figure 8). Equi-DreamSampler with $50$ DDIM steps can generate images that are only possible with DreamSampler when the DDIM steps are increased. We attribute this to EquiReg’s ability to generate images that are closer to the data manifold. For example, the increase of DDIM steps in DreamSampler (from $50$ to $75$ to $100$) has a relatively similar effect to increasing the EquiReg regularizer $\lambda$ at a fixed $50$ DDIM steps. Figure 8 shows that increasing the regularization $\lambda$ results in addition of a car in the background. For DreamSampler, an early notion of the car seat in the background start to arise only when DDIM is increased to $100$ (see also Figures 9 and 10).
Figure 9: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Boy, curly hair.
Figure 10: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Glasses.
Appendix B Additional Experiments on Image Restoration Tasks
We show two additional results. EquiReg improves LatentDAPS performance on ImageNet for box inpainting and motion deblur (Table 5). Moreover, Equi-DPS performs much better than DPS, increasing the PSNR from $20.70$ to $23.45$ dB for super-resolution on ImageNet (Table 6).
Table 5: Equi-LatentDAPS vs. LatentDAPS on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
Method Box inpainting Motion deblur
[…]
LatentDAPS 0.310 20.19 0.305 25.08
Equi-LatentDAPS (ours) 0.299 20.67 0.286 25.86
Table 6: Equi-DPS vs. DPS on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
Method Super-resolution (x4) Random inpainting
[…]
DPS 169.6 20.70 94.83 25.96
Equi-DPS (ours) 102.7 23.45 55.87 28.32
Appendix C Additional Information on Image Restoration Tasks
We report standard deviation on all linear tasks for PSLD, Equi-PSLD, and EquiCon-PSLD (Table 7).
Table 7: Reporting standard deviations. Inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{y}=0.05$.
[…]
Gaussian deblur PSLD 0.357 (0.060) 22.87 (1.67) 0.537 (0.094)
Equi-PSLD (ours) 0.344 (0.055) 24.42 (1.66) 0.657 (0.076)
EquiCon-PSLD (ours) 0.320 (0.048) 24.38 (1.55) 0.664 (0.068)
Motion deblur PSLD 0.322 (0.048) 24.25 (1.47) 0.615 (0.075)
Equi-PSLD (ours) 0.338 (0.052) 24.83 (1.75) 0.669 (0.068)
EquiCon-PSLD (ours) 0.322 (0.054) 25.14 (1.62) 0.699 (0.064)
Super-res. (×$\times 4$) PSLD 0.313 (0.098) 24.51 (2.98) 0.650 (0.140)
Equi-PSLD (ours) 0.289 (0.088) 26.32 (2.24) 0.718 (0.104)
EquiCon-PSLD (ours) 0.277 (0.093) 26.14 (2.74) 0.722 (0.119)
Box inpainting PSLD 0.158 (0.023) 24.22 (2.80) 0.819 (0.031)
Equi-PSLD (ours) 0.098 (0.019) 24.19 (2.67) 0.875 (0.027)
EquiCon-PSLD (ours) 0.092 (0.021) 24.26 (2.80) 0.884 (0.029)
Random inpainting PSLD 0.246 (0.053) 29.05 (1.98) 0.809 (0.049)
Equi-PSLD (ours) 0.188 (0.034) 30.43 (1.76) 0.878 (0.027)
EquiCon-PSLD (ours) 0.204 (0.051) 29.99 (2.06) 0.847 (0.047)
Appendix D Visualizations for Image Restoration Experiments
(a) The equivariance error of the encoder is lower on clean, natural images than corrupted ones.
(b) Example visualizations of used images.
Figure 11: Training induced equivariance for a pre-trained encoder. Experimental setup corresponds to Figure 4(a) in the main text, which uses a pre-trained decoder.
Figure 12: Equivariance error of a pre-trained decoder over different Gaussian noise levels $\sigma$. We apply different Gaussian noise levels to clean images to see the effect of noise on equivariance error. We observe that as Gaussian noise increases, equivariance error increases steadily.
Figure 13: Histogram of EquiReg improvement for DPS. Super-resolution using FFHQ $256\times 256$.
(a) Box inpainting.
(b) Random inpainting.
(c) Gaussian deblur.
(d) Super-resolution (×$\times 4$).
Figure 14: Qualitative comparison of EquiCon-PSLD and PSLD on FFHQ 256 $\times$ 256.
(a) Random inpainting.
(b) Super-resolution (×$\times 4$).
Figure 15: Qualitative comparison of Equi-DPS and DPS on ImageNet 256 $\times$ 256.
Appendix E Implementation Details for Image Restoration Tasks
Hyperparameters. Our method introduces a single hyperparameter $\lambda$ that controls the amount of regularization applied. Below we include a table detailing the use of this hyperparameter in the main experiments (Table 8). For majority of experiments, we keep $\lambda$ constant throughout iterations. However, we note that in our Equi-DPS and Equi-LatentDAPS experiments, we scale $\lambda$ by the noise scheduler at every iteration. For all unscaled experiments, we employ early stopping, setting $\lambda=0$ for the last 10% of sampling.
Table 8: Equivariance regularization weight $\lambda$ used in main experiments.
Method Box Inpainting Random Inpainting Gaussian Deblur Motion Deblur Super-resolution (×$\times 4$)
FFHQ $256\times 256$
Equi-PSLD 0.05 0.05 0.03 0.03 0.02
EquiCon-PSLD 0.01 0.01 0.01 0.01 0.01
Equi-ReSample 0.03 0.05 0.02 0.02 0.05
EquiCon-ReSample 0.001 0.001 0.001 0.001 0.001
Equi-DPS 0.0001 0.001 0.001 0.001 0.1
ImageNet $256\times 256$
EquiCon-PSLD 0.0015 0.05 0.06 0.07 0.001
PSLD. We integrate EquiReg into PSLD by simply adding an additional gradient update step using our regularization term (Algorithm 2, Algorithm 3).
In our experiments, we use the official PSLD implementation from Rout et al. [38], running with its default settings to reproduce the baseline results. We note that in our code, we do not square the norm when computing the gradient, aligning with PSLD’s implementation.
Algorithm 2 Equi-PSLD for Image Restoration Tasks
1:$T,\bm{y},\{\eta_{t}\}_{t=1}^{T},\{\gamma_{t}\}_{t=1}^{T},\{\tilde{\sigma}_{t}\}_{t=1}^{T}$
2:$\mathcal{E},\mathcal{D},\mathcal{A}{\bm{x}}_{0}^{\ast},\mathcal{A},\bm{s}_{\theta},$ $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
3:${\bm{z}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
4:for $t=T-1$ to 00 do
5:    $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{z}}_{t},t)$
6:    ${{\bm{z}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{z}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
7:    $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
8:    ${\bm{z}}^{\prime}_{t-1}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{z}}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{{\bm{z}}}_{0|t}+\tilde{\sigma}_{t}\bm{\epsilon}$
9:    ${\bm{z}}^{\prime\prime}_{t-1}\leftarrow{\bm{z}}^{\prime}_{t-1}-\eta_{t}\nabla_{{\bm{z}}_{t}}\|\bm{y}-\mathcal{A}(\mathcal{D}({{\bm{z}}}_{0|t}))\|^{2}_{2}$
10:    ${\bm{z}}_{t-1}\leftarrow{\bm{z}}^{\prime\prime}_{t-1}-\gamma_{t}\nabla_{{\bm{z}}_{t}}\|{{\bm{z}}}_{0|t}-\mathcal{E}(\mathcal{A}^{T}\mathcal{A}{\bm{x}}_{0}^{\ast}+(\bm{I}-\mathcal{A}^{T}\mathcal{A})\mathcal{D}({{\bm{z}}}_{0|t}))\|^{2}_{2}$
11:    ${\bm{z}}_{t-1}\leftarrow{\bm{z}}_{t-1}-\lambda_{t}\nabla_{{\bm{z}}_{t}}\|S_{g}(\mathcal{D}({{\bm{z}}}_{0|t}))-\mathcal{D}(T_{g}({{\bm{z}}}_{0|t}))\|_{2}^{2}$
12:end for
13:return $\mathcal{D}({{\bm{z}}}_{0|t})$
Algorithm 3 EquiCon-PSLD for Image Restoration Tasks
1:$T,\bm{y},\{\eta_{t}\}_{t=1}^{T},\{\gamma_{t}\}_{t=1}^{T},\{\tilde{\sigma}_{t}\}_{t=1}^{T}$
2:$\mathcal{E},\mathcal{D},\mathcal{A}{\bm{x}}_{0}^{\ast},\mathcal{A},\bm{s}_{\theta},$ $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
3:${\bm{z}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
4:for $t=T-1$ to 00 do
5:    $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{z}}_{t},t)$
6:    ${{\bm{z}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{z}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
7:    $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
8:    ${\bm{z}}^{\prime}_{t-1}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{z}}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{{\bm{z}}}_{0|t}+\tilde{\sigma}_{t}\bm{\epsilon}$
9:    ${\bm{z}}^{\prime\prime}_{t-1}\leftarrow{\bm{z}}^{\prime}_{t-1}-\eta_{t}\nabla_{{\bm{z}}_{t}}\|\bm{y}-\mathcal{A}(\mathcal{D}({{\bm{z}}}_{0|t}))\|^{2}_{2}$
10:    ${\bm{z}}_{t-1}\leftarrow{\bm{z}}^{\prime\prime}_{t-1}-\gamma_{t}\nabla_{{\bm{z}}_{t}}\|{{\bm{z}}}_{0|t}-\mathcal{E}(\mathcal{A}^{T}\mathcal{A}{\bm{x}}_{0}^{\ast}+(\bm{I}-\mathcal{A}^{T}\mathcal{A})\mathcal{D}({{\bm{z}}}_{0|t}))\|^{2}_{2}$
11:    ${\bm{z}}_{t-1}\leftarrow{\bm{z}}_{t-1}-\lambda_{t}\nabla_{{\bm{z}}_{t}}\|{{\bm{z}}}_{0|t}-\mathcal{E}(S_{g}^{-1}(\mathcal{D}(T_{g}({{\bm{z}}}_{0|t}))))\|_{2}^{2}$
12:end for
13:return $\mathcal{D}({{\bm{z}}}_{0|t})$
ReSample. We integrate EquiReg into ReSample by adding our regularization term into the hard data consistency step (Algorithm 4, Algorithm 5). We note that the ReSample algorithm employs a two-stage approach; initially, it performs pixel-space optimization, and later it performs latent-space optimization. We apply EquiReg in the latent-space optimization stage.
In our experiments, we use the official ReSample implementation from Song et al. [37], running with its default settings to reproduce the baseline results.
Algorithm 4 Equi-ReSample for Image Restoration Tasks
1:Measurements ${\bm{y}}$, $\mathcal{A}(\cdot)$, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function ${\bm{s}}_{\theta}(\cdot,t)$, Pretrained LDM Parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\gamma$ to control $\sigma_{t}^{2}$, Time steps to perform resample $C$,  $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
2:${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$triangleright Initial noise vector
3:for $t=T-1,\ldots,0$ do
4:    $\bm{\epsilon}_{1}\sim\mathcal{N}(\mathbf{0},\bm{I})$
5:    $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$triangleright Compute the score
6:    $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$triangleright Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
7:    ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$triangleright Unconditional DDIM step
8:    if $t\in C$ then $\triangleright$ ReSample time step
9:          Initialize $\hat{{\bm{z}}}_{0}({\bm{y}})$ with $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})$
10:         for  each step in gradient descent do
11:              $\bm{g}\leftarrow\nabla_{\hat{{\bm{z}}}_{0}({\bm{y}})}\frac{1}{2}\|{\bm{y}}-\mathcal{A}(\mathcal{D}(\hat{{\bm{z}}}_{0}({\bm{y}})))\|_{2}^{2}+\lambda_{t}\nabla_{\hat{{\bm{z}}}_{0}({\bm{y}})}\|S_{g}(\mathcal{D}(\hat{{\bm{z}}}_{0}({\bm{y}})))-\mathcal{D}(T_{g}(\hat{{\bm{z}}}_{0}({\bm{y}})))\|_{2}^{2}$
12:              Update $\hat{{\bm{z}}}_{0}({\bm{y}})$ using gradient $\bm{g}$
13:         end for
14:         ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$triangleright Map back to $t$
15:    else
16:         ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$triangleright Unconditional sampling if not resampling
17:    end if
18:end for
19:${\bm{x}}_{0}=\mathcal{D}({\bm{z}}_{0})$triangleright Output reconstructed image
20:return ${\bm{x}}_{0}$
Algorithm 5 EquiCon-ReSample for Image Restoration Tasks
1:Measurements ${\bm{y}}$, $\mathcal{A}(\cdot)$, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function ${\bm{s}}_{\theta}(\cdot,t)$, Pretrained LDM Parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\gamma$ to control $\sigma_{t}^{2}$, Time steps to perform resample $C$,  $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
2:${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$triangleright Initial noise vector
3:for $t=T-1,\ldots,0$ do
4:    $\bm{\epsilon}_{1}\sim\mathcal{N}(\mathbf{0},\bm{I})$
5:    $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$triangleright Compute the score
6:    $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$triangleright Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
7:    ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$triangleright Unconditional DDIM step
8:    if $t\in C$ then $\triangleright$ ReSample time step
9:          Initialize $\hat{{\bm{z}}}_{0}({\bm{y}})$ with $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})$
10:         for  each step in gradient descent do
11:              $\bm{g}\leftarrow\nabla_{\hat{{\bm{z}}}_{0}({\bm{y}})}\frac{1}{2}\|{\bm{y}}-\mathcal{A}(\mathcal{D}(\hat{{\bm{z}}}_{0}({\bm{y}})))\|_{2}^{2}+\lambda_{t}\nabla_{\hat{{\bm{z}}}_{0}({\bm{y}})}\|\hat{{\bm{z}}}_{0}({\bm{y}})-\mathcal{E}(S_{g}^{-1}(\mathcal{D}(T_{g}(\hat{{\bm{z}}}_{0}({\bm{y}})))))\|_{2}^{2}$
12:              Update $\hat{{\bm{z}}}_{0}({\bm{y}})$ using gradient $\bm{g}$
13:         end for
14:         ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$triangleright Map back to $t$
15:    else
16:         ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$triangleright Unconditional sampling if not resampling
17:    end if
18:end for
19:${\bm{x}}_{0}=\mathcal{D}({\bm{z}}_{0})$triangleright Output reconstructed image
20:return ${\bm{x}}_{0}$
DPS. Similar to PSLD, we integrate EquiReg into DPS by simply adding an additional gradient update step using our regularization term (Algorithm 6).
In our experiments, we use the official DPS implementation from Chung et al. [32], running with its default settings to reproduce the baseline results.
Algorithm 6 Equi-DPS for Image Restoration Tasks
[…]
2:${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
3:for $t=T-1$ to 00 do
4:    $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
5:    ${{\bm{x}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{x}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
6:    $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
7:    ${\bm{x}}^{\prime}_{t-1}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{{\bm{x}}}_{0|t}+\tilde{\sigma}_{t}\bm{\epsilon}$
8:    ${\bm{x}}_{t-1}\leftarrow{\bm{x}}^{\prime}_{t-1}-\zeta_{t}\nabla_{{\bm{x}}_{t}}\|\bm{y}-\mathcal{A}({{\bm{x}}}_{0|t})\|^{2}_{2}$
9:    ${\bm{x}}_{t-1}\leftarrow{\bm{x}}_{t-1}-\lambda_{t}\nabla_{{\bm{x}}_{t}}\|S_{g}(\mathcal{E}({{\bm{x}}}_{0|t}))-\mathcal{E}(T_{g}({{\bm{x}}}_{0|t}))\|_{2}^{2}$
10:end for
11:return ${{\bm{x}}}_{0}$
LatentDAPS. We integrate EquiReg into LatentDAPS by simply adding an additional gradient update step using our regularization term (Algorithm 7).
In our experiments, we use the official LatentDAPS implementation from Zhang et al. [40], running with its default settings to reproduce the baseline results.
Algorithm 7 Equi-LatentDAPS for Image Restoration Tasks
1:Latent space score model ${\bm{s}}_{\theta}$, measurement ${\bm{y}}$, noise schedule $\sigma_{t}$, $\{t_{i}\}_{i\in\{0,\ldots,N_{A}\}}$, encoder $\mathcal{E}$ and decoder $\mathcal{D}$,  $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
2:Sample ${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\sigma_{T}^{2}\bm{I})$
3:for $i=N_{A},N_{A}-1,\ldots,1$ do
4:    Compute $\hat{{\bm{z}}}_{0}^{(0)}=\hat{{\bm{z}}}_{0}({\bm{z}}_{t_{i}})$ by solving the probability flow ODE with ${\bm{s}}_{\theta}$
5:    for $j=0,\ldots,N-1$ do
6:          $\bm{g}\leftarrow\nabla_{\hat{{\bm{z}}}_{0}}\log p(\hat{{\bm{z}}}_{0}^{(j)}|{\bm{z}}_{t_{i}})+\nabla_{\hat{{\bm{z}}}_{0}}\log p({\bm{y}}|\hat{{\bm{z}}}_{0}^{(j)})+\lambda_{t}\nabla_{\hat{{\bm{z}}}_{0}}\|S_{g}(\mathcal{D}(\hat{{\bm{z}}}_{0}^{(j)}))-\mathcal{D}(T_{g}(\hat{{\bm{z}}}_{0}^{(j)}))\|_{2}^{2}$
7:         $\hat{{\bm{z}}}_{0}^{(j+1)}\leftarrow\hat{{\bm{z}}}_{0}^{(j)}+\eta_{t}\bm{g}+\sqrt{2\eta_{t}}\bm{\epsilon}_{j}$, $\bm{\epsilon}_{j}\sim\mathcal{N}(\mathbf{0},\bm{I})$
8:    end for
9:    Sample ${\bm{z}}_{t_{i-1}}\sim\mathcal{N}(\hat{{\bm{z}}}_{0}^{(N)},\sigma_{t_{i-1}}^{2}\bm{I})$
10:end for
11:return $\mathcal{D}({\bm{z}}_{0})$
Appendix F Experiment Setup for PDE Reconstructions
Helmholtz equation. The Helmholtz equation represents wave propagation in heterogeneous media:
$\nabla^{2}u(x)+k^{2}u(x)=a(x),\quad x\in(0,1)^{2},$ (7)
with $k=1$ and $u|_{\partial\Omega}=0$. Coefficient fields $a(x)$ are generated according to $a\sim\mathcal{N}(0,(-\Delta+9\mathbf{I})^{2})$. We note that this system has reflection equivariance along $x_{1}=\frac{1}{2},x_{2}=\frac{1}{2},x_{1}=x_{2}$ and rotation equivariance by $\frac{\pi}{2},\pi,\frac{3\pi}{2}$.
Navier-Stokes equations. Following the methodology of [135], we model the time evolution of a vorticity field, $u(x,t)$, governed by:
$\partial_{t}u(x,t)+{\bm{w}}(x,t)\cdot\nabla u(x,t)$ $=\nu\Delta u(x,t)+f(x),\quad x\in(0,1)^{2},\,t\in(0,T],$ (8)
$\nabla\cdot{\bm{w}}(x,t)$ $=0,\quad x\in(0,1)^{2},\,t\in[0,T],$ (9)
$u(x,0)$ $=a(x),\quad x\in(0,1)^{2},$ (10)
where ${\bm{w}}$ is the velocity field; $\nu=\frac{1}{1000}$, viscosity; and $f$, a fixed forcing term. The initial condition $a(x)$ is drawn from $\mathcal{N}(0,7^{3/2}(-\Delta+49\mathbf{I})^{-5/2})$ under periodic boundary conditions. The forcing term is $f(x)=0.1\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$. We borrow the dataset from [51]. We note that this system has a reflection symmetry along the $x_{1}=x_{2}$ axis.
Implementation details. EquiReg, as a regularizer for diffusion posterior sampling, can be adapted to many inverse solvers in a plug-and-play manner. For PDE experiments, we use the same model weights and configurations as FunDPS [57]. Error rates are calculated using the $L^{2}$ relative error between the predicted and true solutions, averaged on 100 randomly selected test samples. We provide the information on the EquiReg scaling weights in Table 9.
Table 9: EquiReg loss used in PDE experiments.
Helmholtz Navier-Stokes
Forward Inverse Forward Inverse
EquiReg Norm Type MSE L2 MSE L2
EquiReg Weight $\lambda$ 100 100 100 1000
Appendix G Theoretical Analysis
In this section, we provide the proofs for Proposition 4.1 and Proposition 4.2.
G.1 Preliminary and Notations
We first remind the readers of gradient flow under the Wasserstein-2 metric and introduce the notations related to the diffusion model.
Wasserstein Gradient Flow
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ be a functional of probability distributions. The Wasserstein gradient flow of $\mathcal{F}$ is characterized by the minimizing movement scheme (also known as JKO scheme) introduced by [136]. For a fixed time step $\tau>0$, the sequence $(\rho_{k})_{k\in\mathbb{N}}$ of probability densities is defined recursively by:
$\rho_{k+1}\in\arg\min_{\rho\in\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{\frac{1}{2\tau}W_{2}^{2}(\rho,\rho_{k})+\mathcal{F}(\rho)\right\},$
where $W_{2}$ denotes the 2-Wasserstein distance, and each $\rho_{k}$ is a probability density representing the distribution at time $t=k\tau$. In the limit $\tau\to 0$, this discrete-time scheme recovers the continuous-time gradient flow of $\mathcal{F}$ under the $W_{2}$ metric.
Diffusion Model
A diffusion model defines a forward stochastic process $({\bm{x}}_{t})_{t\in[0,T]}$ governed by the Itô SDE:
$\mathrm{d}{\bm{x}}_{t}=f({\bm{x}}_{t},t)\,\mathrm{d}t+\sqrt{\beta_{t}}\,\mathrm{d}\bm{w}_{t},$ (11)
where $\bm{w}_{t}$ is standard Brownian motion, $\beta_{t}>0$ is a time-dependent variance schedule, and $f({\bm{x}},t)$ is a drift term. For instance, $f\equiv 0$ for a variance-exploding SDE and $f({\bm{x}},t)=-\frac{\beta_{t}}{2}{\bm{x}}$ for a variance-preserving SDE defined in [30]. In this work, we carry out our analysis under a more general setting.
Assumption G.1.
The drift term is a gradient field, $f({\bm{x}},t)=\nabla h({\bm{x}},t)$ for a scalar function $h$.
This process progressively transforms an initial data distribution ${\bm{x}}_{0}\sim p_{0}$ into a tractable reference distribution (e.g., approximately a Gaussian $\mathcal{N}(0,I)$) at time $T$.
Sampling is performed by simulating the reverse-time SDE:
$\mathrm{d}{\bm{x}}_{t}=\left[f({\bm{x}}_{t},t)-\beta_{t}\nabla_{\bm{x}}\log p_{t}({\bm{x}}_{t})\right]\mathrm{d}t+\sqrt{\beta_{t}}\,\mathrm{d}\bm{\bar{w}}_{t},$ (12)
where $p_{t}$ is the marginal density of ${\bm{x}}_{t}$, and $\bm{\bar{w}}_{t}$ is a standard Brownian motion in reverse time.
In practice, the score function $\nabla_{\bm{x}}\log p_{t}({\bm{x}})$ is approximated by a neural network $s_{\theta}({\bm{x}},t)$ trained to estimate the score of the forward process. For conditional sampling, where we sample ${\bm{x}}_{0}$ given some observed variable $y$, the score is replaced by $\nabla_{\bm{x}}\log p_{t}({\bm{x}}|{\bm{y}})$ and decomposed as
$\nabla_{\bm{x}}\log p_{t}({\bm{x}}|{\bm{y}})=\nabla_{\bm{x}}\log p_{t}({\bm{x}})+\nabla_{\bm{x}}\log p_{t}({\bm{y}}|{\bm{x}}),$ (13)
based on Bayes’ rule.
To simplify notation in the sequel, we perform a time reparameterization $t=T-t^{\prime}$, so that the reverse process is written as a forward SDE over $t\in[0,T]$:
$\mathrm{d}{\bm{x}}_{t}=-\left[f({\bm{x}}_{t},T-t)-\beta_{T-t}[\nabla_{\bm{x}}\log p_{T-t}({\bm{x}}_{t})+\nabla_{\bm{x}}\log p_{t}(y|{\bm{x}}_{t})]\right]\mathrm{d}t+\sqrt{\beta_{T-t}}\,\mathrm{d}\bm{w}_{t},$ (14)
This form describes the generative process as evolving forward from $t=0$ to $t=T$, matching the usual direction of analysis in gradient flow frameworks.
G.2 Proof of Proposition 4.1
In this work, we consider Wasserstein gradient flow under the setting where the functional $\mathcal{F}$ depends on time.
Lemma G.1.
Consider a time-dependent functional $\mathcal{F}(\rho,t)=\int\rho({\bm{x}})V({\bm{x}},t)\mathrm{d}x+\int\alpha(t)\rho\log\rho\mathrm{d}{\bm{x}}$. Then the particle description of Wasserstein-2 gradient flow associated with this functional derived by JKO scheme is
$\mathrm{d}{\bm{x}}_{t}=-\nabla V({\bm{x}}_{t},t)\mathrm{d}t+{\sqrt{2\alpha(t)}}\mathrm{d}{\bm{w}}_{t}.$ (15)
Proof.
Consider the following optimization
$\min_{\rho^{\prime}}\mathcal{F}(\rho^{\prime},t+\Delta t)-\mathcal{F}(\rho,t)+\frac{1}{2\Delta t}W_{2}^{2}(\rho,\rho^{\prime}),$ (16)
where the change of density is restricted to the Liouville equation
$\partial_{t}\rho=-\nabla\cdot(\rho v({\bm{x}},t)),\ \text{and}\ \rho^{\prime}(x)=\rho(x)-\Delta t\nabla\cdot(\rho({\bm{x}})v({\bm{x}}))+o(\Delta t).$ (17)
Using the static formulation of $W_{2}$ distance, we have
$W_{2}^{2}(\rho,\rho^{\prime})=\int\rho({\bm{x}})\|{\bm{x}}-T^{*}({\bm{x}})\|^{2}\,\mathrm{d}{\bm{x}}=\Delta t^{2}\int\rho({\bm{x}})\|v^{*}({\bm{x}})\|^{2}\,\mathrm{d}{\bm{x}},$ (18)
where $T^{*}({\bm{x}})$ is the optimal transport map, and $v^{*}({\bm{x}})$ is the associated optimal velocity field.
Thus, we can rewrite the eq. 16 as
$\inf_{v}\ \mathcal{F}(\rho,t)$ $-\Delta t\int\nabla\cdot(\rho({\bm{x}})v({\bm{x}}))\,\frac{\delta\mathcal{F}(\rho,t)}{\delta\rho}({\bm{x}})\,\mathrm{d}{\bm{x}}+\Delta t\int\big{[}\rho({\bm{x}})\partial_{t}V({\bm{x}},t)+\dot{\alpha}(t)\rho\log\rho\big{]}\mathrm{d}{\bm{x}}$ (19)
$-\mathcal{F}(\rho,t)+\frac{\Delta t}{2}\int\rho({\bm{x}})\|v({\bm{x}})\|^{2}\,\mathrm{d}{\bm{x}},$ (20)
which simplifies to
$\min_{v}\ \int\rho({\bm{x}})\left\langle v({\bm{x}}),\nabla\frac{\delta\mathcal{F}(\rho,t)}{\delta\rho}({\bm{x}})\right\rangle\,\mathrm{d}{\bm{x}}+\frac{1}{2}\int\rho({\bm{x}})\|v({\bm{x}})\|^{2}\,\mathrm{d}{\bm{x}},$ (21)
since the last term in the first line of (19) does not depend on $v$. and further to
$\min_{v}\ \int\rho({\bm{x}})\left\|v({\bm{x}})+\nabla\frac{\delta\mathcal{F}(\rho,t)}{\delta\rho}({\bm{x}})\right\|^{2}\,\mathrm{d}{\bm{x}}.$ (22)
From the optimality condition of the above problem, we obtain
$v({\bm{x}},t)=-\nabla\frac{\delta\mathcal{F}(\rho,t)}{\delta\rho}({\bm{x}})=-(\nabla V({\bm{x}},t)+\alpha(t)\nabla\log\rho({\bm{x}},t)).$ (23)
The corresponding evolution of probability density is
$\partial_{t}\rho({\bm{x}},t)$ $=-\nabla\cdot(\rho({\bm{x}},t)v({\bm{x}},t))$ (24)
$=\nabla\cdot(\rho({\bm{x}},t)(\nabla V({\bm{x}},t)+\alpha(t)\frac{\nabla\rho({\bm{x}},t)}{\rho}))$ (25)
$=-\nabla\cdot(\rho({\bm{x}},t)(-\nabla V({\bm{x}},t))+\alpha(t)\Delta\rho({\bm{x}},t),$ (26)
which is exactly the Fokker-Planck equation describing the evolution of the probability density describing the particles following
$\mathrm{d}{\bm{x}}_{t}=-\nabla V({\bm{x}}_{t},t)\mathrm{d}t+{\sqrt{2\alpha(t)}}\mathrm{d}{\bm{w}}_{t}.$ (27)
Now we come back to Proposition 4.1. From eq. 14 we know that choosing
$V({\bm{x}},t)=h({\bm{x}},T-t)-\beta_{T-t}[\log p_{T-t}({\bm{x}})+\log p_{T-t}({\bm{y}}|{\bm{x}})]\ \ \text{and}\ \alpha(t)=\frac{\beta_{T-t}}{2}$ (28)
in Lemma G.1 completes the proof, where $h$ is defined in Assumption G.1.
G.3 Detailed version of Proposition 4.2
In practice, one does not have access to $\log p_{t}({\bm{y}}|{\bm{x}}_{t})$ which appears in the reverse SDE. The most popular approach is do the following approximation,
$p_{t}({\bm{y}}|{\bm{x}}_{t})=\int p({\bm{y}}|{\bm{x}}_{0})p({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}=\mathbb{E}_{{\bm{x}}_{0}\sim p({\bm{x}}_{0}|{\bm{x}}_{t})}[p({\bm{y}}|{\bm{x}}_{0})]\approx p({\bm{y}}|\ \mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]),$ (29)
which can be interpreted as exchanging two operations, the conditional expectation and the measurement $p({\bm{y}}|\cdot)$.
As discussed in the main text, since the conditional expectation is a linear combination over all possible values of ${\bm{x}}_{0}$, it may fall outside the data manifold, resulting in physically invalid samples. One of the central challenges in diffusion-based inverse sampling is guiding the sampling trajectory—generated by the reverse SDE dynamics—toward the data manifold. A common strategy is to incorporate regularization into the reverse SDE to encourage manifold adherence. In this work, building on the perspective of Wasserstein gradient flow as outlined above, we provide a novel interpretation of the role played by such regularization terms.
We show that the regularizer serves to reweight the contribution of different regions in the calculation of the underlying functional being minimized, $\Phi(\rho,t)$ defined in Proposition 4.1. Specifically, it amplifies the influence of regions where the density estimate is reliable (typically near the data manifold), while down-weighting regions with poor approximation quality of based on eq. 29, often corresponding to off-manifold samples.
[…]
$\hat{\Phi}(\rho,t)=\beta_{T-t}\big{[}\int_{{\bm{x}}\in N(\mathcal{M})}\rho({\bm{x}})\hat{\phi}({\bm{x}},t)\mathrm{d}{\bm{x}}+\int_{{\bm{x}}\notin N(\mathcal{M})}\rho({\bm{x}})\hat{\phi}({\bm{x}},t)\mathrm{d}{\bm{x}}+\frac{1}{2}\int\rho\log\rho\mathrm{d}{\bm{x}}\big{]},$ (30)
where $N(\mathcal{M})$ denotes a neighborhood of the data manifold $\mathcal{M}$. Intuitively, we aim to focus on the contribution from regions near $\mathcal{M}$, which corresponds to the first term, while down-weighting the influence of points farther away, where the approximation tends to be unreliable. For instance, we can introduce two positive weights $A\gg B$ and adopt the modified functional
$\tilde{\Phi}(\rho,t)=\beta_{T-t}\big{[}A\int_{{\bm{x}}\in N(\mathcal{M})}\rho({\bm{x}})\hat{\phi}({\bm{x}},t)\mathrm{d}{\bm{x}}+B\int_{{\bm{x}}\notin N(\mathcal{M})}\rho({\bm{x}})\hat{\phi}({\bm{x}},t)\mathrm{d}{\bm{x}}+\frac{1}{2}\int\rho\log\rho\mathrm{d}{\bm{x}}\big{]}.$ (31)
In this work, we further generalize this idea and consider a continuous weight function,
$\tilde{\Phi}(\rho,t)=\beta_{T-t}\big{[}\int\rho({\bm{x}})\hat{\phi}({\bm{x}},t)\lambda({\bm{x}})\mathrm{d}{\bm{x}}+\frac{1}{2}\int\rho\log\rho\mathrm{d}{\bm{x}}\big{]},$ (32)
where the non-negative weight $\lambda({\bm{x}})$ is large for ${\bm{x}}\in N(\mathcal{M})$ and small elsewhere.
In practice, a nonnegative regularization function $\mathcal{R}({\bm{x}})$ is introduced, ideally being nearly zero for ${\bm{x}}$ near the data manifold and much larger elsewhere. We consider the following modified functional with weight function $\lambda({\bm{x}},t):=e^{\frac{\mathcal{R}({\bm{x}})}{\hat{\phi}({\bm{x}},t)}}$,
[…]
Note that $\hat{\phi}<0$, we have that
$\mathcal{R}({\bm{x}})\approx\begin{cases}0,&{\bm{x}}\in N(\mathcal{M})\\ \gg 1,&{\bm{x}}\text{ far away from }N(\mathcal{M})\end{cases}\quad\Rightarrow\quad\lambda({\bm{x}},t)\approx\begin{cases}1,&{\bm{x}}\in N(\mathcal{M})\\ 0,&{\bm{x}}\text{ far away from }N(\mathcal{M})\end{cases}.$
Next, we consider practical algorithms based on this reweighted functional. In practice, we only have the score function instead of the function value of $\log p_{T-t}({\bm{x}})$. Thus, the Wasserstein gradient flow associated with (33) is intractable since we cannot evaluate the weight function. We consider the following approximation based on $e^{\delta}\approx 1+\delta$,
[…]
By Lemma G.1, the dynamics of ${\bm{x}}$ driven by the Wasserstein gradient flow associated with the approximated functional above is
$\mathrm{d}{\bm{x}}=[-f({\bm{x}},T-t)-\beta_{T-t}\nabla_{{\bm{x}}}\big{(}\log{p_{T-t}({\bm{x}})}+\log\hat{p}_{T-t}({\bm{y}}|{\bm{x}})+\mathcal{R}({\bm{x}})\big{)}]\mathrm{d}t+\sqrt{\beta_{T-t}}\mathrm{d}\bm{\bar{w}}.$ (36)
This completes the proof.
Remark 1.
Since $\hat{\phi}<0$, and $e^{A}\geq 1+A$ for any $A\in\mathbb{R}$, the dynamics derived by the approximated functional in (35) is evolving to minimize an upper bound of the reweighted functional $\tilde{\Phi}$.
Appendix H Additional Background Information
Solving inverse problems with deep learning prior to diffusion models. Earlier works [137, 138, 139, 140] used deep neural networks as denoisers to solve inverse problems. Furthermore, deep generative models such as variational autoencoders (VAEs) [141], and generative adversarial networks (GANs) [142] were employed. Notable applications include compressed sensing [143] and MRI [144].
Resources for Definition 4.2 on vanishing-error autoencoders. Manifold constrained distribution-dependent equivariance error uses the notion of vanishing-error autoencoders [145, 146, 62] (Definition H.1), also known as an asymptotically-trained autoencoder [146] or a perfect autoencoder [62]. Vanishing-error autoencoders have previously been employed by diffusion-based inverse solvers to preserve the diffusion process on the manifold [62].
Definition H.1 (Vanishing-Error Autoencoder).
A vanishing-error autoencoder under the manifold $\mathcal{M}$ with encoder $\mathcal{E}:\mathcal{X}\rightarrow\mathcal{Z}$ and decoder $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{X}$ with $\mathcal{Z}=\mathbb{R}^{k}$ where $k<d$, has zero reconstruction error under the support of the data distribution $\mathcal{X}$, i.e., $\forall{\bm{x}}\in\mathcal{X}\subset\mathcal{M}$, ${\bm{x}}=\mathcal{D}(\mathcal{E}({\bm{x}}))$. It follows that the decoder is surjective on the data manifold, $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{M}$ [62], and the encoder-decoder composition forms an identity map, i.e., $\forall{\bm{z}}\in\mathcal{M},{\bm{z}}=\mathcal{E}(\mathcal{D}({\bm{z}}))$.
Equivariance. Let ${\bm{z}}\in\mathbb{R}^{d}$ and ${\bm{x}}=f({\bm{z}})\in\mathbb{R}^{d}$. For rotation and reflection equivariance, the transformations $T_{g}$ and $S_{g}$ can be defined by a rotation matrix $\bm{R}\in\mathbb{R}^{d\times d}$; then, a function $f$ with the rotation equivariant property would satisfy $\bm{R}{\bm{x}}=f(\bm{R}{\bm{z}})$. For translation equivariance, the transformations would be $T_{g}({\bm{z}})={\bm{z}}+g$ and $S_{g}({\bm{x}})={\bm{x}}+g$, where $g\in\mathbb{R}^{d}$. Hence, for a translation equivariance function $f$, we would have ${\bm{x}}+g=f({\bm{z}}+g)$. For the case where the output dimension is larger than the input, $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ with $d>k$, translation equivariance can be defined up to a discrete scale, i.e., $T_{g}({\bm{z}})={\bm{z}}+g$ and $S_{g}({\bm{x}})=T_{sg}({\bm{z}})$ where $s=\nicefrac{{d}}{{k}}$. The equivariance properties of translation, rotation, and reflections, combined, are referred to as E(3) symmetries. Without reflections, the symmetries form a Euclidean group SE(3) [112, 113].
Appendix I Computing Resources
We conduct experiments on two NVIDIA GeForce RTX 4090 GPUs with 24 GB of VRAM. We note that we use pre-trained models and perform inference, so not much compute is required.
Appendix J Assets
Appendix K Broader Impacts
On the positive side, high-fidelity image restoration can improve downstream tasks in medical imaging, remote-sensing and environmental monitoring (e.g., denoising satellite observations to track pollution or deforestation). Likewise, accelerated PDE-solving via learned diffusion priors may enable faster, more accurate simulations for climate modeling, fluid-dynamics research, and engineering design. On the other hand, robust reconstruction methods could be misappropriated for privacy-invasive surveillance or to create deceptive imagery. We emphasize that our method does not amplify these existing risks.
Appendix L Responsible Release
Our approach uses only publicly available datasets and standard pre-trained diffusion models, introducing no novel dual-use or privacy risks. Consequently, no additional safeguards are required.

### v2

@@
-EquiReg: Equivariance Regularized Diffusion for Inverse Problems
+arXiv:2505.22973v2[cs.LG] 01 Mar 2026
 Abstract
 Diffusion models represent the state-of-the-art for solving inverse problems such as image restoration tasks.
-In the Bayesian framework, diffusion-based inverse solvers incorporate a likelihood term to guide the prior sampling process, generating data consistent with the posterior distribution.
-However, due to the intractability of the likelihood term, many current methods rely on isotropic Gaussian approximations, which lead to deviations from the data manifold and result in inconsistent, unstable reconstructions.
-We propose Equivariance Regularized (EquiReg) diffusion, a general framework for regularizing posterior sampling in diffusion-based inverse problem solvers.
-EquiReg enhances reconstructions by reweighting diffusion trajectories and penalizing those that deviate from the data manifold.
-We define a new distribution-dependent equivariance error, empirically identify functions that exhibit low error for on-manifold samples and higher error for off-manifold samples, and leverage these functions to regularize the diffusion sampling process.
-When applied to a variety of solvers, EquiReg outperforms state-of-the-art diffusion models in both linear and nonlinear image restoration tasks, as well as in reconstructing partial differential equations.
-Figure 1: Ground truth, measurement, baseline, and reconstruction. R1 (left-to-right): super-resolution and motion deblur. R2: box and random inpainting. R3: PDE solving and phase retrieval.
+Diffusion-based inverse solvers incorporate a likelihood term to guide prior sampling, generating data consistent with the posterior distribution.
+However, due to the intractability of the likelihood, most methods rely on isotropic Gaussian approximations, which can push estimates off the data manifold and produce inconsistent, poor reconstructions.
+We propose Equivariance Regularized (EquiReg) diffusion, a general plug-and-play framework that improves posterior sampling by penalizing trajectories that deviate from the data manifold.
+EquiReg formalizes manifold-preferential equivariant functions that exhibit low equivariance error for on-manifold samples and high error for off-manifold ones, thereby guiding sampling toward symmetry-preserving regions of the solution space.
+We highlight that such functions naturally emerge when training non-equivariant models with augmentation or on data with symmetries.
+EquiReg is particularly effective under reduced sampling and measurement consistency steps, where many methods suffer severe quality degradation.
+By regularizing trajectories toward the manifold, EquiReg implicitly accelerates convergence and enables high-quality reconstructions.
+EquiReg consistently improves performance in linear and nonlinear image restoration tasks and solving partial differential equations.
 1 Introduction
-Inverse problems consider the problem of recovering an unknown signal ${\bm{x}}^{*}\in\mathbb{R}^{d}$ from undersampled noisy measurements:
+Inverse problems aim to recover an unknown signal ${\bm{x}}^{*}\in\mathbb{R}^{d}$ from undersampled noisy measurements:
 ${\bm{y}}={\mathcal{A}}({\bm{x}}^{*})+\bm{\nu}\in\mathbb{R}^{m},$ (1)
-where ${\mathcal{A}}$ is a known measurement operator, and $\bm{\nu}$ is an unknown noise [1].
-Inverse problems are widely applied and studied in science and engineering, including imaging science [2], medical imaging [3, 4], image restorations [5, 6, 7, 8], computational neuroscience [9], seismology [10], radar imaging [11, 12], source separation [13, 14, 15], and astrophotography [16].
-Figure 2: Equivariance Regularized (EquiReg) diffusion for solving inverse problems.
-a) EquiReg regularizes the posterior sampling trajectory, addressing the errors introduced due to the posterior mean expectation.
-b) The EquiReg loss function is lower for on- or near-manifold (clean) data and higher for off- or far-from-manifold (blurred and missing pixel) data.
-Figure 3: Illustration of off-manifold posterior expectation.
-a) Manifold interpretation.
-While samples from $p({\bm{x}}_{0}|{\bm{x}}_{t})$ may lie on the manifold, their posterior expectation, being a linear combination, can fall off the manifold.
-b) Distribution interpretation.
-In a bimodal distribution, the posterior expectation lies in a low probability region.
-Inverse problems are ill-posed, i.e., the inversion process can have many solutions; hence, they require prior information about the desired solution [17].
-Classical approaches use regularization techniques to improve the stability and offer uniqueness of the inversion, formulated as $\min_{{\bm{x}}\in\mathbb{R}^{d}}\ \|{\bm{y}}-{\mathcal{A}}({\bm{x}})\|^{2}+\mathcal{R}({\bm{x}})$, where the first term encapsulates the forward model error and $\mathcal{R}$ is a function regularizing the solution.
-Bayesian inference formulates the inverse problem as finding ${\bm{x}}$ maximizing the posterior distribution $p({\bm{x}}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}})p({\bm{x}})$ where $p({\bm{y}}|{\bm{x}})$ is the likelihood of the measurements given the signal, and $p({\bm{x}})$ is a prior characterizing the signal structure [18].
-Sparsity [19, 20, 21, 22, 23], low-rank [24, 25], and total variation [26] are examples of hand-crafted regularizers, imposing a prior model.
-This paper focuses on methods that leverage unconditionally pre-trained score-based generative diffusion models as learned priors [27, 28, 29, 30].
-These methods define a sequential noising process ${\bm{x}}_{0}\sim p_{\mathrm{data}}\to{\bm{x}}_{t}\to{\bm{x}}_{T}\sim p_{T}({\bm{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$ and a reverse denoising process parameterized by a neural network score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ [31, 30].
-During sampling, these approaches incorporate gradient signals carrying likelihood-related information to solve inverse problems.
-Applications include image restoration [32, 33, 34, 35, 36, 37, 38, 39, 40, 41], medical imaging [42, 43, 44, 45, 46, 47, 48, 49], and solving partial differential equations (PDEs) [50, 51, 52, 53, 54, 55, 56, 57].
-Solving inverse problems with diffusion [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 58, 59] requires computing the conditional score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t}|{\bm{y}})}$, decomposed into $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$.
-This introduces challenges, as the likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$ is only computationally tractable when $t=0$.
-To handle the intractable likelihood for $t>0$, many state-of-the-art methods approximate the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ with the isotropic Gaussian distribution [32, 38, 39, 40, 60], where the distribution expectation ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ is computed via Tweedie’s formula using the optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$.
-This Gaussian approximation may be inaccurate (e.g., for a bimodal or a highly complex distribution).
-Moreover, this poor estimation can result in very high errors in likelihood score computation, especially when point estimations are employed as in DPS and PSLD [32, 38].
-Conditional expectation, as in the posterior mean expectation, computes a linear combination of all possible ${\bm{x}}_{0}$; hence, from the manifold perspective, the posterior mean expectation may end up being located off-manifold (see Figure 3).
-The approximation error is further exacerbated when using latent diffusion models [61, 38], and can lead to artifacts in image restoration [32].
-To address this error source, prior work explores projection-based methods that project either the diffusion state [62] or the measurement gradient [33, 41] onto the data manifold.
-However, each has limitations: [62] relies on a learned projector at every step, leading to accumulated errors; [33] assumes a noiseless setting; [41] uses SVD, which may not generalize beyond natural images.
-Other works incorporate higher-order information[63, 64, 65], but most large-scale applications still adopt the isotropic Gaussian approximation for its efficiency, scalability, and simplicity [40].
-Our contributions. We propose Equivariance Regularized (EquiReg) diffusion, a symmetry-driven regularization framework for solving inverse problems with diffusion models (Figure 2).
-  • Generalized regularization via Wasserstein gradient flow: We reinterpret the reversed conditional diffusion as a Wasserstein-2 gradient flow minimizing a functional over sample trajectories.
-This framework suggests employing a regularizer that reweights and penalizes trajectories deviating from the data manifold.
-This addresses errors arising in the likelihood score due to poor approximations of the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$, as well as prior score errors resulting from off-manifold trajectories.
-Unlike prior works, this approach provides a generalized regularization strategy over the entire reverse diffusion process.
-To realize such a regularizer, we seek an approach for manifold alignment via global properties of the data distribution.
-We take equivariance as one such global property that can enforce geometric symmetries.
-  • EquiReg with distribution-dependent equivariance errors: We propose a new class of regularizers grounded in distribution-dependent equivariance error, a formalism that quantifies how symmetry violations vary depending on whether samples lie on- or off-manifold.
-We introduce Manifold-Preferential Equivariant (MPE) functions, which exhibit low equivariance error on the support of the data manifold (in-distribution) and higher error off-manifold (out-of-distribution).
-Our proposed regularization, EquiReg, is based on this MPE property.
-  • Identifying manifold-preferential equivariant functions: We present two strategies for finding MPE functions: a) training induced, where equivariance emerges in encoder-decoder architectures trained with symmetry-preserving augmentations, and b) data inherent, where MPE arises from inherent data symmetries, commonly observed in physical systems.
-  • Two practical equivariance regularizers: We develop two flexible plug-and-play losses, Equi and EquiCon (Equi with an implicit manifold constraint), which can be integrated into a variety of pixel- and latent-space diffusion models to improve solving inverse problems.
-Both regularizations guide the sampling trajectory toward symmetry-preserving regions which often lie close to the data manifold, improving posterior sampling.
-  • Improved image restoration performance: We demonstrate that EquiReg significantly improves performance on linear and nonlinear image restoration tasks (Tables 1, 2 and 3).
-For example, Equi-DPS improves the FID of DPS [32] by $59\%$ on super-resolution (×$\times 4$).
-On motion deblur, EquiCon-ReSample improves the LPIPS of ReSample [37] by $51\%$.
-Moreover, EquiReg consistently shows improvement in performance across many noise levels (Figure 5).
-  • Improved PDE reconstruction performance: We show that EquiReg enhances performance on PDE solving tasks (Table 4).
-For instance, Equi-FunDPS reduces the relative $\ell_{2}$ error of FunDPS [56, 57] from $17.16\%$ to $15.91\%$ on Helmholtz inverse problems and from $8.48\%$ to $7.84\%$ on Navier-Stokes inverse problems.
-2 Related Works
-Diffusion models for generation.
-Diffusion generative models [27, 28, 66, 67] are state-of-the-art in computer vision for image [68, 69] and video generation [61, 70, 71, 72, 73], with score-based methods [30] being among the most widely used.
-Recent work has leveraged pre-trained diffusion models for text-guided image generation [74, 75, 76, 77, 78], while others, including variational approaches [79, 80, 81], have employed diffusion models to solve inverse problems by conditioning on measurements [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 58, 59].
-Another line of work uses text conditioning to improve inverse problem solutions [82].
-The most widely adopted text-to-image diffusion models are latent diffusion models (LDMs) [61], which typically use classifier-free guidance [83].
-Large-scale diffusion models for image [84, 85] and video generation [71] are also based on LDMs.
-Diffusion models for inverse problems.
-Solving inverse problems with pre-trained diffusion models requires approximating the intractable likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$.
-Diffusion-based inverse solvers differ in how they approximate the likelihood $p_{t}({\bm{y}}|{\bm{x}}_{t})$ and integrate it with the sampling prior $p_{t}({\bm{x}}_{t})$ [63].
-To make the likelihood tractable, a common assumption is an isotropic Gaussian approximation of the posterior distribution $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ [32, 86, 60, 38].
-Prior work employs the empirical Bayes estimator (known as Tweedie’s formula) [87, 88, 89] to estimate the posterior expectation.
-Peng et al.
-[63] provides a unifying perspective on this approach and proposes a non-isotropic Gaussian approximation, aligning with other works using higher-order information [64, 65].
-Equivariant deep learning.
-Equivariance is a strategy for incorporating symmetries into deep learning [90].
-Prior work has applied it to graph networks [91], convolutional networks [92, 93], and Lie groups for modeling dynamical systems [94].
-The extension of equivariance to probabilistic symmetries [95] has enabled the development of sample efficient diffusion models [96], with applications in molecular generation [97, 98, 99, 100], autonomous driving [101], robotics [102], crystal structure prediction [103], and audio inverse problems [104].
-Prior work has used equivariance as a data prior for solving inverse problems such as super-resolution and deblurring [105], either by training the inverter networks with pre-defined equivariance properties [106], or enforcing it on learned denoisers [107].
-Their empirical benefits are supported by theoretical analysis in compressed sensing, showing that equivariance enables solving inverse problems with fewer measurements [108].
-In diffusion models, equivariance guidance has been applied to improve temporal consistency in video generation [70], and to break symmetry for generation [109].
-These prior studies focus on denoising scores for generation, while we use equivariance to correct likelihood-induced errors in inverse problems and restoration.
-3 Preliminaries
+where $\mathcal{A}$} is a known measurement operator, and $\bm{\nu}$ is an unknown noise (Groetsch, 1993).
+Inverse problems are widely studied in science and engineering, including imaging and astrophotography.
+Inverse problems are ill-posed, i.e., the inversion process can have many solutions; hence, they require prior information about the desired solution (Kabanikhin, 2008).
+In the Bayesian formulation, the solution maximizes the posterior distribution $p({\bm{x}}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}})p({\bm{x}})$, where $p({\bm{y}}|{\bm{x}})$ is the likelihood of the measurements and $p({\bm{x}})$ is a prior describing the signal structure (Stuart, 2010).
+Examples of handcrafted priors include sparsity (Donoho, 2006) and low-rankness (Candès et al., 2011).
+Figure 1: Equivariance Regularized (EquiReg) diffusion for inverse problems.
+(left) Manifold preferential equivariance (MPE) functions whose equivariance error is lower for on-manifold and higher for off-manifold data.
+(right) EquiReg regularizes the posterior sampling trajectory for improved performance.
+It penalizes off-manifold trajectories via MPE-based regularization.
+Figure 2: EquiReg’s broad applicability.
+a-d) image restoration inverse problems and e) text-guided image generation, resulting in artifact reduction and more realistic generation.
+Here, EquiReg refers to our regularization being applied to the diffusion sampling method on the same row.
+This paper focuses on methods that leverage unconditionally pre-trained score‐based generative diffusion models as learned priors (Ho et al., 2020; Song and Ermon, 2019) with applications in image restoration (Chung et al., 2023), medical imaging (Chung et al., 2022a), and solving partial differential equations (PDEs) (Huang et al., 2024; Yao et al., 2025).
+These methods define a sequential noising process ${\bm{x}}_{0}\sim p_{\mathrm{data}}\to{\bm{x}}_{t}\to{\bm{x}}_{T}\sim p_{T}({\bm{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$ and a reverse denoising process parameterized by a neural network score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ (Vincent, 2011).
+During sampling, these approaches incorporate gradient signals carrying likelihood information to solve inverse problems.
+Solving inverse problems with diffusion (Zhang et al., 2025a; Alkhouri et al., 2025) requires computing the conditional score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t}|{\bm{y}})}$, decomposed into $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$.
+This introduces challenges, as the likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}=\nabla_{{\bm{x}}_{t}}\log{\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}}$ is only computationally tractable when $t=0$.
+To handle the likelihood for $t>0$, many methods approximate the posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ with the isotropic Gaussian distribution (Zhang et al., 2025a), where the distribution expectation is computed using the optimal denoising score (Robbins, 1956).
+The Gaussian approximation can be inaccurate for complex distributions (Figure˜3), leading to errors in likelihood computation, especially with point estimations (Chung et al., 2023).
+Since the posterior expectation is a conditional expectation, a linear combination of all possible ${\bm{x}}_{0}$, it may lie off the data manifold even when individual samples remain on it.
+These issues are further amplified in latent diffusion models (LDMs), introducing artifacts (Rout et al., 2023).
+Prior work has attempted to address this challenge via projection-based (He et al., 2024; Zirvi et al., 2025) or decoupled optimization strategies (Zhang et al., 2025a), aimed at reducing the propagation of measurement errors during sampling.
+However, they rely on the isotropic Gaussian assumption, which can lead to failures on difficult tasks or at reduced sampling steps.
+While higher-order statistics reduces errors (Boys et al., 2024), most approaches employ such approximation for its efficiency, scalability, and simplicity (Alkhouri et al., 2025), often coupled with large-scale LDMs (Peebles and Xie, 2023).
+This raises a key question: how can we ensure the reliability and practicality of conditional diffusion models under this approximation?
+Equivariance offers a natural mechanism to keep sampling trajectories close to the data manifold.
+We address this challenge with a regularization scheme that leverages equivariance to improve posterior sampling by guiding trajectories toward symmetry-preserving solution spaces.
+Prior work has enforced equivariance directly within generation or denoising processes (Chen et al., 2023a; Terris et al., 2024), with extensions to probabilistic symmetries (Bloem-Reddy et al., 2020) enabling sample efficiency (Wang et al., 2024).
+Our approach differs as follows: rather than strictly enforcing equivariance within denoising architectures, which can hinder tasks requiring symmetry breaking (Lawrence et al., 2025), we employ equivariance as a plug-and-play regularizer to guide diffusion trajectories toward the data manifold.
+Our contributions.
+We propose Equivariance Regularized (EquiReg) diffusion, an equivariance-based regularization framework for solving inverse problems with diffusion models (Figure˜1).
+EquiReg leverages equivariance to regularize likelihood-induced errors during posterior sampling, guiding diffusion trajectories toward more consistent, on-manifold solutions.
+Crucially, it employs Manifold-Preferential Equivariant (MPE) functions, which discriminate on-manifold from off-manifold data by exhibiting low equivariance error in-distribution and higher error out-of-distribution.
+Figure 3: Off-manifold posterior expectation.
+This impacts the likelihood score $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$ computation achieved via isotropic Gaussian modelling of $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$.
+We formalize that an effective regularizer should capture such a global property, and MPE functions provide a principled way to direct sampling toward plausible solutions.
+This design makes EquiReg architecture-agnostic: the regularizer operates independently of the diffusion model itself.
+With a suitable MPE function, EquiReg improves performance across models, including those with equivariant scores, where likelihood guidance may otherwise push trajectories off the manifold.
+We observe that many practical functions behave as MPEs: their equivariance error is small on the training or data manifold but grows off-manifold.
+This behavior arises in learned models trained with data augmentation, as well as in data with inherent symmetries such as those from physical systems.
+Rather than treating the degradation off-manifold as a limitation, we exploit it as a signal: equivariance error serves as a natural discriminator for identifying undesirable states during diffusion sampling.
+Building on this idea, we construct pre-trained MPEs as the foundation of our EquiReg loss.
+The choice of this function is independent of the denoiser in diffusion models and can be derived separately.
+For instance, if the diffusion architecture is itself equivariant to a set of group actions, it cannot be regularized via EquiReg using these same equivariances, as the MPE cannot discriminate between on- and off-manifold samples (the equivariance loss would remain low).
+In this case, a separate MPE with different equivariance properties can be trained.
+We explore several constructed and pre-trained MPE functions and show that EquiReg is widely robust to the choice of MPE.
+Beyond architectural flexibility, we systematically analyze the diversity properties of EquiReg-guided posterior sampling.
+We demonstrate that EquiReg offers favourable fidelity–diversity trade-offs, and as inverse problems become more ill-posed, diversity increases approximately linearly, indicating that EquiReg naturally expands posterior exploration with growing uncertainty rather than collapsing to a single mode.
+We further demonstrate that MPE functions are not rare or specialized constructs, but emerge broadly across practical neural networks.
+We empirically analyze multiple architectures, including the latent diffusion encoder (Rombach et al., 2022), CNN autoencoders trained with symmetry-based augmentations, pre-trained ResNet-50 (He et al., 2016), and CLIP encoders (Radford et al., 2021), and show that their equivariance error increases systematically as Gaussian noise pushes samples off the data manifold.
+We further show that using these MPE functions for EquiReg guidance improves solutions to inverse problems.
+Overall, this study shows that MPEs are widespread, and EquiReg leverages this intrinsic property as a principled regularization signal.
+We validate the efficacy of EquiReg through extensive experiments across diverse diffusion models, inverse problems, and datasets.
+We demonstrate that EquiReg improves perceptual image quality and remains effective in cases where baselines fail.
+We show that EquiReg improves the performance of SITCOM (Alkhouri et al., 2025) and DPS (Chung et al., 2023) when the number of measurement consistency and sampling steps are reduced, thus moving toward more efficient diffusion-based solvers.
+Our method is particularly useful when applied to LDMs.
+EquiReg reduces failure cases, and consistently improves PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DPS (Chung et al., 2023) on linear and nonlinear image restoration tasks.
+For example, EquiReg significantly improves the LPIPS (Song et al., 2023a) of ReSample by $51\%$ for motion deblur and the FID of DPS (Chung et al., 2023) by $59\%$ on super-resolution.
+We also include diversity analyses, demonstrating that EquiReg maintains diversity without collapse of single mode reconstruction.
+We extend EquiReg’s applicability to function-space diffusion models and demonstrate its added benefit for solving PDEs.
+EquiReg achieves a $7.3\%$ relative reduction in the $\ell_{2}$ error of FunDPS (Mammadov et al., 2024a; Yao et al., 2025) on the Helmholtz equation and a $7.5\%$ relative reduction on the Navier-Stokes equation.
+Lastly, we include preliminary experiments on EquiReg improving the realism and plausibility of text-guided image generation, emphasizing that the benefits of EquiReg extend beyond image restorations.
+Overall, the flexibility of EquiReg as a plug-and-play regularization framework suggests that its utility will extend well beyond the specific methods studied in this paper.
+2 Preliminaries and Related Works
 Diffusion models.
-Diffusion models [27, 28, 29, 66, 67] generate data via a reverse noising process.
-The forward noising process transforms the data sample ${\bm{x}}_{0}\sim p_{\text{data}}$ via a series of additive noise into an approximately Gaussian distribution ($p_{\text{data}}\rightarrow p_{t}\rightarrow\mathcal{N}(0,I)$ as $t\rightarrow\infty$), described by the stochastic differential equation (SDE) [30] $\mathrm{d}{\bm{x}}=-\tfrac{\beta_{t}}{2}{\bm{x}}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}{\bm{w}}$, where $\bm{w}$ is a standard Wiener process, and the drift and diffusion coefficients are parameterized by a monotonically increasing noise scheduler $\beta_{t}\in(0,1)$ in time $t$ [27].
-Reversing the forward diffusion process is described by [110]
-$\mathrm{d}{\bm{x}}=[-\tfrac{\beta_{t}}{2}{\bm{x}}-\beta_{t}\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}]\ \mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{\bar{w}}$ (2)
-with $\mathrm{d}t$ moving backward in time.
-In discrete steps that would be from $T$ to 00.
-This reverse SDE can be used to generate (sample) data from the distribution $p_{\text{data}}$, where the unknown gradient $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}$ can be approximated by a scoring function $s_{\theta}({\bm{x}}_{t},t)$, parameterized by a neural network, and learned via denoising score matching methods [111, 31].
-Solving inverse problems is described as a conditional generation where the data is sampled from the posterior $p({\bm{x}}|{\bm{y}})$, i.e.,
+Diffusion generative models (Ho et al., 2020; Song and Ermon, 2019; Sohl-Dickstein et al., 2015; Kadkhodaie and Simoncelli, 2021) are state-of-the-art in computer vision for image (Esser et al., 2024) and video generation (Brooks et al., 2024; Zhang et al., 2025b), with score-based methods (Song et al., 2021) being among the most widely used.
+Diffusion models generate data via a reverse noising process.
+The forward noising process transforms the data sample ${\bm{x}}_{0}\sim p_{\text{data}}$ via a series of additive noise into an approximately Gaussian distribution ($p_{\text{data}}\rightarrow p_{t}\rightarrow\mathcal{N}(0,I)$ as $t\rightarrow\infty$), described by the stochastic differential equation (SDE) $\mathrm{d}{\bm{x}}=-\tfrac{\beta_{t}}{2}{\bm{x}}\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}{\bm{w}}$, where $\bm{w}$ is a standard Wiener process, and the drift and diffusion coefficients are parameterized by a monotonically increasing noise scheduler $\beta_{t}\in(0,1)$ in time $t$ (Ho et al., 2020).
+Reversing the forward diffusion process is described by (Anderson, 1982)
+$\vskip-5.69054pt\mathrm{d}{\bm{x}}=[-\tfrac{\beta_{t}}{2}{\bm{x}}-\beta_{t}\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}]\ \mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{\bar{w}}$ (2)
+with $\mathrm{d}t$ moving backward in time or in discrete steps from $T$ to 0.
+This reverse SDE is used to sample data from the distribution $p_{\text{data}}$, where the unknown gradient $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}$ is approximated by a scoring function $s_{\theta}({\bm{x}}_{t},t)$, parameterized by a neural network and learned via denoising score matching methods (Hyvärinen and Dayan, 2005; Vincent, 2011).
+Solving inverse problems is described as a conditional generation where the data is sampled from the posterior $p({\bm{x}}|{\bm{y}})$:
 $\mathrm{d}{\bm{x}}=[-\tfrac{\beta_{t}}{2}{\bm{x}}\mathrm{d}t-\beta_{t}(\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{x}}_{t})}+\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})})]\mathrm{d}t+\sqrt{\beta_{t}}\mathrm{d}\bm{\bar{w}}$ (3)
@@
 However, the likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$ is only known at $t=0$, otherwise it is computationally intractable.
-Challenges of training-free diffusion models for inverse problems.
-Given the intractability of the likelihood, $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$, training-free diffusion models account assumptions on $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$.
-The most popular approximation is the isotropic Gaussian approximation $\mathcal{N}({\bm{x}}_{0|t},r^{2}_{t}{\bm{I}})$ [32, 38, 60, 39, 40].
-With an optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$, the posterior expectation ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ can be computed via Tweedie’s formula.
-While this offers an MMSE estimation of the posterior expectation, for complex distributions, there is no guarantee that the distribution $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ is concentrated around its mean.
-From the manifold perspective, this can be interpreted as having a posterior expectation that does not lie on the data manifold $\mathcal{M}$.
-Error in this estimation can propagate beyond the likelihood score into the whole posterior sampling process.
-Our proposed EquiReg aims to regularize the process to account for these errors.
-Equivariance. An equivariant function respects symmetries under group transformations, i.e.,
+Diffusion models for inverse problems.
+Solving inverse problems with pre-trained diffusion models requires approximating the intractable likelihood score $\nabla_{{\bm{x}}_{t}}\log{p_{t}({\bm{y}}|{\bm{x}}_{t})}$.
+Training-free solvers differ in how they approximate $p_{t}({\bm{y}}|{\bm{x}}_{t})$ and combine it with the prior $p_{t}({\bm{x}}_{t})$ (Peng et al., 2024).
+Since $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$, the common choice is to approximate $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ by an isotropic Gaussian $\mathcal{N}({\bm{x}}_{0|t},r_{t}^{2}{\bm{I}})$ (Chung et al., 2023; Song et al., 2023b; Zhu et al., 2023; Zhang et al., 2025a).
+With an optimal denoising score $s_{\theta}({\bm{x}}_{t},t)$, the posterior mean ${\bm{x}}_{0|t}\coloneqq\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ follows from Tweedie’s formula (Robbins, 1956; Miyasawa and others, 1961; Efron, 2011).
+While an MMSE estimate, $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ may not be concentrated around its mean for complex or multimodal distributions, leading to off-manifold solutions (see Figure˜3).
+Equivariance.
+Equivariance describes how functions transform predictably under group actions and provides a principled way to incorporate symmetry into deep learning (Bronstein et al., 2021).
+It has been applied to graphs (Satorras et al., 2021), convolutional networks (Cohen and Welling, 2016; Romero and Lohit, 2022), Lie groups for dynamical systems (Finzi et al., 2020), and diffusion models (Wang et al., 2024), with applications spanning molecular generation (Hoogeboom et al., 2022; Cornet et al., 2024), autonomous driving (Chen et al., 2023b), robotics (Brehmer et al., 2023), crystal structure prediction (Jiao et al., 2023), and audio inverse problems (Moliner et al., 2023).
+Equivariance guidance has also improved temporal consistency in video generation (Daras et al., 2024), and its role as a prior in inverse problems is theoretically supported in compressed sensing (Tachella et al., 2023).
 Definition 2.1 (Equivariance).
-The function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the actions in the group $G$ is equivariant if $S_{g}(f({\bm{z}}))=f(T_{g}({\bm{z}}))$ for all ${\bm{z}}\in\mathcal{Z}$ and $g\in G$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, and $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$, for $g\in G$.
-While prior work leverages equivariance to directly incorporate symmetries into deep neural networks [90, 91, 92, 94, 112, 113], recent studies explore approximate equivariant networks to relax strict mathematical symmetries that may not fully hold in real-world data, aiming to improve performance [114]Wang et al.
-[114] proposes a definition of approximate equivariance (Definition 3.2[114], which reduces to strict equivariance when $\epsilon=0$ (Definition 3.1).
-For such functions, we define the equivariance error (Definition 3.3) to quantify the deviation from perfect symmetry.
-Definition 3.2 (Approximate Equivariance).
-Let $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ define the set of transformations under the symmetry group $G$, where $g\in G$.
-The function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ is $\epsilon$-approximate equivariant if $\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|\leq\epsilon$ for all ${\bm{z}}\in\mathcal{Z}$ and $g\in G$ where $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$.
-Definition 3.3 (Equivariance Error).
-The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ is $\mathcal{L}_{\text{EE}}(f)=\sup_{{\bm{z}},g}\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating on the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$ for $g\in G$.
-Hence, $f$ is $\epsilon$-approximate equivariant iff $\mathcal{L}_{\text{EE}}(f)<\epsilon$.
-Data manifold hypothesis.
-Let data ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be in an ambient space of dimension $d$ with support $\mathcal{X}$ distribution.
-We assume that data are sampled from a low-dimensional manifold $\mathcal{M}$ [115, 116] embedded in a high-dimensional space (Assumption 3.1).
-This data hypothesis is popular in machine learning [117], and has been studied mathematically in the literature [118, 116, 119, 120].
-Moreover, empirical evidence in image processing supports the manifold hypothesis [121, 122], and prior diffusion-based inverse solvers assume this manifold property [62, 33, 32].
-Assumption 3.1 (Manifold Hypothesis).
-Let ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be a data sample.
-The support $\mathcal{X}$ of the data distribution lies on a $k$ dimensional manifold $\mathcal{M}$ within an ambient space $\mathbb{R}^{d}$ where $k\ll d$.
+Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$.
+A function $f:\mathcal{Z}\to\mathcal{X}$ is equivariant if for all $g\in G$ and ${\bm{z}}\in\mathcal{Z}$, $f(T_{g}({\bm{z}}))=S_{g}(f({\bm{z}}))$.
+An equivariant function preserves structure under group transformations (Definition˜2.1).
+While prior work leverages exact equivariance to encode symmetries directly into neural networks, recent studies investigate approximate equivariance to relax strict symmetry assumptions that may not hold in real-world data, aiming to improve performance (Wang et al., 2022).
+These works introduce a formal definition of approximate equivariance (Definition˜2.2) and an equivariance error to quantify deviations from perfect symmetry.
+Definition 2.2 (Approximate Equivariant Functions).
+Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$.
+A function $f:\mathcal{Z}\rightarrow\mathcal{X}$ is $\epsilon$-approximate equivariant if for all $g\in G$ and ${\bm{z}}\in\mathcal{Z}$, $\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|\leq\epsilon$.
+The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ is defined as $\sup_{{\bm{z}},g}\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|$.
+Hence, $f$ is $\epsilon$-approximate equivariant iff its error $\epsilon$.
+We use the term manifold which refers to the data manifold hypothesis (see Assumption˜H.1(Cayton and others, 2005) that assumes data is sampled from a low-dimensional manifold embedded in a high-dimensional space.
+This hypothesis is popular in machine learning (Bordt et al., 2023) and diffusion-based solvers (He et al., 2024; Chung et al., 2022b; 2023), supported by empirical evidence for imaging (Weinberger and Saul, 2006).
 3 EquiReg: Equivariance Regularized Diffusion
-We first discuss our proposed generalized regularization framework for improving diffusion-based inverse solvers.
-Then, we introduce a new class of functions with distribution-dependent equivariance errors.
-Lastly, we propose two objective losses to regularize diffusion models for inverse problems.
-Generalized regularization of diffusion models for inverse problems.
-EquiReg addresses the propagation error caused by the posterior approximation $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ using regularization.
-This regularization can be used as a plug-in on a diverse range of pixel-based and latent diffusion models.
-We explain EquiReg in the context of Wasserstein Gradient Flows.
-Proposition 4.1.
+We begin by presenting a generalized regularization framework for improving diffusion-based inverse solvers.
+We focus on the property of equivariance and introduce a new class of functions whose equivariance errors are distribution-dependent (low for on- or near-manifold samples and high for off-manifold samples).
+We leverage these functions to regularize diffusion models, guiding sampling trajectories toward better inverse solutions.
+This paper addresses the propagation error introduced by the approximation of posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ by incorporating an explicit regularization term.
+The proposed framework is general and can be applied as plug-in on a wide range of pixel and latent-space diffusion models.
+Given $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$, let $\tilde{p}_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ denote an approximation of the posterior to make the likelihood tractable.
+We formulate the regularized reverse diffusion dynamics as
 […]
-The dynamics of $\rho$ remain the same if we replace $\phi({\bm{x}},t)$ with $\phi_{C}({\bm{x}},t):=\phi({\bm{x}},t)-C(t)$ for arbitrary temporal function $C(t)$.
-Without loss of generality, we assume $\phi_{C}({\bm{x}},t)<0$ for all ${\bm{x}}$ and $t$.
-In practice, the density function $p_{T-t}$ is not available and thus $\phi_{C}({\bm{x}},t)$ is approximated as $\hat{\phi}$ with $p_{T-t}({\bm{x}}_{T-t}|{\bm{y}})\approx\tilde{C}p_{T-t}({\bm{x}}_{T-t})p({\bm{y}}|\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}])$ where $\tilde{C}$ only depends on ${\bm{y}}$.
-Because the conditional expectation $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}]$ is a linear combination of all candidate ${\bm{x}}_{0}$, the approximation remains relatively accurate when $T-t$ is small (i.e., ${\bm{x}}_{T-t}$ stays close to the data manifold under low noise) but may incur high error for larger $T-t$, as shown in Figure 3.
-To mitigate this, we reweight the contributions to the first term of $\Phi$, down-weighting unreliable estimates, and amplifying the reliable ones.
-The resulting reweighted functional is
+Algorithm 1 Equi-DPS for Inverse Problems.
 […]
-where $Z_{t}={\int{e^{\frac{\mathcal{R}({\bm{x}})}{\hat{\phi}_{c}({\bm{x}},t)}}\mathrm{d}{\bm{x}}}}$ is the normalizing factor, and $\mathcal{R}({\bm{x}})$ is a positive regularization that is nearly zero near the data manifold and much larger elsewhere.
-Intuitively, since $\hat{\phi}_{C}<0$, the weight is nearly one for ${\bm{x}}$ near the data manifold and much smaller elsewhere.
-Proposition 4.2.
-(Informal) The evolution of $\rho$, the probability distribution of ${\bm{x}}_{T-t}$ driven by the practical and regularized reverse dynamics (eq.
-5), is an approximation of the Wasserstein-2 gradient flow associated with minimizing $\tilde{\Phi}$.
 […]
-In contrast to our work, most prior approaches focus on locally reducing likelihood error and do not adopt a generalized view of correcting the full functional being minimized.
-Next, we introduce distribution-dependent equivariance error and show how to design a regularizer with low error on or near the manifold and high error off-manifold.
-Distribution-dependent equivariance errors.
-While the literature has primarily studied the equivariance properties of functions for general inputs, we propose a new equivariance error for approximately equivariant functions under the support of an input data distribution (see Definition 4.1).
-Definition 4.1 (Distribution-Dependent Equivariance Error).
-The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ and under the distribution $p$ is $\mathcal{L}_{\text{EE}}(f,p)=\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|]$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$, for $g\in G$.
-The above definition enables us to define the equivariance error of trained function mappings for both on and off the manifold data.
-We leverage this definition in the next section to find functions whose equivariance error is low for on-manifold data and high elsewhere.
-Moreover, we introduce a new constrained distribution-dependent equivariance error, where the input is implicitly regularized to lie on the manifold $\mathcal{M}$ in addition to minimizing the equivariance error (Definition 4.2).
-Both equivariance errors are non-local, defined at the distribution level.
-When used to regularize the reverse conditional diffusion process, they are computed via local evaluations over the sampled data.
-Definition 4.2 (Manifold Constrained Distribution-Dependent Equivariance Error).
-Let $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ define the set of transformations under the symmetry group $G$, where $g\in G$.
-The manifold constrained equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ and under the data distribution $p$ is $\mathcal{L}_{\text{EE}}^{\mathcal{M}}(f,p)=\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|{\bm{z}}-h(S_{g}^{-1}(f(T_{g}({\bm{z}}))))\|]$ where $h:\mathcal{X}\rightarrow\mathcal{Z}$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$ of function $f$, and decoder-encoder composition of $f$-$h$ forms a vanishing-error autoencoder under the manifold $\mathcal{M}$ (see Appendix for the definition of vanishing-error autoencoders and their usage in prior work).
-(a) Training induced equivariance.
-(b) Data inherent equivariance.
-Figure 4: MPE function examples.
-Proposed regularization.
-Given the aforementioned formulated errors, we are now ready to discuss the choice the function $f$.
-We term a class of manifold-preferential equivariant (MPE) functions, whose equivariance error is lower for samples on the data manifold than for off-manifold samples.
-We consider two approaches for identifying such functions.
-Training induced equivariance In this case, MPE emerges via symmetry-preserving augmentations applied during training.
-Prior work has studied equivariant properties of learned representations in deep networks [123], showing that data augmentations [124] and representation compression by reducing the model capacity [125] promote equivariant features even when equivariance is not explicitly built into the architecture.
-Importantly, the trained network is only approximately equivariant, and prior studies have sparsely noted that symmetry-preserving properties degrade for inputs deviating from in-distribution data [126].
-A few studies have leveraged this emergence of MPE on trained networks for out-of-distribution detection [127, 128, 129].
-Figure 4(a) illustrates the MPE property, emergent via training induced equivariance, of $\mathcal{E}$-$\mathcal{D}$ of a pre-trained variational autoencoder, currently used in LDMs [61].
-Specifically, it shows that the equivariance error is lower for natural images and increases when images deviate from the clean data distribution.
-Based on Definitions 4.1 and 4.2, we propose Equi and EquiCon losses using a pre-trained encoder-decoder for diffusion-based inverse solvers such as DPS, PSLD, and ReSample.
-$\text{Equi}_{\text{pixel}}\quad\mathcal{R}({\bm{x}}_{t})$ $=\|S_{g}(\mathcal{E}({\bm{x}}_{0|t}))-\mathcal{E}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$ (6)
-$\text{Equi}_{\text{latent}}\quad\mathcal{R}({\bm{z}}_{t})$ $=\|S_{g}(\mathcal{D}({\bm{z}}_{0|t}))-\mathcal{D}(T_{g}({\bm{z}}_{0|t}))\|_{2}^{2}$
-$\text{EquiCon}_{\text{latent}}\quad\mathcal{R}({\bm{z}}_{t})$ $=\|{\bm{z}}_{0|t}-\mathcal{E}(S_{g}^{-1}(\mathcal{D}(T_{g}({\bm{z}}_{0|t}))))\|_{2}^{2}$
-Algorithm 1 Equi-DPS for Image Restoration Tasks
+3:for t=T1t=T-1 to 0 do
+4:  $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
 […]
-2:${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
-3:for $t=T-1$ to 00 do
-4:    $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
-5:    ${{\bm{x}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{x}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
 6:  $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
@@
 8:  ${\bm{x}}_{t-1}\leftarrow{\bm{x}}^{\prime}_{t-1}-\zeta_{t}\nabla_{{\bm{x}}_{t}}\|\bm{y}-\mathcal{A}({{\bm{x}}}_{0|t})\|^{2}_{2}$
-9:    ${\bm{x}}_{t-1}\leftarrow{\bm{x}}_{t-1}-\lambda_{t}\nabla_{{\bm{x}}_{t}}\|S_{g}(\mathcal{E}({{\bm{x}}}_{0|t}))-\mathcal{E}(T_{g}({{\bm{x}}}_{0|t}))\|_{2}^{2}$
+9:  ${\bm{x}}_{t-1}\leftarrow{\bm{x}}_{t-1}-\lambda_{t}\nabla_{{\bm{x}}_{t}}\mathcal{R}({\bm{x}}_{t})$
 10:end for
 11:return ${{\bm{x}}}_{0}$
-We use the above regularization for image restoration tasks (e.g., see Algorithm 1 for Equi-DPS).
-Data inherent equivariance MPE can also emerge naturally from symmetries underlying the data.
-This often occurs in physics systems where the combination of coefficient functions, boundary values, and solution functions of PDEs remains valid after applying an invertible coordinate transformation consistently to them.
-Let $\mathcal{G}(a)\mapsto u$ be a PDE operator that maps from the initial condition $a$ to the solution $u$.
-Then, it follows that $S_{g}(\mathcal{G}(a))=\mathcal{G}(T_{g}(a)).$
-Table 1: Effect of EquiReg on ReSample [37]. Inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.01$.
+where $\mathcal{R}({\bm{x}}_{t})$ is the regularizer.
+Applying this to DPS (Chung et al., 2023) takes the form in Algorithm˜1).
+This formulation brings us to the key contribution: how to design the regularizer to be effective in improving posterior sampling of diffusion models.
+Regularization in optimization aims to penalize undesired solutions, moving the algorithm towards solutions with low regularizer value.
+In the context of diffusion models for sampling from data distributions, we interpret an ideal regularizer as follows: it should yield low values for on-manifold and high values for off-manifold samples, enabling accurate posterior sampling even when the likelihood score is approximated.
+In terms of sampling dynamics, i.e., when applied at each reverse-diffusion step, the regularizer should effectively penalize trajectories leaving the data manifold and reinforce those aligned with high-probability regions.
+This motivates designing a regularizer that applies a global correction to the entire functional, in contrast to prior works that focus only on locally reducing likelihood error.
+The ideal property of a regularizer would be to produce high error on undesirable samples and low error on desirable samples.
+We use equivariance, a global property that enforces geometric symmetries to instantiate such a regularizer to guide the diffusion process toward the data manifold.
+To realize this idea, we seek functions that exhibit approximate equivariance and discriminate on- from off-manifold samples.
+We propose to quantify the equivariance of a function relative to a data distribution.
+Specifically, while the literature has primarily studied the equivariance properties of functions for general inputs, we propose a new definition for functions in which their equivariance error is distribution-dependent and defined under the support of an input data distribution (Definition˜3.1).
+Definition 3.1 (Distribution-Dependent Equivariant Functions).
+Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$.
+The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the distribution $p$ is defined as $\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|]$.
+The above definition enables us to define functions whose equivariance error can differentiate on-manifold samples from off-manifold ones.
+We aim to find functions whose equivariance error is low for on-manifold data and high elsewhere.
+We also introduce a constrained version of equivariance error, where the input is implicitly regularized to lie on the manifold $\mathcal{M}$ in addition to minimizing the equivariance error (Definition˜3.2).
+Both equivariance errors are non-local, defined at the distribution level.
+When used to regularize the reverse conditional diffusion process, they are computed via local evaluations over the sampled data.
+Definition 3.2 (Manifold-Constrained Distribution-Dependent Equivariant Functions).
+Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$.
+The manifold-constrained equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the data distribution $p$ is $\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|{\bm{z}}-h(S_{g}^{-1}(f(T_{g}({\bm{z}}))))\|]$ where $h:\mathcal{X}\rightarrow\mathcal{Z}$, and the pair $(f,h)$ forms a vanishing-error autoencoder (see Appendix I).
+To define our method, we term a class of manifold-preferential equivariant (MPE) functions, whose equivariance error is lower for samples on the data manifold than for off-manifold samples.
+EquiReg is a regularization framework, not a manifold projection method.
+EquiReg penalizes states that deviate from symmetry-preserving regions; when an MPE function is used, these regions align with the data manifold.
+In practice, MPE functions can emerge in different ways, which we illustrate with examples from augmented training and from data symmetries.
+MPE can emerge when functions are trained with symmetry-preserving mechanisms such as data augmentation.
+Prior work has studied equivariant properties of learned representations in deep networks (Lenc and Vedaldi, 2015), showing that data augmentations (Krizhevsky et al., 2012) and representation compression via reduced model capacity (Bruintjes et al., 2023) promote equivariant features even when equivariance is not explicitly built into the architecture.
+Importantly, the trained network is only approximately equivariant, and prior studies have noted that symmetry-preserving properties degrade for inputs deviating from in-distribution data (Azulay and Weiss, 2019).
+A few studies have leveraged this emergent MPE in trained networks for out-of-distribution detection (Zhou, 2022; Kaur et al., 2022; 2023).
+To demonstrate the widespread MPE property of learned mappings, we have considered additional pre-trained models and quantified their equivariance loss for several datasets, i.e., natural images and corrupted ones (see Appendix˜I of Appendix.) Figure˜4(a) illustrates the MPE property, emergent via training with augmentations, of $\mathcal{E}$-$\mathcal{D}$ of a pre-trained autoencoder, currently used in LDMs.
+Specifically, it shows that the equivariance error is lower for natural images and increases when images deviate from the clean data distribution.
+Based on Definitions˜3.1 and 3.2, we propose Equi loss using an MPE function $f$ for diffusion-based inverse solvers:
+$\mathcal{R}_{f_{\text{MPE}}}({\bm{x}}_{t})$ $=\|S_{g}(f_{\text{MPE}}({\bm{x}}_{0|t}))-f_{\text{MPE}}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$ (5)
+where ${\bm{x}}_{0|t}$ and ${\bm{z}}_{0|t}$ are function of ${\bm{x}}_{t}$ and ${\bm{z}}_{t}$, respectively.
+MPE can also emerge due to symmetries present in the data itself during training.
+This often occurs in physics systems where coefficient functions, boundary values, and solution functions of PDEs remain valid under invertible coordinate transformations.
+Formally, let $\mathcal{G}(a)\mapsto u$ be a PDE operator mapping initial condition $a$ to solution $u$, and let $T_{g}$ and $S_{g}$ be invertible transformations that preserve PDE structure and boundary conditions.
+Then, $S_{g}(\mathcal{G}(a))=\mathcal{G}(T_{g}(a))$.
+Neural operators (Kovachki et al., 2021), popular architectures for modelling physics, trained on PDEs with such inherent symmetries can learn equivariance properties.
+Figure˜4(b) shows that an MPE function can be constructed using Fourier Neural Operators (FNOs (Li et al., 2021)) trained on non-augmented Navier–Stokes physics data.
+When a pre-trained FNO is used as the MPE function, the equivariance loss in equation 5 is lower for in-distribution data than for out-of-distribution data under reflection as the group action.
+(a) Equivariance from augmentation, error computed with $f_{\text{MPE}}=\mathcal{E}$ from autoencoder pretrained on FFHQ-256 and evaluated on examples from FFHQ-256.
+(b) Equivariance arising from data symmetries. Error computed with $f_{\text{MPE}}=\text{FNO}$ and evaluated on Navier-Stokes data.
+Figure 4: Equivariance error is consistently lower for clean vs perturbed examples.
+The key message from our MPE examples is that MPE behaviour naturally emerges when a function (e.g., a neural network) is trained with appropriate augmentations or when the data exhibit inherent symmetries.
+Our results extend this analysis and demonstrate that MPE functions are widely present and can be effectively integrated into the EquiReg framework to improve posterior sampling (Table˜8).
+We leverage this property to distinguish on-manifold from off-manifold samples and to regularize the posterior sampling trajectory toward high-probability regions.
+Finally, we note that the choice of symmetry group may often be a challenge depending on application domain, a shared challenge in the broader equivariance literature.
+We provide guidelines on how to choose symmetry groups in Appendix˜H with reference on automatic symmetry discovery from data (Zhou et al., 2021; Quessard et al., 2020; Dehmamy et al., 2021; Mohapatra et al., 2025).
+Finally, we complement our empirical results with a theoretical discussion of the role of the function $\mathcal{R}$ in regularizing conditional diffusion models and its impact on the sampling trajectory.
+In particular, we gain insight into the desirable properties of an optimal regularizer by reinterpreting the reverse conditional diffusion process as a time-inhomogeneous Wasserstein gradient flow (Ferreira and Valencia-Guevara, 2018) (see Propositions˜G.1 and G.2 in Appendix).
+Table 1: EquiReq Characterization under change of EquiReg period, $\lambda_{t}$, and DDIM steps.
+Super Resolution Gaussian Blur
+[…]
+DPS N/A 46.20 22.99 (1.93) 0.20 (0.05) 135.7 46.50 24.59 (2.25) 0.15 (0.03) 88.70
+Equi-DPS 1 51.10 26.73 (1.99) 0.12 (0.03) 87.97 52.20 26.08 (2.25) 0.12 (0.03) 87.11
+Equi-DPS 2 48.90 26.73 (1.99) 0.12 (0.03) 87.98 49.10 26.06 (2.24) 0.12 (0.03) 87.19
+Equi-DPS 5 47.10 26.73 (1.99) 0.12 (0.03) 87.98 47.30 26.06 (2.24) 0.12 (0.03) 87.32
+Equi-DPS 10 46.90 26.73 (1.99) 0.12 (0.03) 87.99 47.00 26.05 (2.24) 0.12 (0.03) 87.04
+(a) Robustness and computational efficiency of applying EquiReg under various periods during sampling.
+EquiReg maintains performance when applied every $\{1,2,5,10\}$ DDIM steps while incurring minimal computational overhead.
+DPS PSLD
+[…]
+0.0 24.34 (1.03) 0.664 (0.061) 0.156 (0.051) 0.0 23.83 (2.61) 0.63 (0.12) 0.315 (0.07)
+0.001 25.44 (1.22) 0.708 (0.057) 0.118 (0.038) 0.01 25.35 (2.24) 0.70 (0.09) 0.280 (0.07)
+0.01 25.44 (1.22) 0.708 (0.057) 0.118 (0.038) 0.1 26.63 (1.68) 0.74 (0.08) 0.337 (0.06)
+0.1 25.44 (1.22) 0.708 (0.057) 0.118 (0.038) 0.25 26.22 (1.57) 0.72 (0.08) 0.366 (0.05)
+1.0 25.44 (1.22) 0.709 (0.057) 0.118 (0.038) 1.0 24.74 (1.28) 0.66 (0.07) 0.438 (0.05)
+(b) Robustness of EquiReg to the choice of $\lambda_{t}$. Sensitivity analysis for DPS and PSLD.
+DPS Equi-DPS (ours)
+[…]
+500 13.89 0.0937 0.955 417.07 20.61 0.366 0.500 238.51
+750 21.77 0.540 0.254 153.74 25.60 0.704 0.160 110.89
+900 22.97 0.628 0.201 148.03 26.52 0.755 0.126 88.46
+1000 22.99 0.649 0.201 135.71 26.73 0.767 0.120 88.00
+(c) EquiReg improves performance under reduced DDIM steps. Pixel-based super-resolution on FFHQ $256\times 256$.
+Table 2: EquiReg improves SITCOM under reduced measurement consistency steps ($K_{\text{meas}}$).
+We reduce $K_{\text{meas}}$ and add an equal amount of EquiReg steps ($K_{\text{EquiReg}}$).
+Evaluated with $50$ DDIM steps on motion deblur for FFHQ.
+[…]
+10 N/A 28.06 0.81 21.57
+5 5 29.26 0.83 11.09
+20 N/A 27.04 0.79 38.85
+10 10 28.93 0.82 20.92
+30 N/A 27.79 0.80 58.84
+15 15 29.63 0.84 30.19
+40 N/A 30.40 0.85 78.08
+20 20 29.50 0.83 41.02
+60 N/A 28.35 0.81 108.57
+30 30 31.36 0.87 59.38
+4 Results
+This section provides experimental results evaluating the performance of EquiReg on inverse problems, including linear and nonlinear image restorations and PDE solving.
+To fairly assess the impact of EquiReg, we adopt a paired comparison setting (e.g., PSLD vs.
+Equi-PSLD) across experiments, keeping all other factors such as architecture, training, and sampling fixed.
+This design ensures that any observed improvements can be attributed specifically to EquiReg rather than to differences in the underlying model or inference procedure.
+For fair comparison, we set a fixed seed for the initial noise and evaluate the impact of EquiReg under reduced measurement consistency and sampling steps, providing a path toward faster diffusion-based inversion.
+Results emphasize EquiReg’s usefulness when the baseline performance deteriorates.
+We analyze the diversity trade-offs, and demonstrate the broad emergence of MPEs and their usefulness within EquiReg.
+Lastly, we provide preliminary analysis on EquiReg improving the realism of text-guided image generation.
+Image restoration tasks.
+We evaluate EquiReg when applied to: SITCOM (Alkhouri et al., 2025), PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DPS (Chung et al., 2023).
+We compare against manifold-preserving or geometry-constraint approaches including MCG (Chung et al., 2022b), MPGD-AE (He et al., 2024), and DiffStateGrad (Zirvi et al., 2025).
+Table 3: EquiReg for diffusion models on FFHQ. $256\times 256$ with $\sigma_{{\bm{y}}}=0.05$.
+Method Gaussian deblur Motion deblur Super-resolution ($\times 4$) Box inpainting Random inpainting
+[…]
+PSLD 0.357 106.2 22.87 0.322 84.62 24.25 0.313 89.72 24.51 0.158 43.02 24.22 0.246 49.77 29.05
+Equi-PSLD 0.344 94.09 24.42 0.338 99.14 24.83 0.289 90.88 26.32 0.098 31.54 24.19 0.188 41.61 30.43
+EquiCon-PSLD 0.320 83.18 24.38 0.322 89.87 25.14 0.277 79.39 26.14 0.092 35.07 24.26 0.204 40.75 29.99
+(a) Latent diffusion.
+Method Gaussian deblur Motion deblur Super-resolution ($\times 4$) Box inpainting Random inpainting
+[…]
+DPS 0.145 104.8 25.48 0.132 99.75 26.75 0.191 125.4 24.38 0.133 56.89 23.10 0.113 51.32 29.63
+Equi-DPS (ours) 0.114 48.76 26.32 0.094 41.71 28.23 0.120 51.00 27.15 0.099 40.47 23.39 0.068 33.65 32.16
+DiffStateGrad-DPS 0.128 52.73 26.29 0.118 50.14 27.61 0.186 73.02 24.65 0.114 47.53 24.10 0.107 49.42 30.15
+MCG 0.340 101.2 6.72 0.702 310.5 6.72 0.520 87.64 20.05 0.309 40.11 19.97 0.286 29.26 21.57
+MPGD-AE 0.150 114.9 24.42 0.120 104.5 25.72 0.168 137.7 24.01 0.138 248.7 21.59 0.172 339.0 25.22
+(b) Pixel-based diffusion.
+Table 4: EquiReg for latent diffusion models on ImageNet. $256\times 256$ with $\sigma_{{\bm{y}}}=0.05$.
+Method Gaussian deblur Motion deblur Super-resolution (x4) Box inpainting Random inpainting
+[…]
+PSLD 263.9 20.70 252.1 21.26 224.3 22.29 151.4 16.28 83.22 26.56
+EquiCon-PSLD 214.5 22.01 196.3 22.69 198.5 22.34 137.6 19.25 65.14 27.03
+Table 5: EquiReg for ReSample on linear and nonlinear tasks. FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.01$.
 […]
@@
 EquiCon-ReSample 0.078 37.61 30.73 0.860
-Super-res. (×$\times 4$) ReSample 0.204 40.46 28.02 0.790
+Super-res. ($\times 4$) ReSample 0.204 40.46 28.02 0.790
 Equi-ReSample 0.098 43.56 29.74 0.849
@@
 EquiCon-ReSample 0.125 54.62 29.55 0.843
-We show that models trained on PDEs with inherent symmetries can indeed learn the equivariance property.
-Neural operators [130] are a popular architecture for modeling physics systems, as their function space setup can preserve the continuum of physics functions and generalize well to different resolutions.
-Among these, the Fourier Neural Operator (FNO [131]) has drawn the most attention.
-We train an FNO on the inverse problem of Navier-Stokes, which respects reflection symmetry (see details in Appendix), without explicit data augmentation.
-As shown in Figure 4(b), FNOs trained on non-augmented physics data yield lower equivariance error ($\text{Equi}_{\text{pixel}}$ loss: $\mathcal{R}({\bm{x}}_{t})=\|S_{g}(\text{FNO}({\bm{x}}_{0|t}))-\text{FNO}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$) on in-distribution as opposed to out-of-distribution data.
-5 Results
-This section provides experimental results on the performance of EquiReg for inverse problems, including linear and nonlinear image restoration tasks and reconstructing PDEs.
-To explore the impact of scaling the noise scheduler, the Appendix includes an experiment on DreamSampler [82], showing preliminary evidence that EquiReg can improve performance on text-to-image guidance.
-Experimental setup for image restorations.
-We evaluate the performance of EquiReg when applied to the following three diffusion-based inverse solvers: PSLD [38], ReSample [37], and DPS [32].
-Of these methods, PSLD and ReSample are latent solvers, while DPS is a pixel-based solver.
-We compare against several other methods including DDRM [34], MCG [33], DiffPIR [39], MPGD-AE [62], and DiffStateGrad [41].
 We measure performance via perceptual similarity (LPIPS), distribution alignment (FID), pixel-wise fidelity (PSNR), and structural consistency (SSIM).
-We test EquiReg on two datasets: a) the FFHQ $256\times 256$ validation dataset [132], and b) the ImageNet $256\times 256$ validation dataset [133].
-For pixel-based experiments, we use i) the pre-trained diffusion model from [32] on the FFHQ dataset, and ii) the pre-trained model from [68] on the ImageNet dataset.
-For latent diffusion experiments, we use i) the unconditional LDM-VQ-4 model [61] on FFHQ, and ii) the Stable Diffusion v1.5 [61] model on ImageNet.
-Table 2: Linear inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
-Method Gaussian deblur Motion deblur Super-resolution (×$\times 4$)
+We test EquiReg on a) the FFHQ $256\times 256$ (Karras et al., 2021) and b) ImageNet $256\times 256$ validation set (Deng et al., 2009).
+For pixel-based experiments, we use the pre-trained model from i) (Chung et al., 2023) on FFHQ, and ii) (Dhariwal and Nichol, 2021) on ImageNet.
+For latent diffusion experiments, we use i) the unconditional LDM-VQ-4 model (Rombach et al., 2022) on FFHQ, and ii) the Stable Diffusion v1.5 (Rombach et al., 2022) model on ImageNet.
+We evaluate EquiReg on linear and nonlinear restoration tasks for natural images (see Appendix˜E for task details).
+We adopt the pre-trained encoder-decoder $\mathcal{E}$-$\mathcal{D}$ as our MPE function.
+For FFHQ, we use vertical reflection as the symmetry group, which preserves upright facial orientation.
+For ImageNet, we define a rotation group $G=\{0,\nicefrac{{\pi}}{{2}},\pi,\nicefrac{{3\pi}}{{2}}\}$, and uniformly at random select the group action for each sample.
+Finally, the loss functions given in Equation˜5 are used to regularize.
+While our main experiment explore the reflection and rotation groups with small cardinality, EquiReg does not rely on full group coverage.
+Sampling even a sparse or randomly chosen subset of group actions is sufficient, as long as the function used for regularization exhibits the MPE property across the group (Table˜10).
+We characterize EquiReg under change of its period, regularization parameter $\lambda_{t}$, and DDIM steps.
+EquiReg can be applied efficiently with longer period’ it preserves performance when applied with lower frequency (Table˜1(c)a); and it is robust to the choice of its hyperparameter $\lambda_{t}$, demonstrated on DPS and PSLD (Table˜1(c)b).
+EquiReg maintains strong performance as the number of DDIM steps is reduced, whereas DPS suffers a significant drop.
+Importantly, Equi-DPS consistently outperforms DPS, with the performance gap widening at lower step counts (Table˜1(c)c, Figure˜12).
+An important advantage of EquiReg is that it allows the user to reduce the number of measurement consistency optimization steps, thereby introducing an implicit acceleration in solving the inverse problem.
+Table˜2 shows that EquiReg regularization enables SITCOM to achieve superior performance with significantly reduced runtime by using fewer measurement consistency steps.
+Table˜3a, Table˜4, and Table˜5 highlights the benefits of EquiReg for latent models by consistently improving the performance of ReSample and PSLD across several tasks on FFHQ and ImageNet.
+We attribute this improvement in part to the reduction of failure cases (Figure˜5d).
+EquiReg also significantly improves the performance of pixel-based methods (see Equi-DPS vs.
+DPS, Table˜3d).
+We observe that EquiReg achieves its largest improvements on perceptual metrics (FID and LPIPS), suggesting it generates more realistic images that lie closer to the data manifold (see Appendix E for supporting qualitative results).
+EquiReg improves performance under high measurement noise (Figure˜5b).
+This result aligns with Figure˜5a, which shows the equivariance error is lower on clean images than noisy ones, indicating that EquiReg enforces an effective denoising.
+Lastly, we note that EquiReg is robust to regularizing hyperparameter $\lambda_{t}$ (Figure˜5c, see Appendix˜C for details).
+For qualitative performance of EquiReg, see Figure˜15 and Figure˜16.
+Table 6: Solving PDEs from sparse observations.
+Steps $(N)$ Helmholtz Navier-Stokes
+Forward Inverse Forward Inverse
+DiffusionPDE 2000 12.64% 19.07% 3.78% 9.63%
+FunDPS 500 2.13% 17.16% 3.32% 8.48%
+Equi-FunDPS (ours) 500 2.12% 15.91% 3.06% 7.84%
+Solving PDEs from sparse observations.
+EquiReg is evaluated on two important PDE problems: the Helmholtz and Navier-Stokes equations (see Appendix˜F).
+The objective is to solve both forward and inverse problems in sparse sensor settings.
+The forward problem involves predicting the solution function or the final state using measurements from only $3\%$ of the coefficient field or the initial state.
+The inverse problem, conversely, aims to predict the input conditions from observations of $3\%$ of the system’s output.
+This task is challenging due to the nonlinearity of the equations, the complex structure of Gaussian random fields, and the sparsity of observations.
+Recent studies (Huang et al., 2024; Mammadov et al., 2024a; Yao et al., 2025) have demonstrated the superiority of diffusion models over deterministic single-forward methods for solving PDEs.
+DiffusionPDE (Huang et al., 2024) decomposes the conditional log-likelihood into a learned diffusion prior and a measurement score.
+FunDPS (Yao et al., 2025) extends the sampling process to a more natural infinite-dimensional spaces, achieving better accuracy and speed via function space models.
+We integrate EquiReg into the state-of-the-art FunDPS framework (Mammadov et al., 2024a; Yao et al., 2025), where the equivariance loss is computed using an FNO trained on the corresponding inverse problem.
+We employ reflection symmetry (i.e., flipping along the $y=x$ axis) and observe no significant performance differences when using alternative transformations such as rotations or alternating flips.
+Equi-FunDPS improves performance (Table˜6), measured by relative $\ell_{2}$ error, across multiple tasks, particularly in inverse problems where a strong data prior is essential.
+Importantly, these results highlight that EquiReg is not limited to conventional neural networks and is naturally applicable to neural operators, enabling regularization in both finite-dimensional and function-space diffusion models.
+Diversity analysis.
+To study posterior sampling diversity of EquiReg, we generated $K=10$ posterior samples for $20$ test images across three inverse problems of box inpainting, Gaussian deblurring, $4\times$ super-resolution, and measured diversity using two complementary metrics: Intra-LPIPS for perceptual diversity and Pixel-Std for spatial diversity.
+Table˜7 demonstrates that Equi-DPS achieves favorable fidelity-diversity trade-offs.
+We further investigated diversity scaling by varying box inpainting mask size from $128\times 128$ to $192\times 192$ pixels (Figure˜17).
+Results show that diversity metrics increase linearly with task difficulty, demonstrating that Equi-DPS naturally expands sampling as problems become more ill-posed rather than artificially constraining solutions.
+This linear relationship indicates healthy, predictable posterior sampling behavior across the difficulty spectrum.
+Lastly, Figures˜6 and 18 provides qualitative confirmation through visual examples showing four posterior samples per image.
+Observable variations in facial features, expressions, and eye gaze validate our quantitative measurements, confirming EquiReg can generate genuinely diverse reconstructions rather than collapsing to a single solution (for additional discussion and analysis, see Appendix˜D).
+Figure 5: EquiReg effectiveness and robustness across a range of measurement noise levels and regularization parameter (a) Equivariance error computed over a pre-trained decoder on increasingly noisy inputs.
+(b) EquiReg performance computed over a range of measurement noise levels on the FFHQ dataset.
+(c) EquiReg reduces failure cases and enhances reconstruction fidelity for super-resolution on FFHQ.
+Figure 6: Qualitative diversity examples for super-resolution.
+We show $K=4$ posterior samples for two test images.
+Samples differ in facial features (i.e., teeth in the first test image, eye color and eyelashes in the second test image) while maintaining high fidelity to the ground truth, demonstrating EquiReg generates diverse plausible reconstructions rather than collapsing to a single mode.
+MPE functions are widespread and effective for EquiReg.
+We empirically evaluate whether MPE behaviour naturally emerges across commonly used neural networks.
+We examine i) the emergence of MPE properties in different functions (neural networks) and ii) the effect of using these functions within EquiReg on identical inverse problem settings.
+For FFHQ and ImageNet, we analyze four function classes: the LDM encoder (Rombach et al., 2022), a CNN autoencoder trained with symmetry augmentations (flip for FFHQ and rotation for ImageNet), pre-trained ResNet-50 (He et al., 2016), and CLIP (Radford et al., 2021).
+Across all architectures, equivariance error increases systematically as Gaussian noise pushes samples off the data manifold, confirming clear MPE behaviour.
+The strength of this effect varies: the CNN autoencoder exhibits the strongest MPE signal, while the LDM encoder shows the weakest; ResNet-50 and CLIP lie between these extremes (see Figures˜14 and 19).
+Notably, the strongest MPE behaviour of the CNN autoencoder is in line with our systematic guidelines for constructing MPE functions; This is precisely the regime where the training distribution of the function matches the distribution of the inverse problem (i.e., training on ImageNet train and evaluating on ImageNet test).
+We then integrate each function into EquiReg under identical inverse problem settings (FFHQ/ImageNet, DPS/SITCOM, super-resolution/motion deblurring).
+In all cases, EquiReg improves reconstruction quality over the corresponding baseline without regularization (Tables˜8 and 13).
+These results demonstrate that i) MPE behaviour naturally arises in widely used pre-trained networks, and ii) EquiReg is robust to the choice of MPE function, even when the MPE signal is relatively weak.
+Notably, our main experiments use the weakest MPE (the LDM encoder), suggesting further gains are possible with stronger MPE constructions.
+Text-to-image guidance.
+Given a “source” image, DreamSampler (Kim et al., 2024) transforms it according to a text prompt.
+When applying EquiReg to DreamSampler, perceptual quality and reduced artifacts are improved in the generated images.
+Figure˜2 shows a source “cat” transformed into a “corgi”: Equi-DreamSampler produces more realistic results and resolves anatomical inconsistencies (e.g., correcting a three-front-legged corgi to two front legs).
+We also observe an implicit acceleration effect when EquiReg is imposed (Figure˜7).
+Equi-DreamSampler with $50$ DDIM steps produces images comparable to DreamSampler with substantially more steps.
+Increasing the regularization strength $\lambda_{t}$ at fixed $50$ DDIM steps results in effects similar to increasing DDIM steps in DreamSampler (from $50$ to $75$ to $100$), suggesting that EquiReg guides sampling trajectories closer to the data manifold.
+These experiments illustrate that EquiReg is readily applicable beyond inverse problems, including text-to-image guidance.
+While not our primary focus, they demonstrate the broader utility and flexibility of EquiReg as a plug-and-play regularization framework.
+Table 7: Fidelity and diversity comparison across inverse problems. EquiReg improves fidelity while largely preserving or enhancing sampling diversity (2020 test images with $K=10$ samples per image.)
+Task Method Fidelity Metrics Diversity Metrics
 […]
-Latent
-PSLD 0.357 106.2 22.87 0.537 0.322 84.62 24.25 0.615 0.313 89.72 24.51 0.650
-Equi-PSLD (ours) 0.344 94.09 24.42 0.657 0.338 99.14 24.83 0.669 0.289 90.88 26.32 0.718
-EquiCon-PSLD (ours) 0.320 83.18 24.38 0.664 0.322 89.87 25.14 0.699 0.277 79.39 26.14 0.722
-Pixel-based
-DDRM 0.239 92.43 24.93 0.732 - - - - 0.282 79.25 26.58 0.782
-MCG 0.340 101.2 6.72 0.051 0.702 310.5 6.72 0.055 0.520 87.64 20.05 0.559
-DiffPIR 0.236 59.65 27.36 - 0.255 65.78 26.57 - 0.260 65.77 26.64 -
-MPGD-AE 0.150 114.9 24.42 0.664 0.120 104.5 25.72 0.723 0.168 137.7 24.01 0.670
-DPS 0.145 104.8 25.48 0.709 0.132 99.75 26.75 0.754 0.191 125.4 24.38 0.675
-DiffStateGrad-DPS 0.128 52.73 26.29 0.739 0.118 50.14 27.61 0.782 0.186 73.02 24.65 0.683
-Equi-DPS (ours) 0.114 48.76 26.32 0.750 0.094 41.71 28.23 0.810 0.120 51.00 27.15 0.776
-(a) Deblurring and super-resolution linear inverse problems.
-Method Box inpainting Random inpainting
+Box inpainting DPS 0.140 70.89 0.112 9.286
+Equi-DPS (ours) 0.112 59.70 0.118 10.59
+Gaussian deblur DPS 0.150 76.71 0.114 6.565
+Equi-DPS (ours) 0.120 63.02 0.092 5.669
+Super-resolution ($\times 4$) DPS 0.683 99.11 0.134 7.956
+Equi-DPS (ours) 0.703 87.52 0.187 23.52
+Table 8: MPE behaviour emerges across diverse network architectures and consistently improves diffusion-based inverse problems with EquiReg.
+SITCOM motion deblurring on FFHQ 256 with $\lambda=0.05$.
+Results are reported as mean (standard deviation).
+MPE function PSNR SSIM LPIPS
+None 27.670 (1.343) 0.790 (0.031) 0.221 (0.040)
+LDM Encoder (FFHQ) 28.357 (1.379) 0.806 (0.031) 0.200 (0.036)
+CNN Autoencoder (FFHQ) 28.852 (1.376) 0.819 (0.044) 0.193 (0.033)
+Pretrained ResNet50 28.682 (1.388) 0.811 (0.036) 0.198 (0.036)
+5 Conclusion
+We introduce Equivariance Regularized (EquiReg) diffusion for inverse problems.
+EquiReg regularizes sampling trajectories to stay closer to the data manifold, leveraging manifold-preferential equivariance (MPE): functions with low equivariance error on-manifold and high error off-manifold.
+Such functions arise naturally in trained networks and can serve as plug-and-play regularizers without modifying the diffusion denoiser.
+EquiReg is agnostic across pixel- and latent-space diffusion models and remains robust under reduced sampling, effectively accelerating convergence.
+Across diverse inverse problems, it consistently improves perceptual and reconstruction metrics while reducing failure cases, highlighting its generality and efficiency.
+EquiReg operates as a plug-and-play regularization framework and therefore builds upon the quality of the underlying diffusion solver.
+It does not modify the diffusion architecture itself, but instead improves sampling by guiding trajectories toward more consistent, on-manifold solutions.
+As a regularization mechanism, its impact is most pronounced in challenging regimes where likelihood guidance alone may degrade or become unstable.
+Applying EquiReg requires selecting appropriate symmetry groups and constructing suitable MPE functions for the task at hand.
+While we provide systematic guidelines for imaging and PDE settings, extending these constructions to new domains remains an important direction for future work; we have provided reference to prior work on how symmetries for various application can be learned and set.
+Finally, although we formalize distribution-dependent equivariant and MPE functions, a deeper theoretical understanding of when MPE behaviour emerges in trained networks and how it may interact with joint training of diffusion models, remains an interesting avenue for further study.
+Broader Impact Statement.
+On the positive side, high‑fidelity image restoration can improve downstream tasks in medical imaging, remote‑sensing and environmental monitoring (e.g., denoising satellite observations to track pollution or deforestation).
+Likewise, accelerated PDE‑solving via learned diffusion priors may enable faster, more accurate simulations for climate modeling, fluid‑dynamics research, and engineering design.
+On the other hand, robust reconstruction methods could be misappropriated for privacy‑invasive surveillance or to create deceptive imagery.
+We emphasize that our method does not amplify these existing risks.
+Appendices for “EquiReg: Equivariance Regularized Diffusion for Inverse Problems”
+We provide our source code when EquiReg. We will provide a publicly available source code upon acceptance. This supplementary materials contain the following:
+  • Appendix˜A includes additional experiments on text-to-image guidance. We regularize DreamSampler (Kim et al., 2024) with EquiReg for an improved performance (see Figures˜7, 8, 9, 10 and 11).
+  • Appendix˜B includes additional experiments on robustness including robustness to $\lambda_{t}$, reduced number of DDIM steps, and reduced number of measurent consistency steps.
+  • Appendix˜C includes qualitative analysis on the performance of methods with and without EquiReg.
+Results show a reduction of artifacts and an improved perceptual quality of the solution.
+This section also includes the equivariance error of a pre-trained encoder used in EquiReg (Figure˜14(a)) and a histogram of Equi’s improvement on DPS (Figure˜13).
+  • Appendix˜D includes diversity experiments. Results show that EquiReg achieves favorable fidelity-diversity tradeoffs (Table˜7, Figure˜17, and Figure˜18).
+  • Appendix˜E demonstrates EquiReg experimental setup and implementation for PSLD, ReSample, and DPS (Algorithms˜2, 3, 4, 5 and 6).
+It also contains information about the EquiReg hyperparameters for image restoration tasks.
+  • Appendix˜F contains information on the PDE reconstruction experiment. It discusses the equations along with implementation details and hyperparameters.
+  • Appendix˜G provides theoretical proofs of Propositions˜G.1 and G.2.
+  • Appendix˜H contains additional background information on solving inverse problems, vanishing-error autoencoders, and equivariance.
+  • Appendix˜I provides additional figures on MPE functions.
+  • Appendix˜J discloses computing resources used to conduct the experiments.
+  • Appendix˜K credits code assets used for our experiments.
+  • Appendix˜L concludes the appendix with a “responsible release” statement.
+The authors acknowledge the usage of LLMs on proofreading and improving the coherency of the manuscript. The authors have not used LLMs for content generation.
+Appendix A EquiReg for Text-to-Image Guidance
+Figure 7: Impact of EquiReg parameter $\lambda_{t}$, implicit acceleration, and introduction of more image details on FFHQ $\times512$. Women, red hair, smiling.
+Figure 8: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance on FFHQ $\times512$. Boy, curly hair.
+Figure 9: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Glasses.
+Figure 10: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Mustache, beard.
+Figure 11: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Sculpture.
+Appendix B Additional Experiments on Robustness
+Figure 12: Advantages of EquiReg under reduced DDIM steps. Super-resolution on FFHQ.
+Table 9: EquiReg improves SITCOM under reduced measurement consistency steps ($K_{\text{meas}}$). Motion deblur on FFHQ sampled with $50$ DDIM steps.
 […]
-Latent
-PSLD 0.158 43.02 24.22 0.819 0.246 49.77 29.05 0.809
-Equi-PSLD (ours) 0.098 31.54 24.19 0.875 0.188 41.61 30.43 0.878
-EquiCon-PSLD (ours) 0.092 35.07 24.26 0.884 0.204 40.75 29.99 0.847
-Pixel-based
-DDRM 0.207 78.62 22.26 0.801 0.218 - 25.75 -
-MCG 0.309 40.11 19.97 0.703 0.286 29.26 21.57 0.751
-MPGD-AE 0.138 248.7 21.59 0.753 0.172 339.0 25.22 0.732
-DPS 0.133 56.89 23.10 0.799 0.113 51.32 29.63 0.849
-DiffStateGrad-DPS 0.114 47.53 24.10 0.831 0.107 49.42 30.15 0.852
-Equi-DPS (ours) 0.099 40.47 23.39 0.835 0.068 33.65 32.16 0.902
-(b) Inpainting linear inverse problems.
-Figure 5: EquiReg consistently improves performance across a range of measurement noise levels. Gaussian deblur using FFHQ $256\times 256$ for $\sigma_{{\bm{y}}}=[0,0.8]$.
+10 N/A 28.06 0.81 21.57
+10 1 28.71 0.82 21.07
+5 5 29.26 0.83 11.09
+20 N/A 27.04 0.79 38.85
+20 1 28.54 0.82 37.74
+10 10 28.93 0.82 20.92
+30 N/A 27.79 0.80 58.84
+30 1 28.35 0.81 55.51
+15 15 29.63 0.84 30.19
+40 N/A 30.40 0.85 78.08
+40 1 30.58 0.85 69.83
+20 20 29.50 0.83 41.02
+60 N/A 28.35 0.81 108.57
+60 1 27.02 0.78 95.62
+30 30 31.36 0.87 59.38
+Table 10: EquiReg Effectiveness with Subset of Group Actions.
+PSLD Equi-PSLD (90, 270 deg)
+$uparrow PSNR\uparrow SSIM\uparrow$
+15.86 (1.19) 0.77 (0.03) 17.60 (1.60) 0.79 (0.03)
+Appendix C Visualizations for Image Restoration Experiments
+Figure 13: Histogram of EquiReg improvement for DPS. Super-resolution using FFHQ $256\times 256$.
+(a) The equivariance error of the encoder is lower on clean, natural images than corrupted ones.
+(b) Example visualizations of used images and corresponding equivariance error computed using the decoder (see Figure˜4(a)).
+Figure 14: Training induced equivariance for a pre-trained function.
+(a) Box inpainting.
+(b) Random inpainting.
+(c) Gaussian deblur.
+(d) Super-resolution ($\times 4$).
+Figure 15: Qualitative comparison of EquiCon-PSLD and PSLD on FFHQ 256 $\times$ 256.
+(a) Random inpainting.
+(b) Super-resolution ($\times 4$).
+Figure 16: Qualitative comparison of Equi-DPS and DPS on ImageNet 256 $\times$ 256.
+Appendix D Diversity Analysis
+In the Bayesian setting, the objective of solving inverse problems with diffusion models is to sample from high-probability regions of the posterior distribution.
+While the goal is not to maximize “diversity”, the true diversity emerges when the posterior admits meaningful variability.
+In practice, diversity-related concerns in inverse problems arise when a method suffers from mode collapse, i.e., the sampler becomes biased and fails to explore multiple plausible modes of the posterior.
+Thus, the relevant question is whether a method properly explores the posterior rather than whether it maximizes diversity in an unconstrained sense.
+Because closed-form posteriors are unavailable for real image restoration tasks, the standard practice in the diffusion inverse-problem literature is to evaluate diversity through variation among plausible reconstructions consistent with the measurement, without collapsing to a single solution.
+This is the notion of “diversity” our work adopts.
+Given the goal of posterior sampling, EquiReg is not designed to maximize diversity for its own sake.
+Its objective is to incorporate data-inherent geometric structure (equivariance) to guide sampling toward high-probability regions of the posterior.
+Hence, diversity arises naturally from the ill-posedness of the inverse problem; it is a consequence of posterior uncertainty, not the goal of the regularizer.
+To quantify this effect, in addition to reconstruction quality, we analyzed the diversity of posterior samples produced by EquiReg.
+We evaluate diversity metrics across multiple tasks and difficulty levels to characterize the sampling behavior of our method.
+D.1 Experimental Setup
+[…]
+D.2 Results and Discussion
+Table 7 shows that Equi-DPS achieves favorable fidelity-diversity trade-offs across three inverse problems.
+For box inpainting and super-resolution, equivariance regularization improves both fidelity and diversity simultaneously.
+For Gaussian deblurring, Equi-DPS achieves 15-20% better fidelity while retaining 80-85% of baseline diversity, representing a modest but justified trade-off.
+These results demonstrate that equivariance constraints do not inherently suppress diversity; rather, they can guide sampling toward regions of higher data fidelity while maintaining posterior exploration.
+Figure 17 reveals linear diversity scaling with task difficulty.
+Diversity metrics grow proportionally with task difficulty, indicating Equi-DPS naturally expands sampling as problems become more ill-posed.
+This linear relationship demonstrates stable, predictable behavior across difficulty levels without artificial diversity suppression.
+Figures 6 and 18 provide qualitative results.
+(a) Box inpainting
+(b) Gaussian deblur
+(c) Super-resolution
+Figure 17: Diversity vs task difficulty across three inverse problems.
+As task difficulty increases (larger inpainting mask, stronger blur, higher SR scale), both diversity metrics increase proportionally, demonstrating that Equi-DPS maintains healthy posterior sampling behavior across a wide difficulty spectrum.
+(a) Gaussian deblur
+(b) Box inpainting
+Figure 18: Qualitative diversity examples across Gaussian deblur and box inpainting.
+Each subfigure shows $K=4$ posterior samples for two different test images.
+(a) Gaussian deblur: samples differ in facial expressions and accessories (i.e., earrings in first test image).
+(b) Box inpainting: Inputs were obstructed with $160\times 160$ masks.
+Samples exhibit perceptually distinct facial features (i.e., expressions, eye gaze, facial structure).
+Across both tasks, EquiReg produces diverse plausible reconstructions rather than collapsing to a single mode.
+D.3 Conclusion
+Finally, we highlight that EquiReg improves both fidelity and diversity on 2 of the 3 considered tasks, an encouraging outcome that is uncommon given the general behavior of classical regularizers.
+Hand-crafted regularizers such as TV and $\ell_{1}$ may suppress diversity by shrinking solutions toward simple structures.
+By contrast, EquiReg leverages data-dependent regularization that captures the richness and structural complexity of the underlying data manifold, enabling it to preserve manifold-consistent variability while suppressing implausible samples.
+High diversity without fidelity is not meaningful for posterior sampling.
+A method that samples the entire solution space, including low-probability and artifacted regions, may score well on diversity but fail to provide useful reconstructions.
+Equi-DPS avoids this failure mode: it maintains meaningful diversity while reducing artifacts and improving perceptual quality.
+In the experiments conducted during the rebuttal, our goal was to demonstrate clearly that EquiReg preserves meaningful diversity, reflecting the posterior uncertainty, rather than unstructured or unconstrained variability.
+Appendix E Implementation Details for Image Restoration Tasks
+Experimental Setup.
 We evaluate EquiReg on a variety of linear and nonlinear restoration tasks for natural images.
@@
 For nonlinear inverse problems, we consider (1) phase retrieval, (2) nonlinear deblur, and (3) high dynamic range (HDR).
-We use an oversampling rate of $2.0$ for phase retrieval, and due to instability of the task, we generate four independent reconstructions and take the best result (as also done in DPS [32], DAPS [40], and DiffStateGrad [41]).
-We use the default setting from [134] for nonlinear deblur, and a scale factor of 2 for HDR.
-Table 3: Linear inverse problems on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
-Method Gaussian deblur Motion deblur Super-resolution (x4) Box inpainting Random inpainting
-[…]
-PSLD 263.9 20.70 252.1 21.26 224.3 22.29 151.4 16.28 83.22 26.56
-EquiCon-PSLD (ours) 214.5 22.01 196.3 22.69 198.5 22.34 137.6 19.25 65.14 27.03
-Figure 6: EquiReg reduces failure cases and enhances reconstruction fidelity. Super-resolution using FFHQ $256\times 256$.
-For these tasks, we adopt the pre-trained encoder-decoder $\mathcal{E}$-$\mathcal{D}$ as our MPE function.
-For FFHQ, the symmetry group contains reflection (across vertical axis).
-For ImageNet, we define a rotation group $G=\{\nicefrac{{\pi}}{{2}},\nicefrac{{3\pi}}{{2}}\}$, and uniformly at random select the group action for each sample.
-Finally, the loss functions given in Equation 6 are used as regularization terms.
-Image restoration performance.
-Our results (Tables 1, 2 and 3) show that EquiReg improves the performance of several methods across a wide variety of linear and nonlinear tasks on both FFHQ and ImageNet.
-EquiReg significantly improves performance on box inpainting when applied to ReSample (Table 1).
-We attribute this improvement in part to the reduction of failure cases (Figure 6).
-We observe that EquiReg achieves its largest improvements on perceptual metrics (FID and LPIPS), suggesting that it generates more realistic images that lie closer to the data manifold (see Appendix for supporting qualitative results).
-Moreover, EquiReg improves performance under high measurement noise (Figure 5).
-This result aligns with Figure 4(a), which shows the equivariance error is lower on clean images than noisy ones, indicating that EquiReg enforces an effective denoising.
-PDE reconstruction tasks.
-We evaluate the performance of EquiReg on two important PDE problems: the Helmholtz and Navier-Stokes equations (see Appendix).
-The objective is to solve both forward and inverse problems in sparse sensor settings.
-The forward problem involves predicting the solution function or the final state using measurements from only $3\%$ of the coefficient field or the initial state.
-The inverse problem, conversely, aims to predict the input conditions from observations of $3\%$ of the system’s output.
-This task is challenging due to the nonlinearity of the equations, the complex structure of Gaussian random fields, and the sparsity of observations.
-Table 4: Partially-observed PDE reconstruction.
-Steps $(N)$ Helmholtz Navier-Stokes
-Forward Inverse Forward Inverse
-DiffusionPDE 2000 12.64% 19.07% 3.78% 9.63%
-FunDPS 500 2.13% 17.16% 3.32% 8.48%
-Equi-FunDPS (ours) 500 2.12% 15.91% 3.06% 7.84%
-Recent studies [51, 56, 57] have demonstrated the superiority of diffusion-based approaches over deterministic single-forward methods for PDE reconstructions.
-DiffusionPDE [51] decomposes the conditional log-likelihood into a data prior (from a pre-trained diffusion model) and a measurement score, thereby providing solutions via conditional sampling.
-FunDPS [56, 57] further extends the sampling process to a more natural infinite-dimensional spaces, achieving better accuracy and speed via function space models.
-We integrate EquiReg into the state-of-the-art FunDPS framework [56, 57], computing the Equi loss using data inherent equivariance derived from an FNO trained on the corresponding inverse problem.
-For our experiments, we use reflection symmetry (i.e., flipping along the $y=x$ axis), and observe no significant performance difference when using other transformations such as rotations or alternating flips.
-As shown in Table 4, Equi-FunDPS improves performance, measured by relative $\ell_{2}$ loss—across various tasks, especially in inverse problems where a strong data prior is critical.
-6 Conclusion
-We introduce Equivariance Regularized (EquiReg) diffusion for solving inverse problems.
-EquiReg corrects errors introduced in posterior approximations by regularizing the sampling trajectories so they lie closer to the data manifold.
-Our method applies to many inverse problem settings across diverse domains and is flexible to the choice of diffusion model and sampling method.
-Limitations and future work.
-EquiReg’s effectiveness relies on the quality of the pre-trained diffusion prior.
-In addition, we do not address the approximations of the underlying diffusion models, but we regularize them for improved performance.
-Moreover, applying EquiReg requires task-specific regularizer design: one must choose an appropriate symmetry group and identify suitable MPE functions for each problem setting.
-As a result of this task-specific design, EquiReg can be explored in many other problem settings where equivariance is training induced or data inherent.
-We leave this for future work.
-7 Acknowledgements and Contributions
-Aditi worked on this project as part of her bachelor’s thesis. Anima was supported by Bren Chair Professor and AI2050 Senior Fellow.
-Bahareh proposed and led the project.
-Aditi and Rayhan implemented the method and executed all the image restoration experiments.
-Abbas conducted the text-to-image experiments, and Jiachen performed the PDE reconstruction experiments.
-Bahareh wrote the original draft of the paper with Aditi, Rayhan, and Jiachen writing the results section.
-The project was originally based on training-induced MPE functions, proposed by Bahareh.
-Later versions of the project include data-inherent MPE functions, proposed by Jiachen; this is used for the PDE experiments.
-Bahareh, Abbas, Jiachen, and Chuwei collaborated on formalizing the theory, and the final formulation of Wasserstein gradient flow was proposed and proven by Chuwei.
-All authors provided editorial comments and feedback on the manuscript.
-Anima supervised the project and provided valuable feedback.
-Appendices
-We provide our source code when EquiReg is implemented on DPS. This supplementary materials contain the following:
-  • Appendix A includes additional experiments on text-to-image guidance. We regularize DreamSampler [82] with EquiReg for an improved performance (see Figures 7, 8, 9 and 10).
-  • Appendix B includes additional image restoration experiments for ImageNet. We show that EquiReg improves LatentDAPS and DPS across several different inverse problems (Table 5, Table 6).
-  • Appendix C reports standard deviations on experiments done on PSLD (see Table 7).
-  • Appendix D includes qualitative analysis on the performance of methods with and without EquiReg.
-Results show a reduction of artifacts and an improved perceptual quality of the solution.
-This section also includes the equivariance error of a pre-trained encoder used in EquiReg (Figure 11(a)), an equivariant loss vs.
-Gaussian noise plot (Figure 12), and a histogram of Equi’s improvement on DPS (Figure 13).
-  • Appendix E demonstrates EquiReg implementation for PSLD, ReSample, DPS, and LatentDAPS (Algorithm 2, Algorithm 3, Algorithm 4, Algorithm 5, Algorithm 6, Algorithm 7).
-It also contains information about the EquiReg hyperparameters for image restoration tasks.
-  • Appendix F contains information on the PDE reconstruction experiment. It discusses the equations along with implementation details and hyperparameters.
-  • Appendix G provides theoretical proof of Propositions 4.1 and 4.2.
-  • Appendix H contains additional background information on solving inverse problems, vanishing-error autoencoders, and equivariance.
-  • Appendix I discloses computing resources used to conduct the experiments.
-  • Appendix J credits code assets used for our experiments.
-  • Appendix K discusses the broader impacts of this paper, the developed method, the conducted experiments, and their overall implications.
-  • Appendix L concludes the appendix with a “responsible release” statement.
-Appendix A EquiReg for Text-to-Image Guidance
-Given the “source” image, DreamSampler [82] is asked to transform the source image using the prompt.
-Applying EquiReg to DreamSampler, we observe perceptual improvement of generated images as well as artifact reduction.
-Figure 7 shows a cat, as the source image, being transformed into “corgi”, “croissant”, “cat statue”, and “fluffy dog”.
-Equi-DreamSampler generates more realistic images than DreamSampler (e.g., EquiReg modifies the three front legged corgi to two front legs).
-Figure 7: EquiReg helps the text-guided DreamSampler generate more realistic images.
-Figure 8: Impact of EquiReg parameter $\lambda$, implicit acceleration, and introduction of more image details. Women, red hair, smiling.
-We have also observed an implicit acceleration of image generation when EquiReg is imposed (Figure 8).
-Equi-DreamSampler with $50$ DDIM steps can generate images that are only possible with DreamSampler when the DDIM steps are increased.
-We attribute this to EquiReg’s ability to generate images that are closer to the data manifold.
-For example, the increase of DDIM steps in DreamSampler (from $50$ to $75$ to $100$) has a relatively similar effect to increasing the EquiReg regularizer $\lambda$ at a fixed $50$ DDIM steps.
-Figure 8 shows that increasing the regularization $\lambda$ results in addition of a car in the background.
-For DreamSampler, an early notion of the car seat in the background start to arise only when DDIM is increased to $100$ (see also Figures 9 and 10).
-Figure 9: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Boy, curly hair.
-Figure 10: Adding EquiReg into the text-to-image guidance method DreamSampler for improved performance. Glasses.
-Appendix B Additional Experiments on Image Restoration Tasks
-We show two additional results.
-EquiReg improves LatentDAPS performance on ImageNet for box inpainting and motion deblur (Table 5).
-Moreover, Equi-DPS performs much better than DPS, increasing the PSNR from $20.70$ to $23.45$ dB for super-resolution on ImageNet (Table 6).
-Table 5: Equi-LatentDAPS vs. LatentDAPS on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
-Method Box inpainting Motion deblur
-[…]
-LatentDAPS 0.310 20.19 0.305 25.08
-Equi-LatentDAPS (ours) 0.299 20.67 0.286 25.86
-Table 6: Equi-DPS vs. DPS on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
-Method Super-resolution (x4) Random inpainting
-[…]
-DPS 169.6 20.70 94.83 25.96
-Equi-DPS (ours) 102.7 23.45 55.87 28.32
-Appendix C Additional Information on Image Restoration Tasks
-We report standard deviation on all linear tasks for PSLD, Equi-PSLD, and EquiCon-PSLD (Table 7).
-Table 7: Reporting standard deviations. Inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{y}=0.05$.
-[…]
-Gaussian deblur PSLD 0.357 (0.060) 22.87 (1.67) 0.537 (0.094)
-Equi-PSLD (ours) 0.344 (0.055) 24.42 (1.66) 0.657 (0.076)
-EquiCon-PSLD (ours) 0.320 (0.048) 24.38 (1.55) 0.664 (0.068)
-Motion deblur PSLD 0.322 (0.048) 24.25 (1.47) 0.615 (0.075)
-Equi-PSLD (ours) 0.338 (0.052) 24.83 (1.75) 0.669 (0.068)
-EquiCon-PSLD (ours) 0.322 (0.054) 25.14 (1.62) 0.699 (0.064)
-Super-res. (×$\times 4$) PSLD 0.313 (0.098) 24.51 (2.98) 0.650 (0.140)
-Equi-PSLD (ours) 0.289 (0.088) 26.32 (2.24) 0.718 (0.104)
-EquiCon-PSLD (ours) 0.277 (0.093) 26.14 (2.74) 0.722 (0.119)
-Box inpainting PSLD 0.158 (0.023) 24.22 (2.80) 0.819 (0.031)
-Equi-PSLD (ours) 0.098 (0.019) 24.19 (2.67) 0.875 (0.027)
-EquiCon-PSLD (ours) 0.092 (0.021) 24.26 (2.80) 0.884 (0.029)
-Random inpainting PSLD 0.246 (0.053) 29.05 (1.98) 0.809 (0.049)
-Equi-PSLD (ours) 0.188 (0.034) 30.43 (1.76) 0.878 (0.027)
-EquiCon-PSLD (ours) 0.204 (0.051) 29.99 (2.06) 0.847 (0.047)
-Appendix D Visualizations for Image Restoration Experiments
-(a) The equivariance error of the encoder is lower on clean, natural images than corrupted ones.
-(b) Example visualizations of used images.
-Figure 11: Training induced equivariance for a pre-trained encoder. Experimental setup corresponds to Figure 4(a) in the main text, which uses a pre-trained decoder.
-Figure 12: Equivariance error of a pre-trained decoder over different Gaussian noise levels $\sigma$.
-We apply different Gaussian noise levels to clean images to see the effect of noise on equivariance error.
-We observe that as Gaussian noise increases, equivariance error increases steadily.
-Figure 13: Histogram of EquiReg improvement for DPS. Super-resolution using FFHQ $256\times 256$.
-(a) Box inpainting.
-(b) Random inpainting.
-(c) Gaussian deblur.
-(d) Super-resolution (×$\times 4$).
-Figure 14: Qualitative comparison of EquiCon-PSLD and PSLD on FFHQ 256 $\times$ 256.
-(a) Random inpainting.
-(b) Super-resolution (×$\times 4$).
-Figure 15: Qualitative comparison of Equi-DPS and DPS on ImageNet 256 $\times$ 256.
-Appendix E Implementation Details for Image Restoration Tasks
+We use an oversampling rate of $2.0$ for phase retrieval, and due to instability of the task, we generate four independent reconstructions and take the best result (as also done in DPS (Chung et al., 2023), DAPS (Zhang et al., 2025a), and DiffStateGrad (Zirvi et al., 2025)).
+We use the default setting from (Tran et al., 2021) for nonlinear deblur, and a scale factor of 2 for HDR.
 Hyperparameters.
-Our method introduces a single hyperparameter $\lambda$ that controls the amount of regularization applied.
-Below we include a table detailing the use of this hyperparameter in the main experiments (Table 8).
-For majority of experiments, we keep $\lambda$ constant throughout iterations.
-However, we note that in our Equi-DPS and Equi-LatentDAPS experiments, we scale $\lambda$ by the noise scheduler at every iteration.
-For all unscaled experiments, we employ early stopping, setting $\lambda=0$ for the last 10% of sampling.
-Table 8: Equivariance regularization weight $\lambda$ used in main experiments.
-Method Box Inpainting Random Inpainting Gaussian Deblur Motion Deblur Super-resolution (×$\times 4$)
+Our method introduces a single hyperparameter $\lambda_{t}$ that controls the amount of regularization applied.
+Below we include a table detailing the use of this hyperparameter in the main experiments (Table˜11).
+For majority of experiments, we keep $\lambda_{t}$ constant throughout iterations.
+For all unscaled experiments, we employ early stopping, setting $\lambda_{t}=0$ for the last 10% of sampling.
+Table 11: Equivariance regularization weight $\lambda_{t}$ used in main experiments.
+Method Box Inpainting Random Inpainting Gaussian Deblur Motion Deblur Super-resolution ($\times 4$)
 FFHQ $256\times 256$
@@
 EquiCon-PSLD 0.0015 0.05 0.06 0.07 0.001
-PSLD. We integrate EquiReg into PSLD by simply adding an additional gradient update step using our regularization term (Algorithm 2, Algorithm 3).
-In our experiments, we use the official PSLD implementation from Rout et al.
-[38], running with its default settings to reproduce the baseline results.
-We note that in our code, we do not square the norm when computing the gradient, aligning with PSLD’s implementation.
 Algorithm 2 Equi-PSLD for Image Restoration Tasks
-1:$T,\bm{y},\{\eta_{t}\}_{t=1}^{T},\{\gamma_{t}\}_{t=1}^{T},\{\tilde{\sigma}_{t}\}_{t=1}^{T}$
-2:$\mathcal{E},\mathcal{D},\mathcal{A}{\bm{x}}_{0}^{\ast},\mathcal{A},\bm{s}_{\theta},$ $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
-3:${\bm{z}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
-4:for $t=T-1$ to 00 do
+[…]
+[…]
+[…]
+4:for t=T1t=T-1 to 0 do
 5:  $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{z}}_{t},t)$
-6:    ${{\bm{z}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{z}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
+[…]
 7:  $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
@@
 13:return $\mathcal{D}({{\bm{z}}}_{0|t})$
+PSLD. We integrate EquiReg into PSLD by simply adding an additional gradient update step using our regularization term (Algorithms˜2 and 3).
+In our experiments, we use the official PSLD implementation from Rout et al.
+(2023), running with its default settings to reproduce the baseline results.
+We note that in our code, we do not square the norm when computing the gradient, aligning with PSLD’s implementation.
 Algorithm 3 EquiCon-PSLD for Image Restoration Tasks
-1:$T,\bm{y},\{\eta_{t}\}_{t=1}^{T},\{\gamma_{t}\}_{t=1}^{T},\{\tilde{\sigma}_{t}\}_{t=1}^{T}$
-2:$\mathcal{E},\mathcal{D},\mathcal{A}{\bm{x}}_{0}^{\ast},\mathcal{A},\bm{s}_{\theta},$ $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
-3:${\bm{z}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
-4:for $t=T-1$ to 00 do
+[…]
+[…]
+[…]
+4:for t=T1t=T-1 to 0 do
 5:  $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{z}}_{t},t)$
-6:    ${{\bm{z}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{z}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
+[…]
 7:  $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
@@
 ReSample.
-We integrate EquiReg into ReSample by adding our regularization term into the hard data consistency step (Algorithm 4, Algorithm 5).
+We integrate EquiReg into ReSample by adding our regularization term into the hard data consistency step (Algorithms˜4 and 5).
 We note that the ReSample algorithm employs a two-stage approach; initially, it performs pixel-space optimization, and later it performs latent-space optimization.
 We apply EquiReg in the latent-space optimization stage.
-In our experiments, we use the official ReSample implementation from Song et al. [37], running with its default settings to reproduce the baseline results.
+In our experiments, we use the official ReSample implementation from Song et al. (2023a), running with its default settings to reproduce the baseline results.
 Algorithm 4 Equi-ReSample for Image Restoration Tasks
 1:Measurements $\bm{y}$}, $\mathcal{A}(\cdot)$, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function ${\bm{s}}_{\theta}(\cdot,t)$, Pretrained LDM Parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\gamma$ to control $\sigma_{t}^{2}$, Time steps to perform resample $C$, $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
-2:${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$triangleright Initial noise vector
+[…]
 3:for $t=T-1,\ldots,0$ do
 4:  $\bm{\epsilon}_{1}\sim\mathcal{N}(\mathbf{0},\bm{I})$
-5:    $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$triangleright Compute the score
-6:    $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$triangleright Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
-7:    ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$triangleright Unconditional DDIM step
+5:  $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$ $\triangleright$ Compute the score
+6:  $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$ $\triangleright$ Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
+7:  ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$ $\triangleright$ Unconditional DDIM step
 8:  if $t\in C$ then $\triangleright$ ReSample time step
@@
 13:    end for
-14:         ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$triangleright Map back to $t$
+14:    ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$ $\triangleright$ Map back to $t$
 15:  else
-16:         ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$triangleright Unconditional sampling if not resampling
+16:    ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$ $\triangleright$ Unconditional sampling if not resampling
 17:  end if
 18:end for
-19:${\bm{x}}_{0}=\mathcal{D}({\bm{z}}_{0})$triangleright Output reconstructed image
+[…]
 20:return ${\bm{x}}_{0}$
@@
 1:Measurements $\bm{y}$}, $\mathcal{A}(\cdot)$, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function ${\bm{s}}_{\theta}(\cdot,t)$, Pretrained LDM Parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\gamma$ to control $\sigma_{t}^{2}$, Time steps to perform resample $C$, $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
-2:${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\bm{I})$triangleright Initial noise vector
+[…]
 3:for $t=T-1,\ldots,0$ do
 4:  $\bm{\epsilon}_{1}\sim\mathcal{N}(\mathbf{0},\bm{I})$
-5:    $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$triangleright Compute the score
-6:    $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$triangleright Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
-7:    ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$triangleright Unconditional DDIM step
+5:  $\hat{\bm{\epsilon}}_{t+1}={\bm{s}}_{\theta}({\bm{z}}_{t+1},t+1)$ $\triangleright$ Compute the score
+6:  $\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}({\bm{z}}_{t+1}-\sqrt{1-\bar{\alpha}_{t+1}}\hat{\bm{\epsilon}}_{t+1})$ $\triangleright$ Predict $\hat{{\bm{z}}}_{0}$ using Tweedie’s formula
+7:  ${\bm{z}}^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0}({\bm{z}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}+\eta\delta\bm{\epsilon}_{1}$ $\triangleright$ Unconditional DDIM step
 8:  if $t\in C$ then $\triangleright$ ReSample time step
@@
 13:    end for
-14:         ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$triangleright Map back to $t$
+14:    ${\bm{z}}_{t}=\text{StochasticResample}(\hat{{\bm{z}}}_{0}({\bm{y}}),{\bm{z}}^{\prime}_{t},\gamma)$ $\triangleright$ Map back to $t$
 15:  else
-16:         ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$triangleright Unconditional sampling if not resampling
+16:    ${\bm{z}}_{t}={\bm{z}}^{\prime}_{t}$ $\triangleright$ Unconditional sampling if not resampling
 17:  end if
 18:end for
-19:${\bm{x}}_{0}=\mathcal{D}({\bm{z}}_{0})$triangleright Output reconstructed image
+[…]
 20:return ${\bm{x}}_{0}$
-DPS. Similar to PSLD, we integrate EquiReg into DPS by simply adding an additional gradient update step using our regularization term (Algorithm 6).
-In our experiments, we use the official DPS implementation from Chung et al. [32], running with its default settings to reproduce the baseline results.
+DPS. Similar to PSLD, we integrate EquiReg into DPS by simply adding an additional gradient update step using our regularization term (Algorithm˜6).
+In our experiments, we use the official DPS implementation from Chung et al. (2023), running with its default settings to reproduce the baseline results.
 Algorithm 6 Equi-DPS for Image Restoration Tasks
 […]
-2:${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
-3:for $t=T-1$ to 00 do
+[…]
+3:for t=T1t=T-1 to 0 do
 4:  $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
-5:    ${{\bm{x}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{x}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
+[…]
 6:  $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
@@
 11:return ${{\bm{x}}}_{0}$
-LatentDAPS. We integrate EquiReg into LatentDAPS by simply adding an additional gradient update step using our regularization term (Algorithm 7).
-In our experiments, we use the official LatentDAPS implementation from Zhang et al. [40], running with its default settings to reproduce the baseline results.
-Algorithm 7 Equi-LatentDAPS for Image Restoration Tasks
-1:Latent space score model ${\bm{s}}_{\theta}$, measurement ${\bm{y}}$, noise schedule $\sigma_{t}$, $\{t_{i}\}_{i\in\{0,\ldots,N_{A}\}}$, encoder $\mathcal{E}$ and decoder $\mathcal{D}$,  $T_{g}$ and $S_{g}$, $\{\lambda_{t}\}_{t=1}^{T}$
-2:Sample ${\bm{z}}_{T}\sim\mathcal{N}(\mathbf{0},\sigma_{T}^{2}\bm{I})$
-3:for $i=N_{A},N_{A}-1,\ldots,1$ do
-4:    Compute $\hat{{\bm{z}}}_{0}^{(0)}=\hat{{\bm{z}}}_{0}({\bm{z}}_{t_{i}})$ by solving the probability flow ODE with ${\bm{s}}_{\theta}$
-5:    for $j=0,\ldots,N-1$ do
-6:          $\bm{g}\leftarrow\nabla_{\hat{{\bm{z}}}_{0}}\log p(\hat{{\bm{z}}}_{0}^{(j)}|{\bm{z}}_{t_{i}})+\nabla_{\hat{{\bm{z}}}_{0}}\log p({\bm{y}}|\hat{{\bm{z}}}_{0}^{(j)})+\lambda_{t}\nabla_{\hat{{\bm{z}}}_{0}}\|S_{g}(\mathcal{D}(\hat{{\bm{z}}}_{0}^{(j)}))-\mathcal{D}(T_{g}(\hat{{\bm{z}}}_{0}^{(j)}))\|_{2}^{2}$
-7:         $\hat{{\bm{z}}}_{0}^{(j+1)}\leftarrow\hat{{\bm{z}}}_{0}^{(j)}+\eta_{t}\bm{g}+\sqrt{2\eta_{t}}\bm{\epsilon}_{j}$, $\bm{\epsilon}_{j}\sim\mathcal{N}(\mathbf{0},\bm{I})$
-8:    end for
-9:    Sample ${\bm{z}}_{t_{i-1}}\sim\mathcal{N}(\hat{{\bm{z}}}_{0}^{(N)},\sigma_{t_{i-1}}^{2}\bm{I})$
-10:end for
-11:return $\mathcal{D}({\bm{z}}_{0})$
+SITCOM.
+We augment the original SITCOM algorithm by introducing an additional equivariant refinement stage at each reverse diffusion step.
+After completing the standard measurement and backward-consistency gradient updates, we perform a second optimization over the equivariance loss, enforcing consistency between $\mathcal{E}(T_{g}(v))$ and $T_{g}(\mathcal{E}(v))$ (Algorithm˜7).
+In our experiments, we use the official SITCOM implementation from Alkhouri et al. (2025), running with its default settings to reproduce the baseline results.
+Algorithm 7 Equi-SITCOM for Image Restoration Tasks
+1:Measurements $\mathbf{y}$, forward operator $\mathcal{A}(\cdot)$, pre-trained DM $\epsilon_{\theta}(\cdot,\cdot)$, diffusion steps NN, schedule $\bar{\alpha}_{i}$, measurement gradient steps $K$, equivariant gradient steps $K_{\text{equi}}$, stop $\delta$, lr $\gamma$, reg.
+$\lambda$.
+2:Restored image $\hat{\mathbf{x}}$.
+3:Initialize $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\Delta t=\left\lfloor\tfrac{T}{N}\right\rfloor$.
+4:for $i=N,N-1,\ldots,1$ do $\triangleright$ Reducing diffusion sampling steps
+5:  $\mathbf{v}_{i}^{(0)}\leftarrow\mathbf{x}_{i}$ $\triangleright$ Init for closeness (C3)
+6:  for $k=1,\ldots,K$ do $\triangleright$ Adam on measurement/backward consistency (C1, C2)
+7:    $\mathbf{v}_{i}^{(k)}\leftarrow\mathbf{v}_{i}^{(k-1)}-\gamma\nabla_{\mathbf{v}_{i}}\!\left[\left\|\mathcal{A}\!\left(\tfrac{1}{\sqrt{\bar{\alpha}_{i}}}\big(\mathbf{v}_{i}-\sqrt{1-\bar{\alpha}_{i}}\,\epsilon_{\theta}(\mathbf{v}_{i},i\Delta t)\big)\right)-\mathbf{y}\right\|_{2}^{2}+\lambda\|\mathbf{x}_{i}-\mathbf{v}_{i}\|_{2}^{2}\right]\Big|_{\mathbf{v}_{i}=\mathbf{v}_{i}^{(k-1)}}$
+8:    if $\left\|\mathcal{A}\!\left(\tfrac{1}{\sqrt{\bar{\alpha}_{i}}}\big(\mathbf{v}_{i}^{(k)}-\sqrt{1-\bar{\alpha}_{i}}\,\epsilon_{\theta}(\mathbf{v}_{i}^{(k)},i\Delta t)\big)\right)-\mathbf{y}\right\|_{2}^{2}<\delta^{2}$ then
+9:     break $\triangleright$ Prevent noise overfitting
+10:    end if
+11:  end for
+12:  $\mathbf{v}_{i}^{(0)}\leftarrow\mathbf{v}_{i}^{(k)}$ $\triangleright$ Initialize to optimized $\mathbf{v}_{i}$
+13:  for $k=1,\ldots,K_{\text{equi}}$ do
+14:    $\mathbf{v}_{i}^{(k)}\leftarrow\mathbf{v}_{i}^{(k-1)}-\gamma\nabla_{\mathbf{v}_{i}}\!\left[\left\|\mathcal{E}\!\big(T_{g}(\mathbf{v}_{i}^{(k)})\big)-T_{g}\!\big(\mathcal{E}(\mathbf{v}_{i}^{(k)})\big)\right\|_{2}^{2}\right]\Big|_{\mathbf{v}_{i}=\mathbf{v}_{i}^{(k-1)}}$
+15:    if $\left\|\mathcal{E}\!\big(T_{g}(\mathbf{v}_{i}^{(k)})\big)-T_{g}\!\big(\mathcal{E}(\mathbf{v}_{i}^{(k)})\big)\right\|_{2}^{2}<\delta^{2}$ then
+16:     break
+17:    end if
+18:  end for
+19:  $\hat{\mathbf{v}}_{i}\leftarrow\mathbf{v}_{i}^{(k)}$ $\triangleright$ Backward diffusion consistency (C2)
+20:  $\hat{\mathbf{x}}^{\prime}_{0}\leftarrow\tfrac{1}{\sqrt{\bar{\alpha}_{i}}}\!\left[\hat{\mathbf{v}}_{i}-\sqrt{1-\bar{\alpha}_{i}}\,\epsilon_{\theta}(\hat{\mathbf{v}}_{i},i\Delta t)\right]$ $\triangleright$ Backward consistency (C2)
+21:  $\mathbf{x}_{i-1}\leftarrow\sqrt{\bar{\alpha}_{i-1}}\,\hat{\mathbf{x}}^{\prime}_{0}+\sqrt{1-\bar{\alpha}_{i-1}}\,\bm{\eta}_{i}$, $\bm{\eta}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ $\triangleright$ Forward consistency (C3)
+22:end for
+23:return $\hat{\mathbf{x}}=\mathbf{x}_{0}$
 Appendix F Experiment Setup for PDE Reconstructions
@@
 We note that this system has reflection equivariance along $x_{1}=\frac{1}{2},x_{2}=\frac{1}{2},x_{1}=x_{2}$ and rotation equivariance by $\frac{\pi}{2},\pi,\frac{3\pi}{2}$.
-Navier-Stokes equations. Following the methodology of [135], we model the time evolution of a vorticity field, $u(x,t)$, governed by:
+Navier-Stokes equations. Following the methodology of (Li et al., 2020), we model the time evolution of a vorticity field, $u(x,t)$, governed by:
 $\partial_{t}u(x,t)+{\bm{w}}(x,t)\cdot\nabla u(x,t)$ $=\nu\Delta u(x,t)+f(x),\quad x\in(0,1)^{2},\,t\in(0,T],$ (7)
@@
 The forcing term is $f(x)=0.1\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$.
-We borrow the dataset from [51].
+We borrow the dataset from (Huang et al., 2024).
 We note that this system has a reflection symmetry along the $x_{1}=x_{2}$ axis.
@@
 EquiReg, as a regularizer for diffusion posterior sampling, can be adapted to many inverse solvers in a plug-and-play manner.
-For PDE experiments, we use the same model weights and configurations as FunDPS [57].
+For PDE experiments, we use the same model weights and configurations as FunDPS (Yao et al., 2025).
 Error rates are calculated using the $L^{2}$ relative error between the predicted and true solutions, averaged on 100 randomly selected test samples.
-We provide the information on the EquiReg scaling weights in Table 9.
+We provide the information on the EquiReg scaling weights in Table˜12.
 Table 12: EquiReg loss used in PDE experiments.
@@
 Appendix G Theoretical Analysis
-In this section, we provide the proofs for Proposition 4.1 and Proposition 4.2.
-G.1 Preliminary and Notations
+G.1 Summary of the Analysis
+The theoretical framework presented in this paper is intended to motivate and guide the design of effective regularizers.
+This perspective, grounded in optimal transport theory (Ferreira and Valencia-Guevara, 2018), serves as an intuitive interpretation of the dynamics and motivates the design of regularization strategies such as EquiReg.
+We note that whether diffusion models follow exact Wasserstein dynamics still remains an open problem (Zheng et al., 2025).
+Proposition G.1.
+Let $\rho({\bm{x}},t)$ be the distribution of ${\bm{x}}_{T-t}$ driven by the ideal reverse dynamics (eq.˜3).
+Then, the evolution of $\rho$ follows the Wasserstein-2 gradient flow associated with minimizing functional $\Phi(\rho,t)$ defined as $\beta_{T-t}\int[\rho\phi({\bm{x}},t)+\frac{1}{2}\rho\log\rho]d{\bm{x}}$, where $\phi({\bm{x}},t)=-(\log p_{T-t}({\bm{x}}|{\bm{y}})+\frac{1}{4}\|{\bm{x}}\|^{2})$.
+The dynamics of $\rho$ remain the same if we replace $\phi({\bm{x}},t)$ with $\phi_{C}({\bm{x}},t):=\phi({\bm{x}},t)-C(t)$ for arbitrary temporal function $C(t)$.
+Without loss of generality, we assume $\phi_{C}({\bm{x}},t)<0$ for all $\bm{x}$} and $t$.
+In practice, the density function $p_{T-t}$ is not available and thus $\phi_{C}({\bm{x}},t)$ is approximated as $\hat{\phi}$ with $p_{T-t}({\bm{x}}_{T-t}|{\bm{y}})\approx\tilde{C}p_{T-t}({\bm{x}}_{T-t})p({\bm{y}}|\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}])$ where $\tilde{C}$ only depends on $\bm{y}$}.
+Because the conditional expectation $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}]$ is a linear combination of all candidate ${\bm{x}}_{0}$, the approximation remains relatively accurate when TtT-t is small (i.e., ${\bm{x}}_{T-t}$ stays close to the data manifold under low noise) but may incur high error for larger TtT-t, as shown in Figure˜1b.
+To mitigate this, we reweight the contributions to the first term of $\Phi$, down-weighting unreliable estimates, and amplifying the reliable ones.
+The resulting reweighted functional is
+[…]
+where $Z_{t}={\int{e^{\frac{\mathcal{R}({\bm{x}})}{\hat{\phi}_{c}({\bm{x}},t)}}\mathrm{d}{\bm{x}}}}$ is the normalizing factor, and $\mathcal{R}({\bm{x}})$ is a positive regularization that is nearly zero near the data manifold and much larger elsewhere.
+Intuitively, since $\hat{\phi}_{C}<0$, the weight is nearly one for $\bm{x}$} near the data manifold and much smaller elsewhere.
+Proposition G.2.
+(Informal) The evolution of $\rho$, the probability distribution of ${\bm{x}}_{T-t}$ driven by the practical and regularized reverse dynamics (eq.˜11), is an approximation of the Wasserstein-2 gradient flow associated with minimizing $\tilde{\Phi}$.
+[…]
+G.2 Preliminary and Notations
 We first remind the readers of gradient flow under the Wasserstein-2 metric and introduce the notations related to the diffusion model.
@@
 Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ be a functional of probability distributions.
-The Wasserstein gradient flow of $\mathcal{F}$ is characterized by the minimizing movement scheme (also known as JKO scheme) introduced by [136].
+The Wasserstein gradient flow of $\mathcal{F}$ is characterized by the minimizing movement scheme (also known as JKO scheme) introduced by (Jordan et al., 1998).
 For a fixed time step $\tau>0$, the sequence $(\rho_{k})_{k\in\mathbb{N}}$ of probability densities is defined recursively by:
@@
 where $\bm{w}_{t}$ is standard Brownian motion, $\beta_{t}>0$ is a time-dependent variance schedule, and $f({\bm{x}},t)$ is a drift term.
-For instance, $f\equiv 0$ for a variance-exploding SDE and $f({\bm{x}},t)=-\frac{\beta_{t}}{2}{\bm{x}}$ for a variance-preserving SDE defined in [30].
+For instance, $f\equiv 0$ for a variance-exploding SDE and $f({\bm{x}},t)=-\frac{\beta_{t}}{2}{\bm{x}}$ for a variance-preserving SDE defined in (Song et al., 2021).
 In this work, we carry out our analysis under a more general setting.
@@
 This form describes the generative process as evolving forward from $t=0$ to $t=T$, matching the usual direction of analysis in gradient flow frameworks.
-G.2 Proof of Proposition 4.1
+G.3 Proof of Proposition˜G.1
 In this work, we consider Wasserstein gradient flow under the setting where the functional $\mathcal{F}$ depends on time.
@@
 where $T^{*}({\bm{x}})$ is the optimal transport map, and $v^{*}({\bm{x}})$ is the associated optimal velocity field.
-Thus, we can rewrite the eq. 16 as
+Thus, we can rewrite the eq.˜17 as
 $\inf_{v}\ \mathcal{F}(\rho,t)$ $-\Delta t\int\nabla\cdot(\rho({\bm{x}})v({\bm{x}}))\,\frac{\delta\mathcal{F}(\rho,t)}{\delta\rho}({\bm{x}})\,\mathrm{d}{\bm{x}}+\Delta t\int\big[\rho({\bm{x}})\partial_{t}V({\bm{x}},t)+\dot{\alpha}(t)\rho\log\rho\big]\mathrm{d}{\bm{x}}$ (20)
@@
 $v({\bm{x}},t)=-\nabla\frac{\delta\mathcal{F}(\rho,t)}{\delta\rho}({\bm{x}})=-(\nabla V({\bm{x}},t)+\alpha(t)\nabla\log\rho({\bm{x}},t)).$ (24)
-The corresponding evolution of probability density is
+We note that By Hörmander’s theorem, a smooth density $\rho({\bm{x}},t)$ exists for $t>0$, ensuring that the above $v$ is well-defined. The corresponding evolution of probability density is
 $\partial_{t}\rho({\bm{x}},t)$ $=-\nabla\cdot(\rho({\bm{x}},t)v({\bm{x}},t))$ (25)
 $=\nabla\cdot(\rho({\bm{x}},t)(\nabla V({\bm{x}},t)+\alpha(t)\frac{\nabla\rho({\bm{x}},t)}{\rho}))$ (26)
-$=-\nabla\cdot(\rho({\bm{x}},t)(-\nabla V({\bm{x}},t))+\alpha(t)\Delta\rho({\bm{x}},t),$ (26)
+$=-\nabla\cdot(\rho({\bm{x}},t)(-\nabla V({\bm{x}},t))+\alpha(t)\Delta\rho({\bm{x}},t)),$ (27)
 which is exactly the Fokker-Planck equation describing the evolution of the probability density describing the particles following
 $\mathrm{d}{\bm{x}}_{t}=-\nabla V({\bm{x}}_{t},t)\mathrm{d}t+{\sqrt{2\alpha(t)}}\mathrm{d}{\bm{w}}_{t}.$ (28)
-Now we come back to Proposition 4.1. From eq. 14 we know that choosing
+Now we come back to Proposition˜G.1. From eq.˜15 we know that choosing
 $V({\bm{x}},t)=h({\bm{x}},T-t)-\beta_{T-t}[\log p_{T-t}({\bm{x}})+\log p_{T-t}({\bm{y}}|{\bm{x}})]\ \ \text{and}\ \alpha(t)=\frac{\beta_{T-t}}{2}$ (29)
-in Lemma G.1 completes the proof, where $h$ is defined in Assumption G.1.
-G.3 Detailed version of Proposition 4.2
+in Lemma˜G.1 completes the proof, where $h$ is defined in Assumption˜G.1.
+G.4 Detailed version of Proposition˜G.2
 In practice, one does not have access to $\log p_{t}({\bm{y}}|{\bm{x}}_{t})$ which appears in the reverse SDE. The most popular approach is do the following approximation,
@@
 As discussed in the main text, since the conditional expectation is a linear combination over all possible values of ${\bm{x}}_{0}$, it may fall outside the data manifold, resulting in physically invalid samples.
-One of the central challenges in diffusion-based inverse sampling is guiding the sampling trajectory—generated by the reverse SDE dynamics—toward the data manifold.
+One of the central challenges in diffusion-based inverse sampling is guiding the sampling trajectory, generated by the reverse SDE dynamics, toward the data manifold.
 A common strategy is to incorporate regularization into the reverse SDE to encourage manifold adherence.
 In this work, building on the perspective of Wasserstein gradient flow as outlined above, we provide a novel interpretation of the role played by such regularization terms.
-We show that the regularizer serves to reweight the contribution of different regions in the calculation of the underlying functional being minimized, $\Phi(\rho,t)$ defined in Proposition 4.1.
-Specifically, it amplifies the influence of regions where the density estimate is reliable (typically near the data manifold), while down-weighting regions with poor approximation quality of based on eq.
-29, often corresponding to off-manifold samples.
-[…]
+We show that the regularizer serves to reweight the contribution of different regions in the calculation of the underlying functional being minimized, $\Phi(\rho,t)$ defined in Proposition˜G.1.
+Specifically, it amplifies the influence of regions where the density estimate is reliable (typically near the data manifold), while down-weighting regions with poor approximation quality of based on eq.˜30, often corresponding to off-manifold samples.
+Following from what we have shown in the main text, $\Phi(\rho,t)$ has the form of $\beta_{T-t}\int[\rho\phi({\bm{x}},t)+\frac{1}{2}\rho\log\rho]d{\bm{x}}$ for a function $\phi({\bm{x}},t)$, which can be derived by (29).
+The $\log p_{t}({\bm{y}}|{\bm{x}})$ term in (29) or $\nabla\log p_{t}({\bm{y}}|{\bm{x}})$ term in (28), equivalently, is computed based on approximation (30).
+We denote the corresponding approximation of $\phi({\bm{x}},t)$ as $\hat{\phi}({\bm{x}},t)$.
+As discussed in the main text, we can assume without loss of generality that $\phi({\bm{x}},t)<0$ and $\hat{\phi}({\bm{x}},t)<0$.
+We have
 $\hat{\Phi}(\rho,t)=\beta_{T-t}\big[\int_{{\bm{x}}\in N(\mathcal{M})}\rho({\bm{x}})\hat{\phi}({\bm{x}},t)\mathrm{d}{\bm{x}}+\int_{{\bm{x}}\notin N(\mathcal{M})}\rho({\bm{x}})\hat{\phi}({\bm{x}},t)\mathrm{d}{\bm{x}}+\frac{1}{2}\int\rho\log\rho\mathrm{d}{\bm{x}}\big],$ (31)
@@
 Note that $\hat{\phi}<0$, we have that
-$\mathcal{R}({\bm{x}})\approx\begin{cases}0,&{\bm{x}}\in N(\mathcal{M})\\ \gg 1,&{\bm{x}}\text{ far away from }N(\mathcal{M})\end{cases}\quad\Rightarrow\quad\lambda({\bm{x}},t)\approx\begin{cases}1,&{\bm{x}}\in N(\mathcal{M})\\ 0,&{\bm{x}}\text{ far away from }N(\mathcal{M})\end{cases}.$
+[…]
 Next, we consider practical algorithms based on this reweighted functional.
@@
 Thus, the Wasserstein gradient flow associated with (34) is intractable since we cannot evaluate the weight function.
-We consider the following approximation based on $e^{\delta}\approx 1+\delta$,
+We consider the following approximation based on $e^{\delta}\approx 1+\delta$ when $\delta$ is sufficiently small,
 […]
 […]
-By Lemma G.1, the dynamics of ${\bm{x}}$ driven by the Wasserstein gradient flow associated with the approximated functional above is
+By Lemma˜G.1, the dynamics of $\bm{x}$} driven by the Wasserstein gradient flow associated with the approximated functional above is
 $\mathrm{d}{\bm{x}}=[-f({\bm{x}},T-t)-\beta_{T-t}\nabla_{{\bm{x}}}\big(\log{p_{T-t}({\bm{x}})}+\log\hat{p}_{T-t}({\bm{y}}|{\bm{x}})+\mathcal{R}({\bm{x}})\big)]\mathrm{d}t+\sqrt{\beta_{T-t}}\mathrm{d}\bm{\bar{w}}.$ (37)
@@
 Solving inverse problems with deep learning prior to diffusion models.
-Earlier works [137, 138, 139, 140] used deep neural networks as denoisers to solve inverse problems.
-Furthermore, deep generative models such as variational autoencoders (VAEs) [141], and generative adversarial networks (GANs) [142] were employed.
-Notable applications include compressed sensing [143] and MRI [144].
-Resources for Definition 4.2 on vanishing-error autoencoders.
-Manifold constrained distribution-dependent equivariance error uses the notion of vanishing-error autoencoders [145, 146, 62] (Definition H.1), also known as an asymptotically-trained autoencoder [146] or a perfect autoencoder [62].
-Vanishing-error autoencoders have previously been employed by diffusion-based inverse solvers to preserve the diffusion process on the manifold [62].
+Earlier works (Metzler et al., 2016; Romano et al., 2017; Zhang et al., 2017; Metzler et al., 2017) used deep neural networks as denoisers to solve inverse problems.
+Furthermore, deep generative models such as variational autoencoders (VAEs) (Kingma, 2013), and generative adversarial networks (GANs) (Goodfellow et al., 2014) were employed.
+Notable applications include compressed sensing (Bora et al., 2017) and MRI (Jalal et al., 2021).
+Applications on diffusion models to solve inverse problems.
+Most popular applications include image restoration (Chung et al., 2023; 2022b; Kawar et al., 2022; Lugmayr et al., 2022; Saharia et al., 2022; Song et al., 2023a; Rout et al., 2023; Zhu et al., 2023; Zhang et al., 2025a; Zirvi et al., 2025), medical imaging (Song et al., 2022; Chung and Ye, 2022; Chung et al., 2022a; Hung et al., 2023; Dorjsembe et al., 2024; Li et al., 2024; Kazerouni et al., 2023; Bian et al., 2024), and solving partial differential equations (PDEs) (Isakov, 2006; Huang et al., 2024; Shysheya et al., 2024; Liu et al., 2023; Li et al., 2025; Baldassari et al., 2023; Mammadov et al., 2024a; Yao et al., 2025).
+On the methodology side, there has been numerous advancements (Chung et al., 2023; 2022b; Kawar et al., 2022; Lugmayr et al., 2022; Saharia et al., 2022; Song et al., 2023a; Rout et al., 2023; Zhu et al., 2023; Zhang et al., 2025a; Zirvi et al., 2025; Song et al., 2022; Chung and Ye, 2022; Chung et al., 2022a; Hung et al., 2023; Dorjsembe et al., 2024; Li et al., 2024; Kazerouni et al., 2023; Bian et al., 2024; Huang et al., 2024; Shysheya et al., 2024; Mammadov et al., 2024b; Cardoso et al., 2024).
+Resources for Definition˜3.2 on vanishing-error autoencoders.
+Manifold constrained distribution-dependent equivariance error uses the notion of vanishing-error autoencoders (Shao et al., 2018; Anders et al., 2020; He et al., 2024) (Definition˜H.1), also known as an asymptotically-trained autoencoder (Anders et al., 2020) or a perfect autoencoder (He et al., 2024).
+Vanishing-error autoencoders have previously been employed by diffusion-based inverse solvers to preserve the diffusion process on the manifold (He et al., 2024).
 Definition H.1 (Vanishing-Error Autoencoder).
 A vanishing-error autoencoder under the manifold $\mathcal{M}$ with encoder $\mathcal{E}:\mathcal{X}\rightarrow\mathcal{Z}$ and decoder $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{X}$ with $\mathcal{Z}=\mathbb{R}^{k}$ where $k<d$, has zero reconstruction error under the support of the data distribution $\mathcal{X}$, i.e., $\forall{\bm{x}}\in\mathcal{X}\subset\mathcal{M}$, ${\bm{x}}=\mathcal{D}(\mathcal{E}({\bm{x}}))$.
-It follows that the decoder is surjective on the data manifold, $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{M}$ [62], and the encoder-decoder composition forms an identity map, i.e., $\forall{\bm{z}}\in\mathcal{M},{\bm{z}}=\mathcal{E}(\mathcal{D}({\bm{z}}))$.
+It follows that the decoder is surjective on the data manifold, $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{M}$ (He et al., 2024), and the encoder-decoder composition forms an identity map, i.e., $\forall{\bm{z}}\in\mathcal{M},{\bm{z}}=\mathcal{E}(\mathcal{D}({\bm{z}}))$.
 Equivariance.
@@
 The equivariance properties of translation, rotation, and reflections, combined, are referred to as E(3) symmetries.
-Without reflections, the symmetries form a Euclidean group SE(3) [112, 113].
-Appendix I Computing Resources
+Without reflections, the symmetries form a Euclidean group SE(3) (Thomas et al., 2018; Fuchs et al., 2020).
+E(3), SE(3), and SO(3) are important symmetry groups in 3D Euclidean space, with well-established applications in physics and chemistry, computer vision, and reinforcement learning (Cohen and Welling, 2016; Thomas et al., 2018; Hoogeboom et al., 2022; Xu et al., 2024; Park et al., 2025).
+Finally, our contributions are complementary to, and can be combined with, the growing literature on meta-learning and automatic symmetry discovery to learn symmetry groups and their actions directly from data (Zhou et al., 2021; Quessard et al., 2020; Dehmamy et al., 2021; Mohapatra et al., 2025).
+Data manifold hypothesis.
+Let data ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be in an ambient space of dimension $d$ with support $\mathcal{X}$ distribution.
+We assume that data are sampled from a low-dimensional manifold $\mathcal{M}$ (Cayton and others, 2005; Ma and Fu, 2012) embedded in a high-dimensional space (Assumption˜H.1).
+This hypothesis is popular in machine learning (Bordt et al., 2023), and has been studied mathematically in the literature (Narayanan and Mitter, 2010; Bortoli, 2022).
+Moreover, empirical evidence in image processing supports the manifold hypothesis (Weinberger and Saul, 2006; Fefferman et al., 2016), and diffusion-based solvers assume this property (He et al., 2024; Chung et al., 2022b; 2023).
+Assumption H.1 (Manifold Hypothesis).
+Let ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be a data sample.
+The support $\mathcal{X}$ of the data distribution lies on a $k$ dimensional manifold $\mathcal{M}$ within an ambient space $\mathbb{R}^{d}$ where $k\ll d$.
+Appendix I Additional figures on MPE functions
+Table 13: DPS superresolution with $\lambda=0.01$ using different MPE functions.
+(a) FFHQ 256.
+MPE function PSNR SSIM LPIPS FID
+None 23.160 (1.923) 0.657 (0.072) 0.193 (0.057) 129.528
+LDM Encoder (FFHQ) 26.581 (2.457) 0.773 (0.044) 0.120 (0.030) 87.437
+CNN Autoencoder (FFHQ) 26.866 (1.943) 0.771 (0.044) 0.116 (0.029) 85.352
+Pretrained ResNet50 26.873 (1.941) 0.771 (0.044) 0.116 (0.029) 85.138
+Pretrained CLIP 26.860 (1.942) 0.771 (0.044) 0.116 (0.029) 85.495
+(b) ImageNet.
+MPE function PSNR SSIM LPIPS FID
+None 19.727 (4.292) 0.407 (0.180) 0.541 (0.182) 446.829
+LDM Encoder (ImageNet) 22.200 (4.295) 0.568 (0.146) 0.384 (0.130) 311.636
+CNN Autoencoder (ImageNet) 22.178 (4.294) 0.568 (0.148) 0.375 (0.125) 312.530
+Pretrained ResNet50 22.176 (4.290) 0.568 (0.148) 0.375 (0.125) 314.590
+Pretrained CLIP 22.177 (4.293) 0.568 (0.148) 0.376 (0.125) 313.468
+(a) FFHQ 256.
+(b) ImageNet.
+Figure 19: Equivariance error vs. $\sigma$ noise added. As more noise is added, equivariance error, computed with all MPE functions, increases.
+Appendix J Computing Resources
 We conduct experiments on two NVIDIA GeForce RTX 4090 GPUs with 24 GB of VRAM. We note that we use pre-trained models and perform inference, so not much compute is required.
-Appendix J Assets
-Appendix K Broader Impacts
-On the positive side, high-fidelity image restoration can improve downstream tasks in medical imaging, remote-sensing and environmental monitoring (e.g., denoising satellite observations to track pollution or deforestation).
-Likewise, accelerated PDE-solving via learned diffusion priors may enable faster, more accurate simulations for climate modeling, fluid-dynamics research, and engineering design.
-On the other hand, robust reconstruction methods could be misappropriated for privacy-invasive surveillance or to create deceptive imagery.
-We emphasize that our method does not amplify these existing risks.
+Appendix K Assets
 Appendix L Responsible Release
-Our approach uses only publicly available datasets and standard pre-trained diffusion models, introducing no novel dual-use or privacy risks. Consequently, no additional safeguards are required.
+Our approach uses only publicly available datasets and standard pre‑trained diffusion models, introducing no novel dual‑use or privacy risks. Consequently, no additional safeguards are required.